A Complete Bibliography of The American Statistician: 2000–2009

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

07 October 2024 Version 1.10

Title word cross-reference

.338 [Fre07c]. .833 [Fre07c]. 1 × 1 × r [Dun03]. 5 [Bar02a, Len01e]. α [HB03]. F [Chr03, Sho03]. gh [HR06]. H_{0} [Bet00]. k [BP02, Tar07]. M [SB02]. n [BHO00]. ⊗ [Hu07]. P [MTA08, BH02, HB03, Rya09, SBB01]. π [Jad00]. s [HJM08]. t [BM01, BHO00, GJLW05, HJM08]. Z [BHO00, HJM08].

-Estimation [SB02]. -Means [Tar07]. -Values [MTA08].

1 [Hil03c]. 1.4 [Cow04c]. 101 [Dob04b]. 102-105 [HHV^{+}02]. 11 [Hil05d]. 11th [Mil03]. 12 [Hil03c]. 12.0 [vH04]. 12.2 [Hil08f]. 121-124 [Fri02a]. 150-152 [Lev02]. 182-186 [Bar02b]. 1999 [DL04d]. 19th [LR06]. 1st [Mil03].

2000 [ESL09, SS04, WJM^{+}03]. 2nd [Cle03b, Coh02, Dob02, Fra00a, Fra00b, Gen03, Gup02, Hai03, Hyu01, Jer03, Kir00, Kre01, Led03, Man03, Sch04f, Ste01, Utt02, Wil02].

3.0 [Hil03c]. 3.1 [Hil05i]. 325 [Fri15]. 3rd [ACA03, Can01, Jam02, Jam03, Mil03, Pie01].

4.0 [Hil06b]. 4.03 [Alt02]. 4th [Cow04a, Gon03, Wal00b].

5 [Hil07a]. 5.1 [Hil02e]. 5.2 [Hil06a]. 5.5 [HH06]. 55 [Bar02b, Fri02a, Fri15, HHV^{+}02, Lev02, ZvD02]. 57 [JH04, RM04]. 58 [Les04, WCLS04]. 5th [Cle00b].

7 [Hil07f].

8.0 [Yal07]. 8.02 [Hil03c].

9 [Hil07b]. 9.0 [Hil05a, Hil05c, Hil06b].

Absolute [BS03, Bon04]. Abuse [HH01]. Academic [Moo01]. Academics [Sta01]. Accident [HS03]. Accuracy [Alt02, CLS07]. Achieving [Cle06a]. Action [Lyn03, Mil04c]. Active [Kva00]. ActivEpi [Sch04f, Sch04f]. Activities [SMC^{+}07]. ActivStats [MJ04]. Actuarial [Wüt08]. Acyclic [HKS06]. Adding [AC00a]. Adjacent [DM01]. Adjust [KP07]. Adjusted [Ann06a, BM01, LM03, LGH09]. Adjustment [And09, BSTM09]. Admission [Les04, WB04]. Ado [HK07]. Adventures [Cle03d, Pic06]. Advice [Gel04, HS04b, RSH01, Rub04b, VM03]. Afifi [Mil01]. Afraid [Tra07]. Age [BB04, Cle05a]. Agencies [Gal03]. Ages [Ano04p]. Agreeing [LS01]. Agreement [LR06, dM07]. Agricultural [ACA03]. Agriculture [Rob07a]. Aid [Lee07]. Aitsahlia [Lun04c]. Albert [Lea04, Cle04]. Algebra [Dav06]. Algorithm [DM01, Hit03, Tar07]. Algorithms [HMRW01, HL04, LR01]. All-Time [Ann06a]. Allegations [BB04]. Altered [KKO^{+}06]. Alternative [Bet00, BH02, FMA^{+}09, Ros00a]. Amelioration [MB06]. America [Tor07]. American [Alb03c, AHL04, Alb05a, Alb05b, Bar02b, Fri02a, Fri15, HHV^{+}02, JH04, Les04, Lev02, RM04, WCLS04, ZvD02, RA04, SMC^{+}07, Wal05]. Among [Bol03, XN03]. Analyses [Gup02, KBH09, Lac00, Lun07b, Man03, OH08a, OH08b]. Analysis [Ann07, Ano04b, Bon02, BGF^{+}08, Car00, Cav07b, Cle08a, Cle08b, Cow04b, DD06, Dav08b, Dav08d, DLM01, Dev07a, Dev07c, Dia05, Fre08, GL00, GGR05, HLW09, HH01, Hu07, IdBPT00, IM01, Jan00, Ker07, Len01a, Lis08, Mac01, McG05, MB06, Men00, Mil03, Nel07, Ost02, Ost03, OSC00, Owe01, PH02, Pri05b, RHCR06, Rob07b, Rue04, Rum01, Vic08, Wie06, Wil02, Woo08, Wul03, Zim07, vH04, Gen03, Hai03, Jam03, War02, AC06a, Pou02]. Analytics [Wie08]. Analyzing [Ano04i, Hal08, Han05, Kel04, Pan02, Was06]. Anderson [Ano04p]. Andreescu [DL04d]. Annalisa [Ano04s]. Anne [Mil04c]. AnotherWorld [Sne08]. ANOVA [AC04b, BSTM09]. ANOVA-Type [BSTM09]. Answer [BHA02]. AppDev [Pra08b]. APPL [GEL01]. Application [DM01, DO04, Hag09]. Applications [Abb02, Bar09, Fer04, GV09, Hai07, Hu03, JTFK04, LHD00, Lun04b, Mor07, RH02, Rid09, Sta08, Tan02, Ver01, Wie07, vDK02, DH05, Lun04a, Dev01]. Applied [And09, BGF^{+}08, Cow03, Dev07a, Kir00, Mis06, RT01, Sim07, Woo08, Fox02, Ano00b, Ano04b, Fot02, Gon03, Jam02, Wil02]. Applying [LG00]. Approach [AC04b, Cle08b, GV09, Har08, Lee08, LR06, LGH09, Mea03, MS02, RM04, Sim07, TTT09, Zim07, dML09, Fra00a, Fra00b, Len01d, Owe01, Ree02]. Approaches [Men04a, Men04b, SB09]. Approximate [FLP07, Rya09]. Approximation [CLPC08, Kuo00, Pro08]. Approximations [JB04, LJ09]. ARIMA [HT02]. Art [Arn07, Gla04]. Arthur [Gla04]. Aspects [IJW06]. Assessing [BUK^{+}08, ESL09, Haz03, Haz04, JH04, PC02, Tob04]. Assessment [FS01, Grö09, KBH09]. Assessments [Tob04]. Assisted [Len04b, Mil04a]. Associated [Wil08]. Association [SMC^{+}07, Wal05]. Assumptions [AL07, SF00]. Assurance [Hyd01]. Asymmetric [BM01, DR01]. Asymmetry [DR07]. Atlas [SW05]. Attention [Alt02]. Auctions [HJS06]. Audits [MSL^{+}08]. Australia [Lud05]. Autoregressive [WGHE09, YW09]. Average [SD01, YW09]. Avoid [AC04a, AC04c].

B [DL04c, War02, BI09]. B-Splines [BI09]. Babies [CS05]. Babylon [Ano04p]. Bachelor [BGNP01]. Back [Ano00c, Ano00d, Ano00e, Ano00f, Ano01b, Ano01c, Ano01d, Ano01e, Ano02a, Ano04c, Ano04d, Ano04e, Sch03c]. Bacon [DH05]. Bag [Cle05b]. Bagged [Ber07]. Balanced [BHA02]. Ballots [WJM^{+}03]. Bands [Ros00a]. Barker [Sch04f]. Bartko [Bar02b]. Baseball [Ann06a, Coc05b, Lea04, SG00, Cle04]. Based [APT07, BA02, CF01, Fre05, Grö07, HQ08, KBH09, LG05, MSL^{+}08, RT01, SV03, Wal00a, dML09]. Basic [AC03, ACL04, Coc05a, Dob02, Dob06a, Wal00b, Mil03, War02]. Basics [Ann05c]. Basis [Bro01]. Basketball [Gil00, RHCR06]. Batting [Ann06a]. Bayes [Chr05, Gel08, Lit06, NVB09, SS02, TGM08]. Bayesian [Ree02, AR08, BM02a, DD06, Dip04, Fre07b, GS08, GGR05, GJLW05, HMRW01, IBP05, KB07, Lee08, MR09, Rum06, Sin08, UJ08]. Be [Fre07a, FG00, BHO00, Fre07b, JW04]. Beam [DT00]. Been [Tra07]. Beginner [Nel01]. Beginners [Gel04, HS04b, Rub04b]. Beginnings [Dav08a]. Behavior [CL02]. Behavioral [Kre01, Lee08]. Behrens [DO04]. Beichelt [Lun04a]. Being [FS06, Fri15, LSO01, WSF02]. Believed [Fre07b]. Bélisle [Lev02]. Bend [WWD03]. Benefit [Hai03]. Benford [CG07, Few09, LSE00]. Benjamin [Gla04]. Berger [Dob04b]. Berlinghoff [Ano04p]. Best [Ann06a, Eli00, Gur06]. Beta [DH05]. Beta-Binomial [DH05]. Bethel [AC04b]. Bets [Alb02]. Better [Fre07c, WE08, WGHE09]. Between [DS08, EW02, GS06, IBP05, SG01, Shi01, Bar02b, Fri02a]. Beyond [Ann05c, Bel05, DS02, Nel07]. Bias [Bar07, ESL09, Gad01, GN02, HLP03, JS03, KH04, MR09]. Biases [Ros01, vH04]. Bimodal [SWW02]. Binary [Fay02]. Binomial [Afs08, BRRB01, CLPC08, CL02, DH05, FM08, HM01, Hwa09, LJ09, Pro08, Ros07]. Bioassay [DO04]. Bioinformatics [Fos08]. Biology [Rob07b, ACA03]. Biomedical [Rue04]. Biosciences [Cle01]. Biostatistics [Fer04, HFDJ08, KYA^{+}07, Zha07b, Mil01]. Biotechnology [Kha04]. Birth [SD04]. Bit [Tur07]. Bivariate [DH05, MS02]. Black [Bro07, Tal07, WH07]. Block [GL00, Pad08]. Board [Coc05a]. Boen [HHV^{+}02]. Bonus [Kes00]. Book [Abb02, AC02, AC03, ACA03, AC04b, ACL04, AC04a, AC05a, AC06a, AC08a, AC08b, Bat04, Ber00a, Ber00b, Ber00c, Bon02, Can01, Car00, Cha04, Che00, Cle00a, Cle00b, Cle03b, Cle03a, Cle03c, Cle03d, Cle04, Cob00, Coh02, Cow03, Cow04a, Cow04b, Dev02, Dix07, Dob02, Dob03b, Dob03a, Dob04b, DL04e, DL04c, DL04d, DL04a, DL04f, DL04b, Dun04, Fer04, Fot02, Fra00a, Fra00b, Gen03, Gla04, Gon03, Gup02, Hai03, Ham00, Har00, Hu03, Jam00, Jam01a, Jam01b, Jam02, Jam03, Jer03, Kan00, Kir00, Lac00, Led03, Len00, Len01a, Len01b, Len01c, Len04a, LF00, Lun02b, Lun02a, Lun04c, Lun04a, Lun07b, Man03, Mil03, Mil04b, Mil04a, Mil04c, Min07, Nel01, Owe01, Pag00, Pie01, Ree00, RH02, Rob03, Rob08, Ros00b, Row02]. Book [Rue04, Sch00, Sch04e, Sch04f, Ste01, Sti00, Tan02, Ter03, Utt02, Ver01, Wal02, Wal00b, Wil02, Wul03]. Books [Ano02g, Ano09g, Ano09h, Ano09i, Ano09j, Cle01, Col01, Dev01, Hai01, Hil01, Hyd01, Hyu01, Jam01d, Jam01c, Kre01, Len01d, Len01e, Len02, Mac01, Mil01, Pou02, Ree02, Rum01, Sha01, Str02, War02]. Bootstrap [AT04, DK07, FF02, JS03, RZR05, TKM04, Fer04]. Borel [SS01]. Borovkov [Sch04e]. Bottles [TH00]. Bowditch [RA04]. Bowler [DPS04]. Box [DR07, FLP07]. Box-and-Whiskers [DR07]. Boxer [Gel06]. Bragged [Ber07]. Breaking [CG07]. Breviary [SW05]. Brian [Sha01]. Bridge [Shu01]. Brief [Ano03a, Ano04f, Ano07b, Ano07c, Ano07d, Ano08b, Ano08c, DL04g, DL05, DAC05, DAC06, RACK07, RD05]. Brown [Les04]. BS [RSH01]. Buffett [FS06]. Build [Bro08]. Bunker [Col01]. Business [AR08, Chi08, Cle08b, Dev06c, Dev07a, Dev07b, Edw07, LH02, Mil01, Owe01, Sim07, Skl07, SB09, Wal02].

C [Dob04b, DL04c, Mon06]. Calculated [Alb02]. Calculating [Bro01, Men04a, Men04b]. Calculation [AC00b]. Calculations [Bro01, HH01, PB03, SRC07]. Calculator [Mon06]. Calculus [Hag09, Hwa09, SB02, Wan02]. Calibrated [Lit06]. Calibration [Fin06, SBB01, YHZ08, Zha00]. California [Tob04]. Called [Fre06]. campaign [CG07]. Can [Ann06a, Coc05a, Fre07a, GN02, Sch03c, Tra07, Utt03a, VM07]. Candid [WSF02]. Can’t [GN02, Ros00b]. Capital [Rum07]. Carlo [FF02, KBH09]. Carolina [Mon04, Ste08]. Cars [Ano04s]. Case [BA02, Cal03, Kos05, PHSN08, Shu01, SB09, Tob04]. Case-Control [PHSN08]. CAST [Len04b, Mil04a]. Categorical [LBTM03, Ost02, Ost03, Was06]. Categories [MFS^{+}03]. Causality [Ros05]. Causation [Fri15]. Cautionary [AH02]. Caveats [Gam09]. CBMS [SS04].

CD [Bon02, Mil03, Sch04f]. CD-ROM [Mil03, Sch04f]. Ceiling [AB03]. Censored [BH08, FM08]. Censoring [Man07, SD01]. Census [Cra08, CHS04, Tor07]. Central [Sko08, Bel01]. Century [DL00, LR06, Cle03c]. Chakraborti [DL04f]. Challenge [AC06b, Bra07, Cle06a, Fin06, Hol07, Nor07]. Championship [Ano04q, Bil04, Col04, Har04, Mas04, Mea04, SMH^{+}04]. Chance [AC05b, Cle06b, Gue06, Sch06b, Hai01, Mac01]. Chances [Cle03d]. Change [BA06, Zho01]. Changing [Tor07]. Characteristic [BK07, DG07, Hal08]. Characterization [Hu08]. Characterizing [EKG02, FRD07]. Charles [MS01]. Chart [RHCR06, Wil06a]. Charts [Bor02, LHD00, SF00]. Chebyshev [AC00b]. Chen [ZvD02]. Chernick [Fer04]. Chi [Mir01, WN07]. Chi-Squared [Mir01]. Chicago [AC05a, AC06a]. Chip [Lee07]. Chisini [GV09]. Chocolate [Lee07]. Choice [BHA02, Wal00a]. Choice-Based [Wal00a]. Cholesky [EW02]. Choosing [LS08]. Christopher [Mac01]. Chung [Lev02, Lun04c, WCLS04]. Citizens [Utt03b]. Claims [Utt02]. Clark [Ano04b]. Clarke [Rue04]. Class [Dip04, HLW09, NVB09]. Classes [AL07, SB09]. Classical [HB03, WMAY09]. Classification [HV03]. Classroom [BG08, Shu06, Ter05, DL04c]. Clear [Ros00b]. Client [Sow03]. Clients [JW04]. Clinical [CRM01, De 06, DL04a]. Clinicians [Lis08]. Clowns [Ano04s]. Cluster [HLW09, WF09]. Clustered [Pan02]. Clustering [Tar07]. Clusters [Moe05a, Moe05b]. Cobb [Mil01, Mil04c]. Coefficient [DR01, Fri02a, Mir01, Shi01, Tju09]. Coefficients [Fri02a, LM03, Men00, Men04a, Men04b, Shi01, Tju09]. Coffee [Ano04s]. Cognitive [LG00]. Coherent [SRMP08]. Coin [CS05, GN02, Gel06, KB07, Shu06]. Coincidence [Ste08]. Collaborators [Ano00a, Ano01a, Ano04a, Ano05a, Ano06a, Ano07a, Ano08a, Ano09a]. Collapse [MFS^{+}03]. Collection [War02]. Collections [BM03, SGK00]. Collector [Kuo00]. College [Alb03a, Ano04q, Bil04, Col04, Dun04, Har04, Mas04, Mea03, Mea04, Par04, RM04, SMH^{+}04]. Collinearity [FK09]. Column [Alb03c]. Combinatorial [Gla04]. Combining [CRM01, SMC^{+}07]. Comment [Bar02b, Fri02a, Fri15, GR09, HS09, Joh09, JH04, Les04, Lev02, MG09, NL09, RM04, WCLS04, ZvD02]. Comments [HHV^{+}02]. Common [AC04c, AC04a, DL04a]. Communication [Rie06]. Community [RT01]. Community-Based [RT01]. Compact [MY05]. Companion [Cow03, Mil01, Sch04f, Fox02]. Comparing [ALTY00, Cal03]. Comparison [AW06, Ano04i, Bar02a, HT02, HL01, IdBPT00, KB07, TGM08, tWGB01]. Comparisons [BA06]. Complex [Rid09]. Complex-Step [Rid09]. Components [GR04]. Computation [Nie05]. Computational [AG04, Car00, Gen04, Har08, Lan04]. Computations [Bar09]. Compute [GV09]. Computer [CF01, Car08, Cow06, GC04, Len04b, Mil04a, Owe01, Rob07b]. Computer-Assisted [Len04b, Mil04a]. Computer-Based [CF01]. Computer-Intensive [Rob07b]. Computers [Alb02]. Computing [Alt02, Cow04b, DC01, Gen04, HM01, Hil02c, Hil02d, Hil04f, Hil07g, HL01, Hun05, LR01, Mon04, Sko08, Zim00]. Conceptions [Alb03a]. Concepts [BM02b, IM01, Kre01, Len02, Lun02b, Mor07, dML09]. Conceptual [Bro08, Kah06]. Concerning [Sen08, SS09]. Concise [Cra05]. Concurrency [HJS06]. Condition [Hu08]. Conditional [BW08, De 08, HHvdWZ06, OI07, Pan02, Upt00, VSL01]. Conditions [CBH09]. Conferences [LH02]. Confidence [Afs08, AC00a, And09, Bar02a, Bar09, Bar02b, BR03, BS03, Bon04, CRM01, CL07, DC01, HM01, HHV^{+}02, LR08, MG04, New01, PE05, Ros03, Ros00a, Rya09, SG01, VH05, WBT01, WS01, ZG00, Cle03b]. Confidentiality [KKO^{+}06]. Confirmatory [Woo08]. Conflict [MR09, Spi07]. Confounder [JTFK04]. Confounding [Lyn03]. Confuse [AC05b]. Confusion [HB03]. Consequences [BHA02]. Conservative [WSF02]. Consistent [LE00]. Constrained [MV07]. Construction [And09]. Consultants [GB06]. Consulting [JW04, AC02]. Contemporary [Wal02]. Context [Bol03]. Contingency [Mir01]. Continuous [LBTM03, Rak06]. Contour [DG07]. Contract [Shu01]. Contrasts [Lan06b]. Contribution [RA04]. Contributions [Kha04]. Control [CLR13, JTFK04, LC09, PHSN08, SF00, AC08a]. Controversies [Utt02]. Convergence [TKM04, dML09]. Conversation [Sin00]. Conversion [Hil07a]. Cookies [Lee07]. Corner [BHM01, CF01, CR01, HJP^{+}01, HM01, HGA01, IM01, Jan00, RT01, Zho01]. Correction [Ano03b, Ano03c, Ano04g, Ano05b, Ano05c, Ano06b, CLR13, JS03, Men04a, Upt00]. Corrections [Ano02b, Ano03d, Ano04h, Ano08d, Bon04, Haz04, Len04b, Str04]. Correlated [Fri15, IdBPT00, Kel04, LSO01, PLC00, SC09]. Correlation [And09, Dav09, DR01, Fri02a, Fri15, Fri02b, HCW07, RA04, Shi01, Zim00]. Correlations [Lan06b, MS02]. Corrgrams [Fri02b]. Count [Gla04]. Counterfactual [SGM07]. Coupon [MS01]. Course [AC08a, BHM01, Cra05, Dav07, Dev02, Dev07a, Gel05, HM01, Len01d, LP00, Mac01, Mar07, Mru07, Pou02, Pri05b, Rob03, SV03, Wal00b, Wul03, Cha04, Row02]. Courses [Ano04s, CF01, FR04, Gen04, Hun05, LJLC05, Par04]. Cousins [BSTM09]. Covariate [AH02, SJI05]. Covariates [Zho01]. Coverage [LS08]. Cox [FLP07, Zho01]. Cramer [Kag01, Nay02]. Crannell [Ano04s]. Crash [Dav07, SR04]. Crashed [Ano04s]. Crisis [Dun04]. Criteria [GR04, HCW07, VH05]. Crossing [FRD07]. Crude [LGH09]. Cryer [Mil01]. Current [Hil03c, WY04]. Curriculum [BGNP01]. Curves [BK07, Hal08].

D [Ano04i, Dob04b, Mil01, RM04, Sch04f]. Daily [BG08]. Damned [AC05b]. Data [AW01, AC07c, Ann07, Ano04i, Arn07, Bon02, BH08, BGF^{+}08, Cav07b, CF07, Cle08a, Cle08b, Cow04b, Dav08b, Dev07c, Dia05, DT00, FM08, GL00, Gen03, GR04, GC04, Han04, Han05, HDE^{+}03, HQ08, HH01, Kah06, KKO^{+}06, Kel04, Kha08, Len01e, Les04, Mac01, Mil03, Ost02, Ost03, OH08a, OH08b, OSC00, Owe01, PLC00, Pan02, Pri05b, Rak06, Ree02, RHCR06, Ric08, Rob07b, Rum01, SS08, SMVD09, Spa08, Su08, WWD03, WB04, War02, Was06, Wie06, Wie08, Wil02, tWGB01, Zim00, Zim07, Mil04c, Rue04, Car00]. Databases [OI07]. Datasets [Bra08a]. David [DL04b, Sch04f]. Dawid [Sen08]. Day [Cha04]. Death [Gue06, Lud05, SD04]. Deaths [Cle03d]. Debate [CR01]. Debating [BB04]. Deborah [Dev01]. Decision [Vic08, WY04]. Decisions [CLS07, Cle08b]. Decline [VD05]. Decomposition [EW02, Grö07]. Deem [JW04]. Deeper [Mil04b]. Degree [JW04]. Degrees [BGNP01, FMA^{+}09]. Degrees-of-Freedom [FMA^{+}09]. Delaney [Ano04i]. Delivery [Pra08b]. Delta [SRMP08]. Deming [Hah02, Sne08, SF00]. Demonstrate [SF00]. Demonstrations [Mil04b]. Density [CC06, FRD07, SA03]. Department [Sti08]. Dependence [FS01, GB03, MG04, Shu01]. Dependency [Sab02]. Derivation [GS02, Sol00]. Derivations [Bet00]. Derive [Kuk09, VE06]. Description [BCB00]. Descriptive [Dav08d]. Design [AC07a, Arn06, BM02a, Ber06, BGF^{+}08, DLM01, DO04, Eas04, Gam09, HS03, IM01, Len01a, Len01d, Lis08, Mru07, OF03, Pie01, Pra08a, Rob07a, Rub04a, Tan02, Wul03]. Designing [FF02, Ano04i]. Designs [HMRW01, Pad08]. Desired [Men09, Stu09]. Detect [AW01]. Detecting [Bar07, GB03]. detection [CG07]. Determination [Afs08, Fri02a, IBP05, Len01f, LM03, Men00, Shi01, Tju09]. Determine [LB00]. Developing [AC06b, AT04]. Development [BGF^{+}08, Zam08]. Deviation [HJP^{+}01, Lev02]. Deviations [BS03, Bon04]. DHEA [EKG02]. Diagnostic [Vic08]. Diagnostics [FK09]. Diagram [Upt01]. Diamond [LBTM03]. Diamond-Shaped [LBTM03]. Diamonds [DL04e]. Dicing [Gue06]. Dickinson [DL04f]. Diebold [Hyu01]. Difference [GS06]. Differences [AC00a, Bar02b, Les04, Man07, SG01, WB04]. Different [BLS00]. Differential [Hwa09, Wan02]. Differentiation [Rid09]. Difficulties [CLS04, WCLS04]. Digit [Rod04]. Digital [BGF^{+}08]. Digits [Jad00]. Dimension [GR04]. Dimension-Reducing [GR04]. Dimensional [CW04]. Dimensions [TKM04]. Direct [HT02]. Directed [HKS06]. Discipline [BM02b]. Disciplines [KC09]. Discovery [Ber06, Kha04, Pic06, Ree02, Ric08]. Discrete [LR08, MF07, Ter05, VH05, Wal00a]. Discriminant [LGH09]. Discrimination [BB04, Tju09]. Discussion [Bel05, Ber03, Bil04, Bra07, Car03, Col04, Gel04, Har04, Mas04, Mea04, Nor07]. Discussions [SB09, SMH^{+}04]. Disease [MB06]. Dispersion [MY05]. Display [BM03, LBTM03, Pri05b]. Displaying [Jac08]. Displays [ALTY00, DT00, Fri02b, HT02]. Dissimilar [CF07]. Distance [HFDJ08]. Distributed [Jad00, JB04, tWGB01]. Distribution [CLPC08, DH05, Dav05, GS02, Hag09, HR06, LM08, LJ09, LR01, LB00, MF07, OSC00, SA00, Sko08, SRMP08, Wil08, vDK02]. Distributions [BM03, Jam00, LS01, LSE00, MS02, OI07, Rod04, SS02, VE06, Wie07, XN03, ZG00]. Diurnal [EKG02]. Do [GB06, Men09, Mon06, SG00, Stu09, TS03, Los04a]. DOE [BA02]. Does [BHO00, FDC09, MFS^{+}03]. Doing [SHS04]. Domains [Tal07]. Double [Rya09]. Doughnut [Bor02]. Draft [DM01]. Drawing [CLS04, WCLS04].

Driven [Zim07]. Drug [Kha04]. Drug-Discovery [Kha04]. Drugs [BHM01]. Dual [Nad09]. Duelling [Cle03d]. Dunn [Ano04b]. Dyadic [Ann07]. Dyk [ZvD02]. Dynamic [AC07c, Col01, IM01].

Early [APT07, Bet00]. Easy [CLR13, LC09, Len01e]. Ecological [For05]. Economic [Dev07a]. Economics [Dev06c, Edw07, Gel08, RH02, Sha01]. Ed [Utt02, ACA03, Can01, Cle03b, Coh02, Cow04a, Dob02, Fra00a, Fra00b, Gen03, Gon03, Hai03, Jam03, Jer03, Kir00, Kre01, Led03, Man03, Mil03, Pie01, Sch04f, Ste01, Gup02]. Edition [Wil02, Hyu01, Sha01, Jam01c, Cle00b, Wal00b, Bon02, Jam02]. Editor [Alb03b, AHL04, Alb05a, Alb05b, Ano00l, Ano00m, Ano00n, Ano01j, Ano02d, Ano02e, Ano02f, Ano03e, Ano03f, Ano03g, Ano03h, Ano04m, Ano04n, Ano04o, Ano05d, Ano05e, Ano06c, Ano06d, Ano06e, Ano06f, Ano07e, Ano07f, Ano07g, Ano08f, Ano08g, Ano08h, Ano08i, Ano09d, Ano09f, Ano09e, BCW01, DBD^{+}01, GGF^{+}01, HV09, Hil00a, Hil00b, Hil02c, Hil02d, Hil02a, Hil02b, Hil03a, Hil03b, Hil04a, Hil04c, Hil04d, Hil04e, Hil05e, Hil05f, Hil05g, Hil05h, Hil06c, Hil06d, Hil06e, Hil06f, Hil07c, Hil07d, Hil07e, Hil08c, Hil08d, Hil08e, Hil08b, Hil09, HB05b, LBH^{+}00, LDH^{+}00, LIdBP^{+}00, Ost09a, Ost09b, SOS^{+}01, SHD^{+}00, Sto01, Sto02, SA02, Stu09, Tuy09, Wes07, Wes08]. Editorial [Ano00a, Ano01a, Ano04a, Ano05a, Ano06a, Ano07a, Ano08a, Ano09a, Wes06]. Educated [Utt03b]. Educating [Sow03]. Education [Cle06a, Gel08, Kre01, LH02, SS04, Wil06b, Dun04]. Edwards [Sne08]. Effect [AB03, Gam09, KP07, Kva00, MG04, Tim04, YT01]. Effective [AC00a, FMA^{+}09, HHV^{+}02, Len01f, LH02, WBT01]. Effects [CK01, GL00, LBTM03, MB00, Ros01, Ros03, SGK00, SF00, ZvD02]. Efficiency [YT01]. Efficient [DO04, NVB09]. Eggs [DH05]. Einstein [Igl07]. Election [CF08, WJM^{+}03]. Electronic [HS03, Mil01]. Elementary [Bro08, Cha08, Mil03, Rum01, YA09, DL04b, Lun04c]. Elements [RH02, Hyu01, Sch04e]. Eleven [Mir01]. Elicitation [BI09]. Eliciting [SGM07]. Else [AC05b, LB00]. Elyn [Ano04s]. Employers [RSH01]. Encounters [Mac01]. Engineering [AC08b, Hu03, Kva00, Tan02, War02, vDK02, Jam00, Kir00]. Engineers [Cle08a, Kie06, Ste01, Kir00, Car00, Hyd01]. Enhance [RT01, Sow03]. Enough [BH08]. Ensembles [Zhu08]. Enterprise [Dix07, Spe08]. Entry [DM01]. Environment [SM09]. Environmental [DL04c]. Epidemiology [Gil08, Str04, SGCS04]. ePrint [Hil07a]. Equal [Sho03]. Equality [RZR05]. Equations [tWGB01]. Equiponderant [LBTM03]. Equitable [BS00a]. Equivalence [BRRB01, CLR13, EW02, LC09]. Erickson [Col01]. Errata [Ano09b]. Error [AB03, KBH09, Nay02, VD05]. Errors [AC04a, AC04c, Dem06, Fre06, Fre05, Gad01, HB03, LE00, SC09, Koe09]. Essence [Dob06a]. Essential [BUK^{+}08, Dev06a, Can01]. Essentials [Edw07, Mil03]. Estimable [PE05]. Estimate [Bar09, WE08]. Estimates [JP01, SA03]. Estimating [LGH09, SD04, Tim04, tWGB01]. Estimation [BS00b, Boy08, BH08, CL07, CK01, HS03, OI07, SB02, Wal00a, ZvD02]. Estimator [Fre06, SD01, SJI05]. Estimators [Grö07, Lee09, YT01]. Ethics [VM03]. Evaluating [GR04, KKO^{+}06]. Evaluation [MJ04, Vic08, VH05, WY04]. Evans [CHS04]. Even [Tra07]. Event [GC04, LHD00]. Events [Bar02a]. Ever [Tra07]. Everyday [Ros00b, Sch06b]. Everyone [Tra07]. Everything [AC05b]. Evidence [APT07, BB04, HB03]. Evidence-Based [APT07]. Evidential [SRC07]. Evolution [UJ08]. Evolutionary [Nie05]. Exact [CL02, Hil02e, KR02, Ost02, Ost03, OH08a, OH08b, Ros03]. Examination [LJLC05, Ost02, Ost03, OH08a, OH08b]. Examining [Bar02b, SG01, WGHE09]. Example [BH08, CLS04, ESL09, Gur08, JS03, LF00, TKM04, WCLS04]. Examples [Dob06b, Pra08b, Pri05b, Ter05, DL04a]. Excel [LF00, Wal02]. Excerpt [Dem06, SD04, SW05]. Exchangeable [LS01]. Exemplar [Bul05]. Exercise [DLM01]. Exercises [ACL04, Bel07, FZ00]. Existence [KC02]. Expected [Bar02a, CRM01]. Experiences [BG08, JW04, SV03]. Experiment [GL00, Pad08, SF00]. Experimental [ACA03, AC07a, DO04, Eas04, HMRW01, IM01, Mru07, Pie01, Tim04, Tan02]. Experimenters [Ber06]. Experiments [Ano04i, Arn06, Koe09, Law03, Len01a, Len01d, Pra08a, Rob07a, Wul03]. Explained [Moo07]. Explaining [CLR01, Lan06b]. Explanation [Few09, Paw00]. Exploration [Cav07b, HBQ04, YCF06]. Exploratory [Dev07c, Fri02b, Tim04]. Explore [DT00]. Exploring [HM01]. Exponential [ML09]. Extend [Sin00]. Extended [SJI05]. Extensions [LHD00]. Extent [SD04]. Exuberance [SR05].

F [Bar02b, Mac01, Rue04]. Factor [GL00, Woo08]. Factorials [SGK00]. Factors [JW04, SS02, WGHE09]. Failures [AC00a]. Fall [VD05]. Fallacy [HH01]. Fallibility [Paw00]. Families [ML09]. Family [Han04, WWD03]. Farid [Lun04c]. Fascination [Coc05b]. Fast [LR01]. Fathom [Col01]. Fathoms [Mil04b]. Fatti [Lun04a]. Favor [Bet00]. Favorites [FDC09]. Feared [Men09, Stu09]. Features [SMC^{+}07]. Fences [DR07]. Feng [DL04d]. Fernando [Ano04p]. Fetterman [AC04b]. Fiducial [Wan00]. Field [Ann06a]. Fields [Utt02]. Fifth [Sha01]. Fifty [Mil04b]. Figure [WY04]. Figures [HQ08]. Filter [EW02]. finance [CG07, Lun04c]. Financial [Dia05, vDK02]. Finding [HMRW01]. Finite [CLS04, KT02, LaM08, Ter05, WCLS04, WS01]. Finite-Mixture [CLS04, WCLS04]. Finzer [Col01]. First [BHM01, Cle03c, Dev02, HM01, Igl07, KC02, Len01d, Mac01, Mar07, Mru07, Rod04, Wul03]. First-Year [BHM01]. Fisher [Box05, Chr05, DO04, Lud05]. Fit [PC02, Ros00a]. Fitting [HV03]. Fixing [Sho03]. Flexible [PHSN08, Rub04a]. Flip [CS05, Gel06]. Florida [WJM^{+}03]. Football [AW06, Ano04q, Bil04, Col04, Gil00, Har04, Mas04, Mea03, Mea04, RM04, SMH^{+}04]. Forecasting [IJW06, LS01, Hyu01]. Forensic [Wei07]. Forest [Grö09]. Forgotten [MF07]. Forms [Eli00].

Formula [Sko08]. Forward [Sch03c]. Foundations [McC01]. Fraction [Fre07b]. Frailty [DJ05]. Frames [PHSN08]. Framework [KKO^{+}06, PO07]. Francis [Hyu01]. Frank [Lun04a]. fraud [CG07]. Freedom [FMA^{+}09]. Frequency [SMC^{+}07, SS08]. Frequentist [IBP05, MR09]. Fresh [LGH09]. Friedman [Fri02a, Fri15]. Friendly [CG06, Ros00b]. Friis [Fer04]. Front [Ano00g, Ano00h, Ano00i, Ano00j, Ano01f, Ano01g, Ano01h, Ano01i, Ano02c, Ano04j, Ano04k, Ano04l]. Frontiers [Loh04]. Fulcrum [DT00]. Function [HT06, LGH09, Pal08]. Functions [DG07, LJ09, PE05, Ros00a, VE06]. Fundamental [Har08]. Fundamentals [Cle08a, Zha07b]. Funnel [SF00]. Fusaro [DL04c]. Future [Moo01].

G [Fri02a, Mil01, Sch04f]. Galton [Han04, WWD03]. Gambling [Alb02, AC05b]. Game [BH08, Coc05a, Coc05b, Gil00, GDL08, Shu01, Shu06, Rum07]. Games [ESL09]. Gamma [Hag09]. GAUSS [Yal07, Cle05a]. Gavin [Ano04s]. GEE [HCW07, tWGB01]. Geisser [BSTM09]. General [DS08, Eli00, FG00, Rub04a, SC09, Ver01, Wil08]. Generalized [DD06, DJ05, Hu08, Hu03]. Generating [CF07, VE06]. Generation [DL00, Jon02]. Generators [Tur07]. Genetic [HMRW01]. Genetics [Sis07]. GenStat [Hil07b]. Gentle [Ano04p, OF03]. Gentleman [Bar02b]. Geometrically [JB04]. George [DL04d, Mac01, Mil04c]. Gerald [Sha01]. Get [Sta01]. Getting [Sow03]. Gibbons [DL04f]. GIS [Cra08]. Give [LF06]. Given [ALTY00]. Glenn [DL04a]. Global [HV03]. Go [Ano04s, VM07]. Goel [JP01]. Gold [HLW09]. Golf [BS00a]. Good [AC04a]. Goodness [Ros00a]. Gosset [HJM08]. Gossip [Cle03d]. Gouvêa [Ano04p]. Grade [Dun04]. Graduate [GS08, Gel08, HR05, Moo05, Ter03]. Graphic [Pic06, Ros00a]. Graphical [FS01, FW05, Haz03, Haz04, IM01, JH04, LBTM03, PC02, SW05, dML09]. Graphics [AC07c, Bra08a, CF07, Pra08b, WC07]. Graphs [GPD02, HKS06]. Greenhouse [BSTM09]. Group [Les04, WB04]. Groups [Par04]. Growth [Bel05, Rum07]. Guide [AC05a, AC05b, AC06a, Bra08b, Cle06b, Dev07c, Dix07, Dob06b, Gen03, Gil08, Gur08, Lis08, Moo07, Ros00b, Spe08, Tra07, Utt02, ACL04, AC03, Hil01, Nel01]. Guided [KC09]. Guidelines [BGNP01, Fin06, Len01f]. GUM [Los04b].

H [Cha04, Fer04, Les04, WCLS04]. Haenszel [Upt00]. Hamada [Gel04]. Handbook [Cle06c, Dav08b, Gup02, Lac00, Lun07b, Man03, Wie07]. Handicapping [BS00a]. Handling [Rub04a]. Hands [DPS04]. Hanley [Lev02]. Hanna [Dob04b]. Happens [Dun03]. Hardin [AC04a]. Harold [Ano04i]. Harvard [Sti08]. Hastings [Hit03]. Hatcher [ACL04, ACL04]. Hawkes [Hil08a]. Hazelton [JH04]. Health [Dav08c, Fer04, Gel08, Gue06, Str04, SGCS04, Utt02]. Heat [WF09]. Height [SWW02]. Helicopter [Ann05b]. Help [MFS^{+}03, Utt03a]. Helping [Bro08]. Heterogeneity [DJ05, Ros05]. Heterogeneous [JB04].

Heteroscedasticity [LE00]. Higgins [Min07]. High [SR04, TKM04]. High-Tech [SR04]. Higher [Cle06a]. Highlighting [Man07]. Hilfertys [Koe09]. Histograms [SS08]. Historical [Dav09, Jam01c]. History [Ano04p, Hit03, Jam01d, Len02, WF09]. Hitter [Fre07c]. Hitters [Ann06a]. HIV [BW08]. Hockey [DM01, Gil00]. Holmes [Ano04p]. Home [Ann06a, LP00]. Homogeneity [GGR05]. Honsberger [DL04e]. Hot [DPS04]. Hour [GS08]. Huber [Fre06]. Human [Han04, SWW02]. Hurricanes [IJW06]. Hyper [NVB09]. Hyper-Efficient [NVB09]. Hypergeometric [AC00b, LR01, MF07]. Hyperstat [Jer03]. Hypotheses [SBB01]. Hypothesis [BH02, BH08, FF02, Jan00, KT02].

I-A [AW06]. Ideas [BUK^{+}08, Dob02, Pag00]. Identical [CF07]. Identically [Jad00]. Idiots [Cle03d]. if [HJM08]. II [Ost03, VD05]. I’ll [LF06]. Illustrated [Les04, WB04]. Illustrates [JS03]. Illustrating [BR03, BP02, SJI05]. Illustrative [Ter05]. Impact [SJI05]. Implementations [FF02]. Importance [Grö07, Grö09, HB05a, Tho04]. Improve [YHZ08]. Improved [CLPC08]. Improvement [AC07b]. Improvements [BS00b, WY04]. Improving [HQ08, YA09]. Imputation [HR06, HL01, HLP03, Sch05e, YHZ08]. In-Sample [Fre05]. Including [Fer04]. Inconsistent [Fre07a]. Incorporate [SA03]. Independence [DS08, Dav09, De 08, Upt00]. Independent [Jad00, Mea03, RM04]. Index [Ano00k, Ano08e, Ano09c]. Indices [dM07]. Industry [Cle08b]. Inequalities [Gho02, Nay02]. Inequality [Fri02a, Kag01, Shi01]. Inference [Cra05, Fre07b, Gad01, Hil02e, HHvdWZ06, NVB09, DL04f, Mac01]. Inferences [CLS04, WCLS04]. Inflation [AB03, Dun04]. Influence [Fay02, SC09]. Information [Ano00o, Ano01k, Ano04r, SA03]. Informative [WSF02]. Informed [Cle08b]. InfoTrac [Mil03]. Innovation [Ber06]. Innovations [Cle06a]. Insignificant [Chr03]. Instruction [BA02, LG00]. Instructors [SB09]. Insurance [Ver01]. Integrated [AC04b, Owe01]. Integration [DG07]. Intensive [Rob07b]. Interactions [SGK00]. Interactive [AC07c]. Interesting [JS03]. Intermediate [Pri05b, Row02, Fra00a, Fra00b, Rum01]. International [DL04d]. Interpretation [Dav08b, HQ08, Les04, Str02, WB04]. Interval [And09, BS00b, Boy08, CRM01, CL07, FM08, HM01, HHV^{+}02, WBT01]. Interval-Censored [FM08]. Intervals [Afs08, AC00a, Bar02a, Bar09, Bar02b, BR03, BS03, Bon04, BHO00, LS08, LR08, MG04, New01, PE05, Ros03, Rya09, SG01, SS08, VH05, Wan00, WS01, YA09, ZG00]. Intro [McD04]. Introduce [KS08, KC09]. Introduction [CG06, Chi08, Cle08a, Cle08b, Coh02, Cow04b, Dav06, Dev07b, Fos08, Har06, Ker07, Len01a, Lun04c, Mäc07, McG05, Nel07, OF03, Pra08a, Rob07b, Rum06, Sch03a, Sch03b, Sch04a, Sch04b, Sch04c, Sch04d, Sch05a, Sch05b, Sch05c, Sch05d, Sch06a, Sin08, SW05, Str02, SRC07, Stu09, Wei07, Cle00b, Mil03, Jam03, Lun02a, Min07, Owe01, Ste01, Ver01]. Introductory [AL07, CF01, CR01, FR04, HR00, Moo07, SB09, Bat04, Fer04]. Intuitive [Rus08]. Inventions [SW05]. Inverse [Hu08, New01, SD01]. Inverse-Probability-of-Censoring [SD01]. Inverting [Ros03]. Investigating [HJS06, Mor07]. Investigation [KR02]. Investigations [Cle01]. Invitation [Stu09]. Involving [HS03]. Iowa [FDS05]. iPod [FDC09]. Iraq [Spi07]. Irrational [SR05]. Irregular [SS08]. Island [Meh09]. Issue [Gra08]. Issues [AC05b]. Itching [MB06]. Itself [GS06].

J [Bar02b, Gla04, HHV^{+}02, Lev02, Mac01, Mil01, Min07, War02, WCLS04]. Jacobian [DD06]. James [AC04a, Min07]. Java [Pra08b]. Jean [Ano04b, DL04f]. Jeffreys [TGM08]. Jennifer [Gla04]. Jill [Col01]. Jim [Cle04, Lea04]. JMP [Alt02]. Job [Sta01]. Jobe [War02]. Johnson [Dun04]. Joint [GC04]. Joseph [Lev02]. Journalists [Cle06c]. Journals [TS03]. Jr. [DL04d]. Judging [Bar02b, ESL09, SG01]. Jury [Bar07]. Just [AC05b, Mil03].

Kai [Lun04c]. Kalman [EW02]. Kaplan [Bar09, SD01, SJI05]. Kappa [dM07]. Kappa-Type [dM07]. Karl [Cle05a]. Katz [Ano04p]. Keith [AC04b]. Keller [Sha01]. Kendall [DF07, GB03]. Kenschaft [DL04c]. Kernel [PW06]. Kernels [Zhu08]. Kevin [Sch04f]. Key [BHA02]. Key-Balanced [BHA02]. Khintchine [Jon02]. Kleinbaum [Sch04f]. Know [Ter03, Utt03b]. Known [DS08]. Kuczma [DL04d]. Kuo [ZvD02]. Kurtosis [KS08].

L [JH04, Les04, Lev02, Lun04a, Mil04c, WCLS04, ZvD02]. Lab [CF01]. Labs [Dev01, KC09]. Lai [Lun04c]. Langford [Fri15]. Language [Bro07, GEL01]. Laplace [TGM08]. Laptops [BG08]. Large [BR03, BHO00, Bra08a, OI07]. Larose [Ano04s]. Larry [ACL04]. Last [SS06]. Late [Gil00]. Late-Game [Gil00]. Latent [HLW09, HLW09]. Law [BR03, CG07, Few09, Koe09, LSE00]. Laws [Gra08]. Leaf [SMC^{+}07]. League [DM01]. Learn [Coc05a, Tra07]. Learning [Arn07, BGF^{+}08, Gur08, Hil01, Hil08a, Kva00, MJ04, RT01]. Least [Lee09, Tar00, deL06]. Lee [DL04d, Mil01]. Lego [HHvdWZ06]. Lemma [BP02]. Length [BCB00]. Lesser [Les04]. Let [GPD02]. Letter [Ano09d, Ano09e, HV09, SS09, Tuy09]. Letters [Ano00l, Ano00m, Ano00n, Ano01j, Ano02d, Ano02e, Ano02f, Ano03e, Ano03f, Ano03g, Ano03h, Ano04m, Ano04n, Ano04o, Ano05d, Ano05e, Ano06c, Ano06d, Ano06e, Ano06f, Ano07e, Ano07f, Ano07g, Ano08f, Ano08g, Ano08h, Ano08i, Ano09f, BCW01, DBD^{+}01, GGF^{+}01, HB05b, LBH^{+}00, LDH^{+}00, LIdBP^{+}00, SOS^{+}01, SHD^{+}00]. Level [Ann06a, CF08, Gel05]. Levels [GL00]. Levine [Lev02]. Licensing [Les04, WB04]. Licks [SS06]. Lies [AC05b]. Life [Cle05a, Cle06b, Kha08, Lud05, Moo07, Sch06b, Ros00b]. Lifelong [Coc05b]. Lifetime [OSC00]. Likelihood [Boy08, JP01, KR02, LG05, Mea03, MV07, Pan02, Paw00, RM04, WS01]. Likelihoods [BM03]. Likely [PP09]. LIMDEP [Hil06b]. Limit [Bel01]. Limited [Rak06]. Linear [CBH09, Cow04a, DD06, Eli00, FW05, Grö07, Grö09, Gur06, Hu03, Jam03, Len01d, LE00, MB00, PLC00, PE05, Pri05a, SC09, Tar07, Tim04, VSL01, XN03]. Lines [MB00]. Link [CG06]. Link-Probit-Normal [CG06]. Literacy [AC06b, Cle06a, Gal03, Utt03a]. Literary [GGR05, PH02]. Little [Dix07, Len01e, She07]. Liver [MB06]. Lives [She07]. Load [GN02]. Local [HV03, Iri01]. Locally [Kuk09]. Location [HJP^{+}01, Lev02]. Logistic [Fay02, KR02, LM03, Men00, Men04a, Men04b, MFS^{+}03, PC02, Tju09]. Logit [CK01, New01, ZvD02]. Lognormal [SA00, Sko08]. Logrank [Cal03]. LogXact [Har00]. Long [BH08, HR00]. Longitudinal [GC04, Hor06, SMVD09, Zim00]. Look [Kag01, LGH09, Mir01, Nas06, Sch03c]. Looking [Sch03c]. Lottery [Ste08]. Love [AC05b]. Lower [GP09]. Lucullus [Mis06]. Lyell [MS01].

M [Ano04b, Fri15, Gel04, HHV^{+}02, JH04, Les04, Lev02, Mil01, Sch04f]. Macintosh [Han05]. Main [SGK00]. Making [CS05, Dev07c, LH02, Sne08]. Management [Sha01, Tan02]. Managing [SB09]. Mann [BLS00, Ran01]. Mantel [Upt00]. Manufacturing [SM09]. Many [ALTY00]. Map [WF09]. Mar [SS09]. Marcin [DL04d]. Marcus [War02]. Marginal [JTFK04, PLC00, Pan02, SA03]. Margins [Mea03, RM04]. Markers [Vic08]. Market [AC05b, Ann06b]. Marketing [AR08]. Markov [Gho02]. Marlow [Ano04p]. Matching [PMO08]. Material [BGF^{+}08, FR04]. Materials [Ano02g, Ano03a, Ano04f, Ano07b, Ano07c, Ano07d, Ano08b, Ano08c, Ano09g, Ano09h, Ano09i, Ano09j, Cle01, Col01, Dev01, DL04g, DL05, DAC05, DAC06, Hai01, Hil01, Hyd01, Hyu01, Jam01d, Jam01c, Kre01, Len01d, Len01e, Len02, Mac01, Mil01, Pou02, Ree02, RACK07, RD05, Rum01, Sha01, Str02, War02]. Math [Lan06a, Len01e, Ano04p]. Mathematica [Hil06a, Sti00]. Mathematical [Alb02, Ano04p, Coh02, Dev01, DL04d, HBQ04, Lun04c, Sta08, Dob02, DL04e]. Mathematics [AC03, Ano04s, Dev06a, DL04c, Ter03, Hai01]. MATLAB [Rus08]. Matrices [Fri02b]. Matrix [Dav06]. Matter [Ano00c, Ano00d, Ano00e, Ano00f, Ano00g, Ano00h, Ano00i, Ano00j, Ano01b, Ano01c, Ano01d, Ano01e, Ano01f, Ano01g, Ano01h, Ano01i, Ano02a, Ano02c, Ano04c, Ano04d, Ano04e, Ano04j, Ano04k, Ano04l]. Matters [HV03, Kah06]. Maurer [Dob04b]. Maurice [DF07]. Maximal [Pal08]. Maximization [LG05]. Maximum [JP01, JB04, KR02, Mea03, RM04]. Maxwell [Ano04i]. MCLUST [HLW09]. Me [LF06, Bon02]. Mean [Bar09, Ber07, BS03, Bon04, DK07, Gra08, GV09, HHV^{+}02, JB04, KP07, LF06, Muk06, Nay02, SG00, WBT01, Zha07a]. Meaningful [Iri01]. Means [Tar07, ZG00]. Mease [RM04]. Measured [Cle03c]. Measures [HB03, SC09]. Measuring [Fay02, HB05a, Sab02, Str04, SGCS04]. Median [Bar09, Ber07, BH02, CLS07, FLP07, FKLL03, FG00, HJP^{+}01, Koe09, LR08, Lev02]. Medical [Dav08c, Les04, SZM^{+}07, WB04, Wie06]. Medicine [APT07, Zam08, Abb02]. Meier [Bar09, SD01, SJI05]. Men [Han04]. Message [Koe09]. Method [FMA^{+}09, Fre05, LF06, McC01, McI05, PC02, Rid09, SD04]. Methods [ACA03, AT04, Bel07, CHS04, Car00, CL02, DL04a, Dob06a, DS02, Gen03, Har06, Hyd01, Jam02, Kva00, Lee08, Mar07, Mor07, Ost02, Ost03, OH08a, OH08b, PW06, Rie06, Rob07b, Rue04, Su08, Wüt08, YA09, Len01e, Len02]. Metropolis [Hit03]. Michael [Fer04]. Mickey [Ano04b]. Microsoft [LF00, Wal02, Bor02]. Mighty [Kes00]. Milk [Pic06]. Million [Bra08a]. Mind [Cav07a]. Minded [dML09]. Minimizing [Fre05]. Minimum [BCB00, HJP^{+}01, Lev02]. Minimum-Deviation [HJP^{+}01, Lev02]. Mining [DHN^{+}05, Dev07c, HDE^{+}03, Su08]. MINITAB [AC07b, Arn06]. MINITAB^{T M} [Cle01]. MINQUE [tWGB01]. Missed [Ter03]. Missing [HL01, Spa08, vH04]. Mixed [DJ05, Gur06, MB00, Nel07, PO07, VSL01, Wul07]. Mixed-Effects [MB00]. Mixture [CLS04, LR06, MS02, VE06, WCLS04]. Mixtures [Rod04]. ML [tWGB01]. MLE [FM08]. MM [HL04]. Model [AW01, AH02, Ano04i, BCB00, CK01, Eli00, Gur06, HQ08, Jan00, JP01, JTFK04, KB07, LR06, LE00, MF07, MB00, PC02, PE05, PO07, SA00, Tim04, ZvD02]. Model-Based [HQ08]. Model/Data [AW01]. Modeling [Alb02, GC04, Hai07, Hor06, KP07, Rak06, Sim07, Sch04e]. Modelling [Nel07]. Models [AT04, CG06, CBH09, CLS04, CW04, DD06, DJ05, HV03, HT02, HL01, Hu03, JTFK04, Len01d, Mam06, OF03, PO07, Pri05a, SMVD09, SC09, Su08, Tju09, VSL01, Vic08, Wal00a, WCLS04, WGHE09, Wul07, YW09, Zho01, Cow04a, Fot02]. Modern [Abb02, Fer04, Fra00a, Fra00b, Min07, Sta08, AC08b]. Molecular [Vic08]. Moment [KC02, VE06]. Moments [De 08, LB00, Sko08, WN07]. Monotone [RT03]. Monte [FF02, KBH09]. Moore [Hu08]. Most [Kuk09]. Motivational [FR04]. Movies [ZL09]. Moving [Bel05, YW09]. Much [HK07, Len01e, LB00]. Mudgeon [Mon06]. Muller [AC04b]. Multicollinearity [FW05]. Multidimensional [Dav08d, Par04]. Multilevel [Hor06]. Multinomial [CK01, GGR05, ZvD02]. Multiple [BHA02, FW05, HR06, HL01, HLP03, KC09, Men00, Sch05e]. Multiple-Choice [BHA02]. Multiplicative [YW09]. Multiplicity [CLR13, LC09]. Multistage [CF08]. Multivariable [Lis08]. Multivariate [Ald00, AC06a, Ano00b, Dav08b, Dob06a, GR04, Gon03, Hu07, Sko08, SA03, Wil02]. Murder [Gra08]. MVUE [Muk06].

N [Bar02b, Fri15]. Named [ML09]. Nancy [Sch04f]. National [Bol03, You05, DM01]. Natural [Jam01d, ML09, Rob07a]. NCAA [AW06]. Near [HMRW01]. Near-Optimal [HMRW01]. Nearly [SA03]. Need [Mon06, Ter03]. Negative [KC02, MF07, MPR07]. Negligible [HS03]. Neil [Cha04]. Network [SS01]. Neutral [Ran01]. News [Utt02]. Next [Men09, Stu09]. Neyman [BP02, Chr05]. NFL [WB02]. NHL [BH08]. Nine [Bar02a]. NLOGIT [Hil06b]. Nolan [Dev01]. Non [Kha08]. Non-Simple-Random [Kha08]. Noncentral [LR01, WN07]. Nonclinical [Kha04]. Nonconstant [Ros03]. Nonhomogeneous [Hag09]. Nonlinear [Dav08d, SMVD09]. Nonlinearity [WWD03]. Nonparametric [CHJ08, Jam02, Min07, OH08a, OH08b, SA03, DL04f]. Nonrandomized [TTT09]. Nonregular [NVB09]. Nonresponse [HH04, KH04, Rub04a]. Nonstatisticians [UJ08]. Nontransitivity [Fri15]. Normal [CG06, CL07, CLPC08, Dav05, LJ09, Pro08, Wil08]. Normality [Haz03, Haz04, JH04]. Normally [tWGB01]. North [Mon04, Ste08]. Note [BI09, CLPC08, CK01, Dav09, Hil08b, PLC00, Pan02, Sen08, Tar00, Wes08, ZvD02]. Notes [Hil00a, Hil00b, Hil02c, Hil02d, Hil02a, Hil02b, Hil03a, Hil03b, Hil04a, Hil04c, Hil04d, Hil04e, Hil05e, Hil05f, Hil05g, Hil05h, Hil06c, Hil06d, Hil06e, Hil06f, Hil07c, Hil07d, Hil07e, Hil08c, Hil08d, Hil08e, Hil09, Ost09a, Ost09b]. Nothing [HK07, Jam01d]. Null [SBB01]. Number [Bar02a, DL00, Fre05, WE08]. Numbers [AC05b, BR03, Cle06c, Coc05b, She07, Utt02, AC05a]. Numeracy [Sow03]. Numerator [WSF02]. Numerical [Alt02, Rid09, Yal07]. Nursing [GS08].

Observation [Muk06]. Occupancy [WMAY09]. Odds [LGH09, Rob03]. Office [Cle03d]. Official [Bra07, Hol07, Nor07]. Okumoto [JP01]. Olive [Ano04b]. Olympiads [DL04d]. Olympic [ESL09, Bra07, Hol07, Nor07]. One [CLR13, Cha04, CW04, GS08, HHV^{+}02, LC09, Law03, MS06, Muk06, WBT01, ZG00]. One-Day [Cha04]. One-Dimensional [CW04]. One-Hour [GS08]. One-Sided [CLR13, LC09, ZG00]. One-Step [Law03]. Online [HQ08, HJS06, LJLC05]. Operating [BK07, Hal08]. Opinion [IdBPT00]. Opportunities [KYA^{+}07]. Optimal [HMRW01, OF03]. Optimization [Lan04, LG05, Law03]. Optimizing [Nie05]. Order [BP02, Mon06]. Ordered [MFS^{+}03]. Ordering [Nie05, SRMP08]. Origins [BM02b]. Orthogonal [Lan06b]. Orthogonality [deL06]. Other [Ano04p, CF08, Cle03d, Pic06, TGM08]. Our [She07]. Outcome [MFS^{+}03]. Outcomes [BLS00, CF08, LBTM03]. Outlier [Fin06]. Output [Pra08b]. Overlap [Bar02b, SG01]. Overtime [BH08]. Owens [Fri15].

P [Ano04p, DL04c, Lev02, KP07]. Packages [BLS00, Cal03, DHN^{+}05, GC04, HDE^{+}03, HKS06, HLW09, HL01, Kel04, Ost02, Ost03, OH08a, OH08b]. Pages [LP00]. Paired [OSC00]. Paired-Data [OSC00]. Paleontology [MS01]. Paper [Ann05b, Igl07]. Paradox [CBH09, KS05, Sen08, Shu06, Sin00, SS01, Bel05, CBH09, PP09, RT03, SR04, SR05]. Paradoxes [BS00b, Les04, WB04]. Parameter [Bar02a, HM01, LaM08, MV07, SGM07]. Parameters [Iri01]. Parametric [CHJ08, Hil02e, OH08a, OH08b, Ros00a, SA03]. Pareto [Wil06a]. Part [Hil04b, Hil05b, Hil03c, Ost03]. Partial [CW04]. Passing [Lan09]. Passion [Lan09]. Patterns [Rod04]. Paul [Dob04b, Lun04a]. PDF [HT06]. PDFs [Nad09]. Peace [You05]. Peak [KS08]. Pearson [BP02, Chr05, Cle05a]. Pedagogy [BM02b]. Peirce [BK07, LR06, RA04]. Penalized [Mea03, PW06, RM04]. Penn [HR05]. Penrose [Hu08]. Perceive [TS03].

Percent [BA06]. Percentage [MSL^{+}08]. Percentage-Based [MSL^{+}08]. Percentile [CL07]. Perfect [CLR01]. Performance [Ann06a, Tob04]. Period [BH08]. Permutation [FF02]. Personal [BG08]. Persons [Moe05a, Moe05b]. Perspective [Ano04i, Jam01c]. Perturbing [Wil08]. Pervasive [HH01]. Petersburg [Bel05, Rum07, SR04, SR05]. Petition [WE08]. Pharmaceutical [Bra08b]. Phillip [AC04a]. Philosophical [Cle06b]. Picture [FS01]. Picturing [Cha08]. Pieces [Dob03b]. Pilot [MB06]. Place [Jon02]. Planning [IJW06]. Platt [Lev02]. Play [FDC09]. Player [Shu06]. Players [SG00]. Playfair [SW05]. Playing [Ann06a, Coc05a]. Playoff [AW06]. Plot [AH02, CC06, Jan00, Pad08, SMC^{+}07]. Plots [BM03, DR07, GB03, MB00, SRMP08]. Plotting [Lan06b]. PLUS [Cow03, Har06, Dia05, Fox02, Man03, Pri05b, Lun02b]. Plutarch [Mis06]. Pocket [Gil08]. Point [DC01, Kes00, Sin00]. Points [FRD07, Spi07]. Poisson [Bar02a, LJ09]. poLCA [HLW09]. Political [Gel08]. Polls [CF08]. Polychotomous [WB02]. Polynomials [AC00b]. Pool [DM01]. Pool-Adjacent-Violators [DM01]. Pooling [Jan00]. Population [LR08, MG04]. Populations [KT02, RZR05, Str04, SGCS04, WS01]. Positively [Fri15, LSO01, ZG00]. Possible [WY04]. Post [Zha00]. Post-Stratification [Zha00]. Posttest [YT01]. Potential [AW06, HLP03]. Power [Bro01, CRM01, CL02, De 06, HH01, MSL^{+}08, vDK02]. Powerful [Kuk09]. Practical [Bra08b, Dev07c, Gen03, Len01f, Lis08, SB09, LF00]. Practice [BRRB01, Bro01, CRM01, CK01, DLM01, Dob03b, Dob04b, Dob04a, GPD02, HMRW01, HL01, Len01f, MS01, Nas06, Ran01, SG01, WSF02]. Practices [Fin06]. Pre [CF08]. Pre-Election [CF08]. Preach [GPD02]. Precise [SBB01]. Predict [Len01e]. Predicting [CF08]. Prediction [Ald00, LS08, Nay02, Tar00, Vic08, YA09]. Predictive [AT04]. Predictor [Eli00, GP09]. Predictors [LBTM03]. Preliminary [KR02]. Preparation [Wie08]. Preparing [Moo05]. Preschool [HGA01]. Presence [AB03, BH08, FKLL03, KP07]. Presentation [HQ08]. Presidential [CF08, WJM^{+}03]. Pretest [YT01]. Pretest-Posttest [YT01]. Previously [WWD03]. Primer [For05]. Principle [BCB00]. Principles [Str04, SGCS04, Sch00]. Prior [SS02]. Priors [De 06, TGM08, WSF02]. PRISM [Zim00]. Proactive [Hah02]. Probabilistic [Ros07, Wüt08]. Probabilities [AC00b, She07]. Probability [Alb03a, Ann05a, BG08, BW08, Car08, Cle03d, Cle06b, Cle08a, Coc05a, Coh02, Cow06, Dev06b, Dob04b, Dob04a, FR04, Fot02, Gho02, GEL01, Hai07, Har08, Kie06, Kir00, LS08, Lun04c, Mil03, Rob03, Rus08, SD01, Sch06b, Ste01, Utt03b, DL04b, Cle03d, Dev02, Dob03a, Nel01]. Probit [CG06]. Problem [Cle08b, DO04, HH04, Meh09, Sny05, SS09, WMAY09, Kuo00, MS01]. Problems [Dob04b, DL04d, Dob04a, HV03, NVB09, WSF02, Dob03a]. PROC [Ann05c]. Procedure [And09, CRM01, KH04]. Procedures [Hil02e].

Process [AC08a, Law03, Sin00]. Processes [Dev06b, Hai07, Lun02a, Lun04c, Lun04a, Lun04b, Rus08]. Products [Hil03c]. Profession [Lan09]. Professional [Gil00, Hil07a]. Professor [Mon06]. Profitable [Len01e]. Program [Zim07]. Program-Driven [Zim07]. Programmer [Gur08]. Programming [Cha04, GEL01, Spe08, Zim07]. Programs [Moo01]. Project [Dip04, HR00]. Projects [Ano04s, RT01]. Promoting [Bol03]. Proof [DS08, Hwa09, Ros07, Gla04]. Proofs [Gla04]. Propensity [PMO08]. Properties [DR01, JS03, vDK02]. Property [Fri15, LSO01, Wil08]. Proportion [Afs08, CRM01, CL02, Fre07b, MG04]. Proportions [AC00a, IdBPT00, KT02, WS01]. Proposal [Tju09]. Prospective [RSH01]. Prosperity [Sny05]. Protect [KKO^{+}06]. Psychiatry [Ham00]. Psychologists [Row02]. Psychology [Jam01c, Mru07]. Public [AC05b, Gel08]. Published [Igl07]. Putting [Dob03b]. Puzzlers [Cle03d].

Q [Ano04p]. Quality [AC07a, AC07b, Hyd01, Zam08]. Quantify [SR05]. Quantile [Ros00a]. Quantitative [Cle06a, PH02, Cle01]. Quantum [Tur07]. Quarter [GP09]. Quarterbacks [WB02]. Questions [Mon06, TTT09]. Quick [AC08a, MY05]. Quinn [Gla04].

R [Cow03, Fer04, Gel04, HHV^{+}02, Lev02, Mon06, Rue04, Fox02, HJM08, Har06, HBQ04, Mäc07, Pri05a, Pri05b, Rob08, WC07, Bat04, Lun07b]. R. [Box05, Lud05]. R/S [Har06]. R/S-PLUS [Har06]. Raindrop [BM03]. Random [BRRB01, Bel01, CK01, DL00, FR04, GL00, Grö09, Hai07, HT06, JB04, Jon02, Kah06, Kha08, MB00, MTA08, Nel01, Pal08, Rus08, SHS04, Ter05, Tur07, Wil08, ZvD02]. Randomization [Dav08a, Gad01, Moe05a, Moe05b]. Randomized [Dip04]. Range [Bro01, Fay02]. Rank [Ald00, Kuk09, Ros03]. Ranked [McI05]. Ranking [Fre05, Mea03, RM04, WB02]. Rankings [GDL08]. Rao [Kag01, Nay02]. Rate [ZL09]. Rater [LR06]. Raters [ZL09]. Rates [CLS07, Gra08, HS03, SD04, VD05]. Rating [ZL09]. Ratings [HQ08]. Ratio [MV07, PE05]. Ratios [LGH09]. Ratliff [Ano04s]. Real [DLM01, JS03, Kha08]. Real-Life [Kha08]. Realism [Ste07]. Really [Coc05a, FDC09, Gla04]. Reasoning [AC06b]. Reassessing [CLS07]. Receiver [BK07, Hal08]. Record [SS05]. Recurrence [WN07]. Recursive [Sko08]. Reduced [Ald00]. Reduced-Rank [Ald00]. Reducing [GR04, KH04]. Reduction [SM09]. Reformulation [Lee09]. Regarding [Ter05]. Regions [DC01]. Registration [SD04]. Regress [SG00]. Regression [Ald00, Ano04b, CBH09, CW04, Dev07a, DR01, Eli00, Fay02, FLP07, FW05, Gra08, Grö07, Grö09, HL01, Iri01, Jam03, Ker07, KR02, KP07, LM03, LE00, Men00, Men04a, Men04b, Mon06, MFS^{+}03, Nel07, OF03, PLC00, PC02, Rak06, RT03, RA04, Sim07, Tju09, WWD03, WB02, XN03, Zho01, Fox02, AC04b, Cow03]. Regrouping [RT03]. Rejoinder [Ano03i, BK09a, OH08b]. Related [Dav05, Gho02, Utt02, XN03]. Relating [JW04]. Relation [CBH09, WN07]. Relationship [IBP05, SGK00]. Relationships [LM08]. Relative [Grö07, HB05a]. Relatives [ML09]. Reliability [JP01, SS01, WJM^{+}03, Yal07]. Remain [SR05]. Reminder [Paw00]. Reminiscence [Box05]. REML [Gur06, tWGB01]. Replacement [KH04]. Replicating [Ros01]. Replications [BM02a]. Reply [Ano04q, HHV^{+}02, SMH^{+}04, JH04, RM04, WCLS04]. Report [AHL04, Alb05a, Alb05b, Dob06b, SS04, Sto01, Wes07]. Reporting [Utt02]. Reports [Alb03b, Sto02, SA02]. Representation [YW09]. Representing [Bor02]. Reproducing [PW06]. Republication [SW05]. Resampling [Har06, Gen03]. Research [Ann06b, DL04a, Gel04, HS04b, Kha04, KC09, Los04a, Rub04b, SHS04, SZM^{+}07, Woo08, Str02]. Residual [Ste07]. Residue [Hag09]. Response [DC01, Dip04, HS04a, PLC00, Rak06, SHS04, SR05, TTT09, Fay02]. Responses [Ran01]. Restricted [Tar00]. Result [AC00a]. Results [DS08]. Retention [BUK^{+}08, Kva00]. Rethinking [Ann05b]. Retired [FG00]. rev [McG05]. Revealing [Sch00]. Revenge [Lun07a]. Reversals [Gil00]. Review [Abb02, Alt02, AC02, ACA03, AC04b, ACL04, AC04a, AC05a, AC06a, AC08a, AC08b, Bat04, Ber00a, Ber00b, Ber00c, Bon02, Can01, Car00, Cha04, Che00, Cle00a, Cle00b, Cle03b, Cle03a, Cle03c, Cle04, Cob00, Coh02, Cow03, Cow04a, Cow04b, Cow04c, DHN^{+}05, Dev02, Dob02, Dob03b, Dob03a, Dob04b, DL04e, DL04c, DL04d, DL04a, DL04f, DL04b, Dun04, Fer04, Fot02, Fra00a, Fra00b, Gen03, Gla04, Gon03, Gup02, Hai03, Ham00, Har00, HH06, HDE^{+}03, HKS06, HLW09, Hil03c, Hil04b, Hil05b, Hil05d, Hil05a, Hil05c, Hil06b, Hu03, Jam00, Jam01a, Jam01b, Jam02, Jam03, Jan00, Jer03, JTFK04, Kan00, Kel04, Kir00, Lac00, LJLC05, Led03, Len00, Len01a, Len01b, Len01c, LF00, Los04b, Lun02b, Lun02a, Lun04c, Lun04a, Lun07b, Man03]. Review [Mil04b, Mil04a, Mil04c, Nel01, Owe01, Pag00, Pie01, Ree00, RH02, Rob03, Ros00b, Row02, Rue04, Sch00, Sch04e, Sch04f, Ste01, Sti00, Tan02, Ter03, Utt02, Ver01, Wal02, Wal00b, Wil02, Wul03, AC03, Min07]. Reviews [Alt02, Ano02g, Ano03a, Ano04f, Ano07b, Ano07c, Ano07d, Ano08b, Ano08c, Ano09g, Ano09h, Ano09i, Ano09j, Cle01, Cle03d, Col01, Dev01, DL04g, DL05, DAC05, DAC06, Hai01, Hil01, Hil02c, Hil02d, Hil04f, Hil07g, HL01, Hyd01, Hyu01, Jam01d, Jam01c, Kre01, Len01d, Len01e, Len02, Len04a, Mac01, Mil01, Mil03, Pou02, Ree02, RACK07, RD05, Rum01, Sha01, Str02, War02, Zim00]. Revising [Wil06a]. Revisited [KC02, Meh09]. Revisiting [WMAY09]. Reviving [MF07]. Rhythm [EKG02]. Richard [Mil04c]. Risk [Gue06, Hai03]. Risk-Benefit [Hai03]. Robert [Dob04b, Fer04, Rue04]. Robin [Ano04p]. Robust [Fre06]. Rock [BHM01]. Role [Boy08, HF04, Wal05, WSF02]. Roll [BHM01]. Rolled [Dun03]. ROM [Mil03, Sch04f]. Ross [DL04e]. Rounding [HLP03, YHZ08]. Rule [Bet00, BP02, She07, WY04]. Rules [Cle03a, Sch06b]. Runs [Ann06a]. Ruth [Ano04b]. Rykken [Ano04s].

S [Cow03, Dia05, Fox02, Man03, Pri05b, Lun02b]. S-PLUS [Cow03, Har06, Dia05, Fox02, Man03, Pri05b, Lun02b]. Saba [SS09]. Saddlepoint [Kuo00]. Sample [Afs08, BM01, Bro01, CRM01, CL02, DK07, FMA^{+}09, Fre05, GJLW05, HH04, IBP05, Kha08, LaM08, Len01f, MS06, MY05, PB03, Rub04a, SRC07, TTT09, Zha07a]. Sampler [CLR01]. Samples [HHV^{+}02, Kha08, MFS^{+}03, Wal00a, WBT01]. Sampling [AL07, Bel01, Bel07, Bul05, HS03, LR01, Loh04, McI05, Mis06, PHSN08, Rya09]. Samuel [Dob04b, DM06]. Sandwich [Fre06]. SARS [YCF06]. SAS [AC04b, ACL04, Ano00b, DL04a, Gup02, Ann05c, Bra08b, Cot06, Dix07, Dob06b, Gur08, Hal08, Pra08b, Pri05b, Spe08, Wie06, Wie08, Wul07, Zim07, Cha04, Cow04a]. Satisfying [LSE00]. Saw [CL02]. Saw-Toothed [CL02]. Scaling [Par04]. Scatter [Upt01]. Schaeffer [Mil04c]. Schenker [Bar02b]. School [Bol03, Les04, Ter03, Tob04, WB04]. Schools [LH02]. Schwertman [Fri15]. Science [Ann05a, Arn07, BGNP01, Cle05a, Cow06, Dev06a, Gel08, Rie06, Wüt08, BGNP01]. Sciences [Dav08c, Fer04, Lee08, Rob07a, Gon03, Kre01, Hu03, RH02, Tan02]. Scientific [Cle05a, HH06]. Scientists [Car00, Car08, Kie06, Moo07, Wei07, Ste01]. Score [Fre07a, MV07, MPR07, PMO08, VM07]. Scott [Ano04i]. Screening [Law03]. Seasonal [YW09]. Seber [Mac01]. Second [Dev07a, Jam01c]. Secondary [Bol03]. Section [AG04, Utt03a, Hil00a, Hil00b, Hil02c, Hil02d, Hil02a, Hil02b, Hil03a, Hil03b, Hil04a, Hil04c, Hil04d, Hil04e, Hil05e, Hil05f, Hil05g, Hil05h, Hil06c, Hil06d, Hil06e, Hil06f, Hil07c, Hil07d, Hil07e, Hil08c, Hil08d, Hil08e, Hil08b, Hil09, Ost09a, Ost09b]. Sectioned [CC06]. Seeing [AC07c, Cav07a]. Seek [BHA02]. Selected [Dob02]. Selecting [Gur06, Par04]. Selection [Bar07, BCB00, HCW07, JTFK04, MR09, SA00, Sen08]. Selective [McI05]. Semielasticity [HF04]. Seminar [GS08]. Semiparametric [BI09]. Sense [Dev07c]. Sensitive [TTT09]. Sensitivity [SGM07, SS02]. Separate [BM02b, GC04]. Separation [XN03]. Sequence [GGR05, HR00, Jad00]. Sequences [BHA02]. Sequencing [CR01]. Sequential [LS01]. Series [Pou02]. Service [BHM01, RT01]. Service-Learning [RT01]. Services [Gal03]. Sets [McI05]. Setting [DR01]. Sex [BHM01]. Shading [Jac08]. Shall [Gra08]. Shall-Issue [Gra08]. Shaped [LBTM03]. Sharpening [Cot06]. Sherlock [Ano04p]. Shewhart [SF00]. Shieh [Fri02a]. Shot [RHCR06]. Should [FG00, Sin00, Utt03b]. Shrunk [Ald00]. Sided [CLR13, KB07, LC09, ZG00, vDK02]. Sigma [AC07b]. SigmaPlot [Hil03c, Hil05a]. SigmaStat [Hil03c, Hil05i]. Sign [FKLL03, Ran01]. Signatures [WE08]. Signed [Ros03]. Significance [Bar02b, SG01]. Significant [GS06, Rod04]. Significantly [Chr03]. Simple [AC00a, And09, Bel01, CLS04, Eli00, Few09, Fre07b, Fri02a, GS02, Hwa09, Kha08, Kos05, Pal08, PO07, Ros07, SGK00, Shi01, Sol00, WCLS04]. Simpson [CBH09, PP09, RT03]. Simulating [MS02]. Simulation [KBH09, KP07, Shu01, dML09]. Simulation-Based [KBH09, dML09]. Simulations [AW01]. Simultaneous [PE05]. Single [Bon02, CL02]. Sinh [New01]. Sir [DF07]. Sitter [Gel04]. Six [AC07b, LJLC05, Men04a, Men04b]. Size [Afs08, Bro01, CRM01, CL02, FMA^{+}09, HHV^{+}02, IBP05, Len01f, PB03, SRC07, WBT01]. Sizes [Tim04]. Skating [WY04]. Skew [CLPC08]. Skew-Normal [CLPC08]. Skewed [ZG00]. Skills [Cot06]. Sleuth [Led03]. Slow [TKM04]. Sluggers [Ann06a]. Small [FMA^{+}09, MFS^{+}03, Par04]. Small-Sample [FMA^{+}09]. Smoothed [SS08]. Smoothing [Sol00]. So-Called [Fre06]. Social [Gon03, Lee08, Rak06, RH02]. Sociology [Gel08]. Softly [Ald00]. Software [Alt02, Cal03, CL02, HDE^{+}03, HKS06, Hil02c, Hil02d, Hil04f, Hil07g, HL01, JP01, Kel04, Ost02, Ost03, OH08a, OH08b, Zim00, AC04b, Ano00b, Col01]. Solution [CLS04, WCLS04]. Solutions [Bel07, CL02, DL04d]. Solving [Cle08b]. Some [DS08, Gel04, HS04b, Len01f, LJ09, SHS04, Ter05, VM03, YA09]. Sometimes [Jan00]. Sources [HB05a]. Space [MV07, PO07]. Spaced [GL00]. Spaces [LaM08]. Spatial [RHCR06]. SPC [BA02]. Special [AG04, Alt02, Dob04b, Dob04a, Utt03a]. Specified [MS02]. Speed [Dev01]. Spencer [Cha04]. Spline [Sol00, TH00]. Splines [BI09, PW06]. Split [AH02, Jan00, Pad08]. Split-Plot [AH02, Jan00]. Splitting [GP09]. SPlus [Cow04b]. Sports [Coc05a, Fel08]. Spreadsheets [Nas06]. SPSS [Dav07, Dav08b, Han05, Hil03c, Hil04b, Hil05b, MJ04, vH04, Str02]. SQL [Ann05c]. Square [Abb02, DK07, WN07]. Squared [Fri02a, Mir01, Nay02, Shi01]. Squares [Lee09, Tar00, deL06]. St. [Bel05, Rum07, SR04, SR05]. Stable [LR01]. Standard [Fre06, Gad01, GC04, LE00, vDK02]. Standardized [Men04a, Men04b]. Stanley [DM06]. Stat [Dev01]. Stata [Hor06, Hil05c, McG05, Lac00]. StatConcepts [Pag00]. State [ALTY00, CF08, FDS05, HR05, Mon04, PO07, SS04]. State-Level [CF08]. State-Space [PO07]. Stationary [DC01]. Statistic [BSTM09, MPR07, Paw00, Tar00]. STATISTICA [Hil07f]. Statistical [Abb02, Alt02, AC06b, AC08a, BRRB01, BB04, BUK^{+}08, Bol03, Bro01, BGNP01, Cal03, CHS04, Car00, CRM01, CK01, Cle05a, Cle08a, Cle08b, Cra05, DLM01, Dia05, DL04a, DL04f, DS02, Fel08, Gal03, Gel04, Gen04, GB06, Gup02, HMRW01, HS04b, HM01, Hil02c, Hil02d, Hil02e, Hil04f, Hil07g, HL01, HB03, Hun05, IJW06, Jam02, JW04, Kha04, KBH09, KC09, Lac00, Led03, Len01f, Len02, Les04, Lis08, Loh04, Los04a, Lun02b, Lun07b, Mam06, Man03, Mar07, MSL^{+}08, MS01, McC01, Mon04, Mor07, Nas06, Ost02, Ost03, OH08a, OH08b, Pag00, Pri05b, Ran01, Rid09, Rie06, Rub04b, Sch00, SG01, SMC^{+}07, SS09, SW05, Str04, SGCS04, Utt02, Utt03a, WB04, Wal05, Wie06, Wie07, Wil06b, WSF02, Wüt08, YCF06, Zam08, Zim00, CG07, AC02, ACA03]. Statistical [Cle03a, Cow04b, Hyd01, Jam00, Kre01, Rue04]. Statistical-Power-Based [MSL^{+}08]. Statistically [GS06]. Statistician [Bar02b, Fri02a, Fri15, Hah02, HHV^{+}02, JH04, Les04, Lev02, RM04, WCLS04, ZvD02, Alb03c, AHL04, Alb05a, Alb05b]. Statisticians [RSH01, Sow03, TS03, VM03]. Statistics [AG04, AL07, AR08, ACL04, AC04a, AC04c, AC05b, AC07c, Ann05a, Ann06a, Ann06b, Ano00b, Ano04b, Ano04q, Arn07, Bat04, BG08, Ber06, BLS00, BM02b, Bil04, BHM01, Bra07, Bra08b, Bro08, BK09a, BK09b, CF01, Car08, Cav07b, Cav07a, CR01, Cha08, CF07, Chi08, CHJ08, Cle04, Cle05b, Cle06a, Cle06c, Cle08a, Cle08b, Coc05b, Col04, Cow06, Cra05, DS08, Dav06, Dav08c, Dev01, Dev06a, Dev06b, Dev06c, Dev07a, Dev07b, Dev08, Dob02, Dob03b, Dob04a, Edw07, FR04, For05, Fra00a, Fra00b, FDS05, GS08, Gal03, Gel05, Gen04, GR09, Gon03, Har06, HR05, Har04, HF04, HS09, HGA01, HR00, Hol07, HBQ04, Joh09, Kah06, Kha08, Kie06, Kir00, Kva00, Lan06a, Lan04, LJLC05, Lea04, Len04b, LP00, LF00, LH02, LG00, Mäc07]. Statistics [MG09, Mas04, Mea04, Mil03, Mil04b, Mil04a, Moo01, Moo05, Moo07, Mru07, NL09, Nor07, Owe01, Ree02, RH02, Ric08, RT01, Ros00b, Rum06, SS04, Sin08, Skl07, SB09, Sta08, Ste01, SMH^{+}04, Sti08, SZM^{+}07, SV03, Tal07, TS03, Tra07, Utt03b, UJ08, VM03, Ver01, Wal02, Wan02, Wei07, You05, deL06, Can01, Coh02, Dob04b, Mil01, Mil03, Wal00b, Mil01, Min07, Abb02, AC08b, Cle00a, Cle00b, Cle03b, Ham00, Jam01c, Len02, Mil04c, Pie01, Rob03, Ros00b, Row02, Rum01, Sha01, Sti00]. Stats [McD04]. Stature [Han04, WWD03]. StatXact [Cal03]. Steadfast [SR05]. Stem [SMC^{+}07]. Stem-and-Leaf [SMC^{+}07]. Step [ACL04, CRM01, Law03, Rid09, ACL04]. Step-by-Step [ACL04, ACL04]. Stephen [War02]. Stochastic [Dev06b, Lun04c, Lun04b, RZR05, Sch04e, Lun02a, Lun04a]. Stock [AC05b, Bel05]. Stocks [SR04]. Stopping [Bet00, BP02]. Strategic [SMC^{+}07]. Stratification [Zha00]. Streakiness [AW01]. Strikeouts [Ann06a]. Strizaker [DL04b]. Strong [Bro07]. Structural [JTFK04]. Structure [HCW07, Zim00]. Student [Alb03a, ACL04, BUK^{+}08, Kva00, RT01, HJM08]. Students [AR08, Bol03, Bro08, GS08, Gel08, HR05, KYA^{+}07, Mon06, Moo05]. Studies [PHSN08, Tim04]. Studio [Pra08b]. Study [BA02, Jan00, Lis08, MB06, Shu01, YT01]. Style [GGR05]. Styles [PH02]. Subhabrata [DL04f]. Success [JW04]. Successes [AC00a]. Successful [BM02a]. Sufficiency [KS05, LaM08]. Sufficient [CBH09]. Suggestions [LJ09]. Sullivan [Sch04f]. Sum [Fri02a, Hag09, Shi01, Tar00]. Summarize [SMC^{+}07]. Sums [Ter05]. Supplement [CF01, Sch04f]. Support [Ter05]. Suppression [FW05, Lyn03]. Surface [DC01]. Surprising [FM08, SS05]. Survey [BHM01, Loh04]. Surveys [Dem06, HH04, Kha08, Rub04a, TTT09]. Survival [Bar09, Fre08, Kel04, LSE00, McG05]. Swan [Lun07a, WH07]. Swans [Bro07, Tal07]. Swing [IdBPT00]. Symmetric [Pal08]. Symmetrized [BA06]. Synthesis [Zha00]. Systat [Hil05d, Hil08f]. System [HHvdWZ06, Pra08b, Rub04a]. Systems [AW06, Col01, Hil08a].

T [Gla04, MS06]. Table [Len02, SMC^{+}07]. Tables [Dav05, GPD02, Kos05, Mir01, Nie05]. Tabulation [MSL^{+}08]. Tail [LB00]. Tails [KS08]. Tale [AH02]. Tales [Ano04p]. Teach [CR01, Moo05, Bon02]. Teach/Me [Bon02]. Teacher [BHM01, CF01, CR01, HJP^{+}01, HM01, HGA01, IM01, Jan00, RT01, Zho01]. Teachers [Ano04p, FDS05]. Teaching [AG04, AR08, Ano02g, Ano03a, Ano04f, Ano07b, Ano07c, Ano07d, Ano08b, Ano08c, Ano09g, Ano09h, Ano09i, Ano09j, BG08, BH08, Cle01, Cle05b, Cle06a, Col01, Dev01, DL04g, DL05, DAC05, DAC06, Eas04, Gal03, Gel05, Gel08, Hai01, HR05, Hil01, Hun05, Hyd01, Hyu01, IM01, Jam01d, Jam01c, Kre01, KB07, Lea04, Lee07, Len01d, Len01e, Len02, Len04b, LP00, LJ09, Mac01, Mil01, Mil04a, MJ04, Mon04, Pou02, Ree02, RACK07, RD05, Rum01, Sha01, SM09, Str02, Str04, SGCS04, SV03, UJ08, War02, Cle04]. Teams [Mea03, RM04]. Tech [SR04, BM05]. Technique [DD06]. Techniques [Cle08a, KC09]. Tell [LF06]. Term [HR00]. Term-Long [HR00]. Ternary [Upt01]. Terry [Dev01]. Test [BM01, BRRB01, BLS00, Cal03, FKLL03, Fre07a, FG00, GJLW05, MS06, MY05, MPR07, Ran01, RZR05, Ros03, Sho03, Upt00, VM07]. Testing [BRRB01, BW08, BH08, CLR13, CRM01, Chr05, HB03, KT02, LC09, Ros00a, SBB01]. Tests [BHA02, CLR13, Chr03, Dav08a, FF02, FS01, Jan00, Kuk09, LC09, MV07, Sab02, Vic08]. Text [DHN^{+}05]. Text-Mining [DHN^{+}05]. Textbook [Dav08c, Sch04f]. Their [BHA02, De 06, JW04, Lun04a, Lun04b, ML09, PE05, SGK00, Sow03]. Them [AC04a, AC04c, ZL09]. Themes [Dob06a]. Theorem [Gho02, Hwa09, Jon02, Ros07, Bel01]. Theory [Coh02, Hun05, Lan04, LG00, Lun04c, Zim07]. Think [Cle00a]. Thinking [AC06b, Bol03, Dob06a, Fel08, Ros00b]. Third [GP09]. Thomas [Ano04s]. Thoughts [BM02b, SHS04]. Three [HKS06, HLW09, Hu08, KB07]. Three-Condition [Hu08]. Three-Sided [KB07]. Thumb [Cle03a]. Tiered [WB02]. Ties [FKLL03, Ran01]. Tim [Col01]. Time [Ann06a, GC04, Pou02, SJI05, Zho01]. Time-Change [Zho01]. Time-Varying [SJI05]. Titan [Cle05a]. Titu [DL04d]. Toast [Sti08]. Together [Dob03b]. Tool [DC01, Haz03, Haz04, JH04]. Toolbox [HBQ04]. Tools [IM01, LG05]. Toothed [CL02]. Topics [CR01, Dob02]. Toroidal [Upt01]. Tour [Pag00]. Toxicologists [Pie01]. Toys [HS03]. Traditional [DS02]. Training [BM05, FDS05, GS08, HR05, RSH01]. Transformation [FLP07, New01]. Transformations [SA00, Tar07]. Transitive [Fri15, LSO01]. Transmuting [Han04]. Treatment [BA06, YT01]. Trees [Bor02]. Trends [Len01e]. Trial [YT01]. Trials [CRM01, De 06]. Tricks [Cle05b]. Trout [Pic06]. Truncation [Man07]. Trust [Ros00b]. Tukey [HR06]. Tunable [JB04]. Turning [GPD02, Spi07]. Tutorial [HL04]. Tweedie [HHV^{+}02]. Twins [BSTM09]. Two [Abb02, AC00a, BM01, CLR13, CRM01, DS08, DHN^{+}05, GJLW05, HHV^{+}02, IdBPT00, KT02, LC09, Les04, LBTM03, MS06, MY05, PE05, RZR05, Shu06, WB04, WBT01, vDK02]. Two-Player [Shu06]. Two-Sample [BM01, GJLW05, MS06, MY05]. Two-Sided [vDK02]. Two-Step [CRM01]. Type [BSTM09, Nay02, dM07, AB03, VD05].

U.S. [CHS04]. UCLA [Lan04]. Unbalanced [GL00]. Unbiased [Eli00, McI05]. Uncertainty [Ann05a, Jac08]. Uncertified [WJM^{+}03]. Undergraduate [Moo01, SS04, Wil06b]. Underlying [SF00]. Understand [Bel05, WGHE09]. Understanding [Abb02, Bro08, DJ05, Han05, Mil04c, RT01, Sch06b, Skl07, Zho01, dML09, Lun02b, Mil04b]. Undiscovered [WWD03]. Unequally [GL00]. Uniform [Rod04]. Uniformly [NVB09]. Unifying [ML09]. Unique [Wil08]. Univariate [Dav08b, LM08]. University [Gel05, GDL08, FDS05, HR05, Mon04]. Unknown [MS06]. Unlocking [Cra08]. untitled [Ber00a, Ber00b, Ber00c, Che00, Cob00, Jam01a, Jam01b, Kan00, Len00, Len01b, Len01c, Ree00]. Updated [Dav07]. Upper [GP09]. Urgent [Cle06a]. Use [BRRB01, BB04, CHS04, De 06, FKLL03, FG00, HBQ04, LP00, Pan02, Sch04f, SZM^{+}07, Str02]. Used [KC09]. User [Bon02, CG06]. User-Friendly [CG06]. Users [Spe08]. Uses [Cle06b]. Using [AW01, Ald00, AC00b, AC04b, ACL04, Ann05c, Bor02, Bra08b, Bro01, BCB00, CLR13, CL02, CF08, Cle04, Cle08b, DD06, DT00, GC04, Han05, HB05a, HR00, Hor06, Hwa09, Lac00, LG05, LC09, Lea04, LF00, LE00, Lun02b, Lun07b, McG05, McI05, Ost02, Ost03, OH08a, OH08b, Pra08b, RT01, Shu06, SF00, SM09, Str04, SGCS04, Tim04, VE06, Wie06, Wie08, YHZ08, Cle01, Cow04b, Gup02, Mäc07, Man03, Rus08]. Utah [CHS04]. Utility [Hil07a, KKO^{+}06].

v [CHS04]. Valen [Dun04]. Valid [WE08]. Valuation [Bel05]. Value [BH02, vH04]. Values [KP07, MTA08, SBB01]. Vardeman [War02]. Variable [Grö09, HT06, Wil08]. Variables [BRRB01, HL01, JB04, MTA08, Nel01, Rak06, Ter05, Upt01]. Variance [Ano04b, DK07, Grö07, HB05a, Jan00, JB04, MS06, Nel07, Zha07a]. Variances [Sho03]. Variate [Jon02]. Variates [Hag09]. Variation [HB05a, SM09]. Variations [Pad08]. Varying [SJI05]. vec [Hu07]. Vector [Wan02, YW09]. Versions [Dav07]. Versus [CL02, HV03, HB03, Moe05a, Moe05b, WSF02, Grö09, KR02, MR09, MSL^{+}08]. Very [Cle06b]. via [FLP07, HMRW01]. Victor [Ano04p]. Victory [Mea03, RM04, deL06]. Viewing [Zim00]. Viewpoints [MR09]. Views [FW05]. Violating [SF00]. Violators [DM01]. Virginia [Ano04b, BM05]. Virtual [SM09]. Visual [AC07c, Pag00, Pic06, dML09]. Visual-Minded [dML09]. Visualization [HJS06]. Visualizing [Bra08a, FK09, HJP^{+}01, KS08, Lev02]. vol [Dob02]. Volume [AHL04, Alb05a, Alb05b, Ano00k, Ano00o, Ano01k, Ano04r, Ano09c, Sab02, Ano08e]. Voodoo [Cle03d]. Vote [MSL^{+}08].

W [AC04a, Lev02, Mil01, Mil04c]. W. [Sne08]. Wainer [Les04]. Wald [MV07, Paw00]. Waldo [FK09]. Walk [Pal08]. Walker [DL04a]. Walks [FR04]. Wall [HHV^{+}02]. War [You05]. Warrack [Sha01]. Warren [FS06]. Watkins [Mil04c]. Way [BM03, Kuk09]. Ways [Mir01]. Web [LG05, SV03]. Web-Based [LG05, SV03]. Weighing [Rob03]. Weighted [DM01, Kha08, Lee09, SD01]. Whence [BHA02]. Where [FK09, HV03, Sin00]. Which [BRRB01, JW04]. Whiskers [DR07]. White [Lun07a]. Whitney [BLS00, Ran01]. Who [Tra07, ZL09]. Width [CRM01]. Wilcoxon [BLS00, Ran01]. Wild [Mac01]. Wilks [DM06]. William [Ano04p, Col01, Rue04, SW05]. Wilson [Ano04p, Koe09]. Win [Alb02]. WinBUGS [Cow04c]. Windows [Dav07, Str02, Han05, Har00]. Wisdom [Sow03]. Wishart [GS02]. Within [Moe05a, Moe05b]. Without [DG07, HB05a, Hu07, Lan06a]. Witmer [Mil01, Mil01]. Wittkowski [HHV^{+}02]. Women [Han04]. Woolson [Rue04]. Workbench [Los04b]. Working [Cle06c, HCW07]. Working-Correlation-Structure [HCW07]. WorkPlace [HH06]. Workshop [Ric08, Ree02]. World [Cha08, DLM01, DL04d, Kah06, Mil04c]. Worth [FS01]. Wrestler [Gel06]. Writing [AC05a, AC06a, Dob06b, Mon06, Ano04s]. Wrong [VM07].

X [Hyu01]. XploRe [Hil01]. XPro [Hil02e].

Yates [Upt00]. Year [BHM01]. Years [Men09, Stu09]. Youden [BK07]. Young [VM03].

Z [AC03, ZvD02]. Zero [Dav09, WSF02, Jam01d]. Zero-Numerator [WSF02]. Zeros [Ran01]. Zuming [DL04d].

References

Austin:2003:TEI

[AB03] Peter C. Austin and Lawrence J. Brunner. Type I error inflation in the presence of a ceiling effect. The American Statistician, 57 (2):97–??, May 2003. CODEN ASTAAJ. ISSN 0003-1305 (print),

            1537-2731 (electronic). URL http://oberon.ingentaselect.
            com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/00031305/v57n2/
            s8/p97.

Abbott:2002:BRB

[Abb02] Robert D. Abbott. Book review: Statistics at Square Two: Understanding Modern Statistical Applications in Medicine. The American Statistician, 56(4):331–??, November 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

            tronic). URL http://oberon.ingentaselect.com/cgi-bin/
            linker?ini=asa&reqidx=/cw/asa/00031305/v56n4/s16/p331.

Agresti:2000:SEC

[AC00a] Alan Agresti and Brian Caffo. Simple and effective confidence intervals for proportions and differences of proportions result from adding two successes and two failures. The American Statistician, 54(4):280–288, November 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685779.

Alvo:2000:CHP

[AC00b] M. Alvo and P. Cabilio. Calculation of hypergeometric probabilities using Chebyshev polynomials. The American Statistician, 54(2):141–144, May 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.amstat.org/publications/tas/Alvo.htm;   http://
www.jstor.org/stable/2686033.

Anderson-Cook:2002:BRB

[AC02] C. M. Anderson-Cook. Book review: Statistical Consulting. The American Statistician, 56(4):329–??, November 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v56n4/s13/p329.

Anderson-Cook:2003:BRBa

[AC03] Christine M. Anderson-Cook. Book review: The A to Z of Mathematics: A Basic Guide. The American Statistician, 57(3):218–??, August 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://oberon.ingentaselect.com/
cgi-bin/linker?ini=asa&reqidx=/cw/asa/00031305/v57n3/
s16/p218.

Anderson-Cook:2004:BRBc

[AC04a] Christine M. Anderson-Cook. Book review: Common Errors in Statistics (and How to Avoid Them), by Phillip I. Good; James W. Hardin. The American Statistician, 58(4):359, November 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/27643597.

Anderson-Cook:2004:BRBa

[AC04b] Christine M. Anderson-Cook. Book review: Regression and ANOVA: An Integrated Approach Using SAS Software, by Keith E. Muller; Bethel A. Fetterman. The American Statistician, 58 (2):172–173, May 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/27643532.

Anderson-Cook:2004:CES

[AC04c] Christine M. Anderson-Cook. Common errors in statistics (and how to avoid them). The American Statistician, 58(4):359, November 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anderson-Cook:2005:BRB

[AC05a] Christine M. Anderson-Cook. Book review: The Chicago Guide to Writing About Numbers. The American Statistician, 59(3): 274, August 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anderson-Cook:2005:MDL

[AC05b] Christine M. Anderson-Cook. More damned lies and statistics: How numbers confuse public issues and chance: a guide to gambling, love, the stock market, & just about everything else. The American Statistician, 59(3):274–275, August 2005. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anderson-Cook:2006:BRB

[AC06a] Christine M. Anderson-Cook. Book review: The Chicago Guide to Writing About Multivariate Analysis. The American Statistician, 60(2):203–204, May 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anderson-Cook:2006:CDS

[AC06b] Christine M. Anderson-Cook. The challenge of developing statistical literacy, reasoning and thinking. The American Statistician, 60(1):99, February 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anderson-Cook:2007:QED

[AC07a] Christine M. Anderson-Cook. Quality by experimental design. The American Statistician, 61(2):183–184, May 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anderson-Cook:2007:SSQ

[AC07b] Christine M. Anderson-Cook. Six sigma quality improvement with MINITAB. The American Statistician, 61(2):185, May 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anderson-Cook:2007:VSS

[AC07c] Christine M. Anderson-Cook. Visual statistics: Seeing data with dynamic interactive graphics. The American Statistician, 61(3): 278, August 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anderson-Cook:2008:BRBa

[AC08a] Christine M. Anderson-Cook. Book review: A Quick Course in Statistical Process Control. The American Statistician, 62(2):183, May 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anderson-Cook:2008:BRBb

[AC08b] Christine M. Anderson-Cook. Book review: Modern Engineering Statistics. The American Statistician, 62(3):270–271, August 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anderson-Cook:2003:BRBb

[ACA03] Christine M. Anderson-Cook and Marcus M. Alley. Book review: Statistical Methods in Agricultural and Experimental Biology (3rd ed.). The American Statistician, 57(4):317–??, November 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v57n4/s18/p317.

Anderson-Cook:2004:BRBb

[ACL04] Christine M. Anderson-Cook and Russell V. Lenth. Book review: Step-by-Step Basic Statistics Using SAS: Student Guide, by Larry Hatcher; Step-by-Step Basic Statistics Using SAS: Exercises by Larry Hatcher. The American Statistician, 58(3):262, August 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/27643572.

Afshartous:2008:SSD

[Afs08] David Afshartous. Sample size determination for binomial proportion confidence intervals. The American Statistician, 62(1): 27–31, February 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Albert:2004:SST

[AG04] James H. Albert and James E. Gentle. Special section: Teaching computational statistics. The American Statistician, 58(1):1–??, February 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://oberon.ingentaselect.com/
cgi-bin/linker?ini=asa&reqidx=/cw/asa/00031305/v58n1/
s1/p1.

Aldworth:2002:SPM

[AH02] Jeremy Aldworth and Wherly P. Hoffman. Split-plot model with covariate: a cautionary tale. The American Statistician, 56(4):284–??, November 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://oberon.

ingentaselect.com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/
00031305/v56n4/s4/p284.

Albert:2004:ERB

[AHL04] James H. Albert, Joseph M. Hilbe, and Russell V. Lenth. Editor’s report for The American Statistician, volume 57 (2003). The American Statistician, 58(2):91–92, May 2004. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/27643515.

Alf:2007:SAI

[AL07] Cherié Alf and Sharon Lohr. Sampling assumptions in introductory statistics classes. The American Statistician, 61(1):71–77, February 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Albert:2002:CBC

[Alb02] James H. Albert. Calculated bets: Computers, gambling, and mathematical modeling to win. The American Statistician, 56 (4):329–??, November 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://oberon.

ingentaselect.com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/
00031305/v56n4/s14/p329.

Albert:2003:CSC

[Alb03a] James H. Albert. College students’ conceptions of probability. The American Statistician, 57(1):37–??, February 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v57n1/s6/p37.

Albert:2003:ER

[Alb03b] James H. Albert. Editor’s reports. The American Statistician, 57(3):153–??, August 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://oberon.

ingentaselect.com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/
00031305/v57n3/s1/p153.

Albert:2003:NCB

[Alb03c] James H. Albert. A new column in The American Statistician. The American Statistician, 57(2):94–??, May 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v57n2/s6/p94.

Albert:2005:ERBa

[Alb05a] James H. Albert. Editor’s report for The American Statistician, volume 58 (2004). The American Statistician, 59(3):205–206, August 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Albert:2005:ERBb

[Alb05b] James H. Albert. Editor’s report for The American Statistician, volume 59 (2005). The American Statistician, 59(4):281, November 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Aldrin:2000:MPU

[Ald00] Magne Aldrin. Multivariate prediction using softly shrunk reduced-rank regression. The American Statistician, 54(1):29– 34, February 2000. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685607.

Altman:2002:SCS

[Alt02] Micah Altman. Statistical computing software reviews: a review of JMP 4.03 with special attention to its numerical accuracy. The American Statistician, 56(1):72–75, February 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/3087331.

Almond:2000:DCG

[ALTY00] Russell G. Almond, Charles Lewis, John W. Tukey, and Duanli Yan. Displays for comparing a given state to many others. The American Statistician, 54(2):89–93, May 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.amstat.org/publications/tas/
Almond.htm;   http://www.jstor.org/stable/2686023.

Andersson:2009:SCA

[And09] Per Gösta Andersson. A simple correlation adjustment procedure applied to confidence interval construction. The American Statistician, 63(3):258–262, August 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Annis:2005:PSS

[Ann05a] Dave H. Annis. Probability and statistics: The science of uncertainty. The American Statistician, 59(3):276–277, August 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Annis:2005:RPH

[Ann05b] David H. Annis. Rethinking the paper helicopter. The American Statistician, 59(4):320–326, November 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Annis:2005:PSB

[Ann05c] Izabela E. Annis. PROC SQL: Beyond the basics using SAS. The American Statistician, 59(4):350, November 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Annis:2006:BAT

[Ann06a] David H. Annis. Baseball’s all-time best hitters: How statistics can level the playing field and baseball’s all-time best sluggers: Adjusted batting performance from strikeouts to home runs. The American Statistician, 60(1):97, February 2006. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Annis:2006:SMR

[Ann06b] Izabela E. Annis. Statistics in market research. The American Statistician, 60(4):345, November 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Annis:2007:DDA

[Ann07] David H. Annis. Dyadic data analysis. The American Statistician, 61(4):371, November 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2000:EC

[Ano00a] Anonymous. 2000 editorial collaborators. The American Statistician, 54(4):329–330, November 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685792.

Anonymous:2000:AMS

[Ano00b] Anonymous. Applied Multivariate Statistics with SAS Software. The American Statistician, 54(2):154–??, May 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2000:BMa

[Ano00c] Anonymous. Back matter. The American Statistician, 54(1): ??, February 2000. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685625.

Anonymous:2000:BMb

[Ano00d] Anonymous. Back matter. The American Statistician, 54(2):??, May 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2686043.

Anonymous:2000:BMc

[Ano00e] Anonymous. Back matter. The American Statistician, 54(3): ??, August 2000. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685600.

Anonymous:2000:BMd

[Ano00f] Anonymous. Back matter. The American Statistician, 54(4): ??, November 2000. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685793.

Anonymous:2000:FMa

[Ano00g] Anonymous. Front matter. The American Statistician, 54(1): ??, February 2000. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685601.

Anonymous:2000:FMb

[Ano00h] Anonymous. Front matter. The American Statistician, 54(2):??, May 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2686022.

Anonymous:2000:FMc

[Ano00i] Anonymous. Front matter. The American Statistician, 54(3): ??, August 2000. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685583.

Anonymous:2000:FMd

[Ano00j] Anonymous. Front matter. The American Statistician, 54(4): ??, November 2000. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685771.

Anonymous:2000:IV

[Ano00k] Anonymous. Index of volume 54. The American Statistician, 54 (4):331–??, November 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2000:LEa

[Ano00l] Anonymous. Letters to the Editor. The American Statistician, 54(2):156–??, May 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2000:LEb

[Ano00m] Anonymous. Letters to the Editor. The American Statistician, 54(3):227–??, August 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2000:LEc

[Ano00n] Anonymous. Letters to the Editor. The American Statistician, 54 (4):325–??, November 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2000:VI

[Ano00o] Anonymous. Volume information. The American Statistician, 54(4):331–332, November 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685770.

Anonymous:2001:EC

[Ano01a] Anonymous. 2001 editorial collaborators. The American Statistician, 55(4):377, November 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685710.

Anonymous:2001:BMa

[Ano01b] Anonymous. Back matter. The American Statistician, 55(1): ??, February 2001. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685541.

Anonymous:2001:BMb

[Ano01c] Anonymous. Back matter. The American Statistician, 55(2):??, May 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2686011.

Anonymous:2001:BMc

[Ano01d] Anonymous. Back matter. The American Statistician, 55(3): ??, August 2001. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685817.

Anonymous:2001:BMd

[Ano01e] Anonymous. Back matter. The American Statistician, 55(4): ??, November 2001. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685711.

Anonymous:2001:FMa

[Ano01f] Anonymous. Front matter. The American Statistician, 55(1): ??, February 2001. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685521.

Anonymous:2001:FMb

[Ano01g] Anonymous. Front matter. The American Statistician, 55(2):??, May 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685991.

Anonymous:2001:FMc

[Ano01h] Anonymous. Front matter. The American Statistician, 55(3): ??, August 2001. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685794.

Anonymous:2001:FMd

[Ano01i] Anonymous. Front matter. The American Statistician, 55(4): ??, November 2001. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685685.

Anonymous:2001:LE

[Ano01j] Anonymous. Letters to the Editor. The American Statistician, 55(1):85–??, February 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2001:VI

[Ano01k] Anonymous. Volume information. The American Statistician, 55(4):378–380, November 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685684.

Anonymous:2002:BM

[Ano02a] Anonymous. Back matter. The American Statistician, 56(1): ??, February 2002. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
3087342.

Anonymous:2002:C

[Ano02b] Anonymous. Corrections. The American Statistician, 56(3):258– ??, August 2002. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://oberon.ingentaselect.
com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/00031305/v56n3/
s22/p258.

Anonymous:2002:FM

[Ano02c] Anonymous. Front matter. The American Statistician, 56(1): ??, February 2002. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
3087319.

Anonymous:2002:LEa

[Ano02d] Anonymous. Letters to the Editor. The American Statistician, 56(2):159–??, May 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://oberon.

ingentaselect.com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/
00031305/v56n2/s14/p159.

Anonymous:2002:LEb

[Ano02e] Anonymous. Letters to the Editor. The American Statistician, 56(3):252–??, November 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://oberon.

ingentaselect.com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/
00031305/v56n3/s21/p252.

Anonymous:2002:LEc

[Ano02f] Anonymous. Letters to the Editor. The American Statistician, 56(4):337–??, November 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://oberon.

ingentaselect.com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/
00031305/v56n4/s24/p337.

Anonymous:2002:RBT

[Ano02g] Anonymous. Reviews of books and teaching materials. The American Statistician, 56(2):156–??, May 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://oberon.ingentaselect.com/cgi-bin/linker?
ini=asa&reqidx=/cw/asa/00031305/v56n2/s13/p156.

Anonymous:2003:BRT

[Ano03a] Anonymous. Brief reviews of teaching materials. The American Statistician, 57(4):319–??, February 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v57n4/s20/p318.

Anonymous:2003:Ca

[Ano03b] Anonymous. Correction. The American Statistician, 57(2):151– ??, May 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://oberon.ingentaselect.com/
cgi-bin/linker?ini=asa&reqidx=/cw/asa/00031305/v57n2/
s21/p151.

Anonymous:2003:Cb

[Ano03c] Anonymous. Correction. The American Statistician, 57(3):227– ??, August 2003. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://oberon.ingentaselect.
com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/00031305/v57n3/
s21/p227.

Anonymous:2003:Cc

[Ano03d] Anonymous. Corrections. The American Statistician, 57(4):322– ??, November 2003. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://oberon.ingentaselect.
com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/00031305/v57n4/
s21/p322.

Anonymous:2003:LEa

[Ano03e] Anonymous. Letters to the Editor. The American Statistician, 57(1):71–??, February 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://oberon.

ingentaselect.com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/
00031305/v57n1/s19/p71.

Anonymous:2003:LEb

[Ano03f] Anonymous. Letters to the Editor. The American Statistician, 57(2):147–??, February 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://oberon.

ingentaselect.com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/
00031305/v57n2/s20/p147.

Anonymous:2003:LEc

[Ano03g] Anonymous. Letters to the Editor. The American Statistician, 57(3):221–??, February 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://oberon.

ingentaselect.com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/
00031305/v57n3/s20/p221.

Anonymous:2003:LEd

[Ano03h] Anonymous. Letters to the Editor. The American Statistician, 57(4):319–??, February 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://oberon.

ingentaselect.com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/
00031305/v57n4/s20/p319.

Anonymous:2003:R

[Ano03i] Anonymous. Rejoinder. The American Statistician, 57(3):189–??, August 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://oberon.ingentaselect.com/

cgi-bin/linker?ini=asa&reqidx=/cw/asa/00031305/v57n3/
s10/p181.

Anonymous:2004:EC

[Ano04a] Anonymous. 2004 editorial collaborators. The American Statistician, 58(4):365–368, November 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2004:ASA

[Ano04b] Anonymous. Applied Statistics: Analysis of Variance and Regression by Ruth M. Mickey; Olive Jean Dunn; Virginia A. Clark. The American Statistician, 58(3):262, August 2004. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/27643573.

Anonymous:2004:BMa

[Ano04c] Anonymous. Back matter. The American Statistician, 58(2):??, May 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/27643548.

Anonymous:2004:BMb

[Ano04d] Anonymous. Back matter. The American Statistician, 58(3): ??, August 2004. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
27643577.

Anonymous:2004:BMc

[Ano04e] Anonymous. Back matter. The American Statistician, 58(4): ??, November 2004. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
27643604.

Anonymous:2004:BRT

[Ano04f] Anonymous. Brief reviews of teaching materials. The American Statistician, 58(3):262–263, August 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2004:Cb

[Ano04g] Anonymous. Correction. The American Statistician, 58(4):364, November 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2004:Ca

[Ano04h] Anonymous. Corrections. The American Statistician, 58(2):178, May 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2004:DEA

[Ano04i] Anonymous. Designing Experiments and Analyzing Data: a Model Comparison Perspective by Scott E. Maxwell; Harold D. Delaney. The American Statistician, 58(3):262, August 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/27643574.

Anonymous:2004:FMa

[Ano04j] Anonymous. Front matter. The American Statistician, 58(2):??, May 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/27643514.

Anonymous:2004:FMb

[Ano04k] Anonymous. Front matter. The American Statistician, 58(3): ??, August 2004. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
27643550.

Anonymous:2004:FMc

[Ano04l] Anonymous. Front matter. The American Statistician, 58(4): ??, November 2004. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
27643580.

Anonymous:2004:LEa

[Ano04m] Anonymous. Letters to the Editor. The American Statistician, 58(1):90–??, February 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://oberon.

ingentaselect.com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/
00031305/v58n1/s23/p90.

Anonymous:2004:LEb

[Ano04n] Anonymous. Letters to the Editor. The American Statistician, 58(2):176–177, May 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2004:LEc

[Ano04o] Anonymous. Letters to the Editor. The American Statistician, 58 (4):362–363, November 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2004:MTA

[Ano04p] Anonymous. Math through the Ages: a Gentle History for Teachers and Others by William P. Berlinghoff; Fernando Q. Gouvêa; Sherlock Holmes in Babylon and Other Tales of Mathematical History by Marlow Anderson; Victor Katz; Robin Wilson. The American Statistician, 58(3):262, August 2004. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/27643575.

Anonymous:2004:RSC

[Ano04q] Anonymous. Reply: Statistics and the college football championship. The American Statistician, 58(3):194–195, August 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2004:VI

[Ano04r] Anonymous. Volume information. The American Statistician, 58(4):??, November 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/27643579.

Anonymous:2004:WPM

[Ano04s] Anonymous. Writing Projects for Mathematics Courses: Crashed Clowns, Cars, and Coffee to Go by Annalisa Crannell; Gavin Larose; Thomas Ratliff; Elyn Rykken. The American Statistician, 58(3):262–263, August 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/27643576.

Anonymous:2005:EC

[Ano05a] Anonymous. 2005 editorial collaborators. The American Statistician, 59(4):355–358, November 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2005:Ca

[Ano05b] Anonymous. Correction. The American Statistician, 59(1):119, February 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2005:Cb

[Ano05c] Anonymous. Correction. The American Statistician, 59(3):280, August 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2005:LEa

[Ano05d] Anonymous. Letters to the Editor. The American Statistician, 59 (1):116–118, February 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2005:LEb

[Ano05e] Anonymous. Letters to the Editor. The American Statistician, 59 (3):279, August 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2006:EC

[Ano06a] Anonymous. 2006 editorial collaborators. The American Statistician, 60(4):350–353, November 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2006:C

[Ano06b] Anonymous. Correction. The American Statistician, 60(2):212, May 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2006:LEa

[Ano06c] Anonymous. Letters to the Editor. The American Statistician, 60 (1):101–103, February 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2006:LEb

[Ano06d] Anonymous. Letters to the Editor. The American Statistician, 60(2):208–211, May 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2006:LEc

[Ano06e] Anonymous. Letters to the Editor. The American Statistician, 60(3):295–297, August 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2006:LEd

[Ano06f] Anonymous. Letters to the Editor. The American Statistician, 60 (4):347–349, November 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2007:EC

[Ano07a] Anonymous. 2007 editorial collaborators. The American Statistician, 61(4):374–379, November 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2007:BRTa

[Ano07b] Anonymous. Brief reviews of teaching materials. The American Statistician, 61(2):186, May 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2007:BRTb

[Ano07c] Anonymous. Brief reviews of teaching materials. The American Statistician, 61(3):279, August 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2007:BRTc

[Ano07d] Anonymous. Brief reviews of teaching materials. The American Statistician, 61(4):372, November 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2007:LEa

[Ano07e] Anonymous. Letters to the Editor. The American Statistician, 61(2):187–188, May 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2007:LEb

[Ano07f] Anonymous. Letters to the Editor. The American Statistician, 61(3):280–284, August 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2007:LEc

[Ano07g] Anonymous. Letters to the Editor. The American Statistician, 61(4):373, November 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2008:EC

[Ano08a] Anonymous. 2008 editorial collaborators. The American Statistician, 62(4):366–367, November 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2008:BRTa

[Ano08b] Anonymous. Brief reviews of teaching materials. The American Statistician, 62(1):94, February 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2008:BRTb

[Ano08c] Anonymous. Brief reviews of teaching materials. The American Statistician, 62(2):185, May 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2008:C

[Ano08d] Anonymous. Corrections. The American Statistician, 62(2):188, May 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2008:IV

[Ano08e] Anonymous. Index to volume 62. The American Statistician, 62 (4):368–370, November 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2008:LEa

[Ano08f] Anonymous. Letters to the Editor. The American Statistician, 62(1):95–96, February 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2008:LEb

[Ano08g] Anonymous. Letters to the Editor. The American Statistician, 62(2):186–187, May 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2008:LEc

[Ano08h] Anonymous. Letters to the Editor. The American Statistician, 62(3):274–278, August 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2008:LEd

[Ano08i] Anonymous. Letters to the Editor. The American Statistician, 62 (4):364–365, November 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2009:EC

[Ano09a] Anonymous. 2009 editorial collaborators. The American Statistician, 63(4):407–408, November 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2009:E

[Ano09b] Anonymous. Errata. The American Statistician, 63(2):200, May 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2009:IV

[Ano09c] Anonymous. Index to volume 63. The American Statistician, 63 (4):409–411, November 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2009:LEa

[Ano09d] Anonymous. Letter to the Editor. The American Statistician, 63 (1):102–103, February 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2009:LEc

[Ano09e] Anonymous. Letter to the Editor. The American Statistician, 63(4):406, November 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2009:LEb

[Ano09f] Anonymous. Letters to the Editor. The American Statistician, 63(2):197–199, May 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2009:RBTa

[Ano09g] Anonymous. Reviews of books and teaching materials. The American Statistician, 63(1):92–101, February 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2009:RBTb

[Ano09h] Anonymous. Reviews of books and teaching materials. The American Statistician, 63(2):189–196, May 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2009:RBTc

[Ano09i] Anonymous. Reviews of books and teaching materials. The American Statistician, 63(3):281–293, August 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2009:RBTd

[Ano09j] Anonymous. Reviews of books and teaching materials. The American Statistician, 63(4):401–405, November 2009. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Athanasios:2007:EEB

[APT07] Diamandopoulos A. Athanasios, Goudas C. Pavlos, and Kassimatis I. Theodoros. Early evidence-based medicine. The American Statistician, 61(2):154–158, May 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Allenby:2008:TBS

[AR08] Greg M. Allenby and Peter E. Rossi. Teaching Bayesian statistics to marketing and business students. The American Statistician, 62(3):195–198, August 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Arnold:2006:DEM

[Arn06] Steven F. Arnold. Design of experiments with MINITAB. The American Statistician, 60(2):205, May 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Arnholt:2007:SAS

[Arn07] Alan T. Arnholt. Statistics: The art and science of learning from data. The American Statistician, 61(2):181–182, May 2007. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Austin:2004:BMD

[AT04] Peter C. Austin and Jack V. Tu. Bootstrap methods for developing predictive models. The American Statistician, 58(2):131–137, May 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/27643521.

Albert:2001:UMD

[AW01] Jim Albert and Patricia Williamson. Using model/data simulations to detect streakiness. The American Statistician, 55(1):41–50, February 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

amstat.org/publications/tas/Albert.htm;   http://www.
jstor.org/stable/2685528.

Annis:2006:CPP

[AW06] David H. Annis and Samuel S. Wu. A comparison of potential playoff systems for NCAA I-A football. The American Statistician, 60(2):151–157, May 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Brady:2002:CSB

[BA02] James E. Brady and Theodore T. Allen. Case study based instruction of DOE and SPC. The American Statistician, 56(4):312–??, November 2002. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://oberon.ingentaselect.
com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/00031305/v56n4/
s9/p312.

Berry:2006:SPC

[BA06] Donald A. Berry and Gregory D. Ayers. Symmetrized percent change for treatment comparisons. The American Statistician, 60(1):27–31, February 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Barker:2002:CNC

[Bar02a] Lawrence Barker. A comparison of nine confidence intervals for a Poisson parameter when the expected number of events is 5. The American Statistician, 56(2):85–??, May 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v56n2/s2/p85.

Bartko:2002:SGJ

[Bar02b] John J. Bartko. Schenker, N., and Gentleman, J. F. (2001), “On Judging the Significance of Differences by Examining the Overlap between Confidence Intervals,” The American Statistician, 55, 182-186: Comment by Bartko. The American Statistician, 56(1): 81, February 2002. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
3087339.

Barrett:2007:DBJ

[Bar07] Bruce E. Barrett. Detecting bias in jury selection. The American Statistician, 61(4):296–301, November 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Barker:2009:MMC

[Bar09] Chris Barker. The mean, median, and confidence intervals of the Kaplan–Meier survival estimate — computations and applications. The American Statistician, 63(1):78–80, February 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bates:2004:BRB

[Bat04] Samantha C. Bates. Book review: Introductory Statistics with R. The American Statistician, 58(1):88–??, February 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v58n1/s20/p88.

Barry:2004:DUS

[BB04] Daniel Barry and Philip J. Boland. Debating the use of statistical evidence in allegations of age discrimination. The American Statistician, 58(2):102–109, May 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/27643517.

Bryant:2000:MSU

[BCB00] Peter G. Bryant and Olga I. Cordero-Brana. Model selection using the minimum description length principle. The American Statistician, 54(4):257–268, November 2000. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.amstat.org/publications/tas/
Bryant.htm;   http://www.jstor.org/stable/2685777.

Berger:2001:LE

[BCW01] Roger L. Berger, Basil W. Coutant, and Don G. Wardell. Letters to the Editor. The American Statistician, 55(1):85, February 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685540.

Bellhouse:2001:CLT

[Bel01] D. R. Bellhouse. The Central Limit Theorem under simple random sampling. The American Statistician, 55(4):352–357, November 2001. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685700.

Belden:2005:DMB

[Bel05] Susan Belden. Discussion: Moving beyond the St. Petersburg Paradox to understand growth stock valuation. The American Statistician, 59(3):233–234, August 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bellhouse:2007:SME

[Bel07] David R. Bellhouse. Sampling methods: Exercises and solutions. The American Statistician, 61(3):276, August 2007. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Berk:2000:BRBa

[Ber00a] Kenneth N. Berk. Book review: [untitled]. The American Statistician, 54(1):82, February 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685619.

Berk:2000:BRBb

[Ber00b] Kenneth N. Berk. Book review: [untitled]. The American Statistician, 54(1):82–83, February 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685620.

Berk:2000:BRBc

[Ber00c] Kenneth N. Berk. Book review: [untitled]. The American Statistician, 54(2):154–155, May 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2686040.

Berk:2003:D

[Ber03] Kenneth N. Berk. Discussion. The American Statistician, 57(3): 178–??, August 2003. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://oberon.ingentaselect.
com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/00031305/v57n3/
s6/p178.

Berengut:2006:SED

[Ber06] David Berengut. Statistics for experimenters: Design, innovation, and discovery. The American Statistician, 60(4):341–342, November 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Berrendero:2007:BMB

[Ber07] José R. Berrendero. The bagged median and the bragged mean. The American Statistician, 61(4):325–330, November 2007. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Betensky:2000:ADR

[Bet00] Rebecca A. Betensky. Alternative derivations of a rule for early stopping in favor of H_{0}. The American Statistician, 54(1):35– 39, February 2000. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685608.

Bell:2008:ETP

[BG08] Richard C. Bell, Jr. and Andrew G. Glen. Experiences teaching probability and statistics with personal laptops in the classroom daily. The American Statistician, 62(2):155–160, May 2008. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Busstra:2008:DDD

[BGF^{+}08] M. C. Busstra, A. Geelen, E. J. Feskens, R. J. M. Hartog, and P. van ’t Veer. Design and development of digital learning material for applied data analysis. The American Statistician, 62(4): 329–339, November 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bryce:2001:CGB

[BGNP01] G. Rex Bryce, Robert Gould, William I. Notz, and Roxy L. Peck. Curriculum guidelines for Bachelor of Science degrees in statistical science. The American Statistician, 55(1):7–13, February 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685523.

Bhattacharya:2002:MVU

[BH02] Bhaskar Bhattacharya and Desale Habtzghi. Median of the p value under the alternative hypothesis. The American Statistician, 56(3):202–??, August 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://oberon.

ingentaselect.com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/
00031305/v56n3/s7/p202.

Brimberg:2008:OPN

[BH08] Jack Brimberg and W. J. Hurley. Is the overtime period in an NHL game long enough? an example for teaching estimation and hypothesis testing in the presence of censored data. The American Statistician, 62(2):151–154, May 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bar-Hillel:2002:SWA

[BHA02] Maya Bar-Hillel and Yigal Attali. Seek whence: Answer sequences and their consequences in key-balanced multiple-choice tests. The American Statistician, 56(4):299–??, November 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v56n4/s6/p299.

Bolstad:2001:TCS

[BHM01] W. M. Bolstad, Lyn A. Hunt, and Judith L. McWhirter. Teacher’s corner: Sex, drugs, and rock & roll survey in a first-year service course in statistics. The American Statistician, 55(2):145–149, May 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2686002.

Boos:2000:HLD

[BHO00] Dennis D. Boos and Jacqueline M. Hughes-Oliver. How large does n have to be for Z and t intervals? The American Statistician, 54(2):121–128, May 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.amstat.org/publications/tas/Boos.htm;   http://
www.jstor.org/stable/2686030.

Bornkamp:2009:NBS

[BI09] Björn Bornkamp and Katja Ickstadt. A note on B-splines for semiparametric elicitation. The American Statistician, 63(4):373– 377, November 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Billingsley:2004:DSC

[Bil04] Richard Billingsley. Discussion: Statistics and the college football championship. The American Statistician, 58(3):190, August 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Baker:2007:PYR

[BK07] Stuart G. Baker and Barnett S. Kramer. Peirce, Youden, and receiver operating characteristic curves. The American Statistician, 61(4):343–346, November 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Brown:2009:RWS

[BK09a] Emery N. Brown and Robert E. Kass. Rejoinder: What is statistics? The American Statistician, 63(2):122–123, May 2009. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Brown:2009:WS

[BK09b] Emery N. Brown and Robert E. Kass. What is statistics? The American Statistician, 63(2):105–110, May 2009. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bergmann:2000:DOW

[BLS00] Reinhard Bergmann, John Ludbrook, and Will P. J. M. Spooren. Different outcomes of the Wilcoxon–Mann–Whitney test from different statistics packages. The American Statistician, 54(1):72– 77, February 2000. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685616.

Balkin:2001:AAT

[BM01] Sandy D. Balkin and Colin L. Mallows. An adjusted, asymmetric two-sample t test. The American Statistician, 55(3):203–206, August 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685800.

Bayarri:2002:BDS

[BM02a] M. J. Bayarri and A. M. Mayoral. Bayesian design of “successful” replications. The American Statistician, 56(3):207–??, August 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://oberon.ingentaselect.com/
cgi-bin/linker?ini=asa&reqidx=/cw/asa/00031305/v56n3/
s8/p207.

Bessant:2002:TOC

[BM02b] Kenneth C. Bessant and Eric D. MacPherson. Thoughts on the origins, concepts, and pedagogy of statistics as a “separate discipline”. The American Statistician, 56(1):22–28, February 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/3087323.

Barrowman:2003:RPN

[BM03] Nicholas J. Barrowman and Ransom A. Myers. Raindrop plots: a new way to display collections of likelihoods and distributions. The American Statistician, 57(4):268–??, November 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v57n4/s10/p268.

Birch:2005:TVT

[BM05] Jeffrey B. Birch and J. P. Morgan. TA training at Virginia Tech. The American Statistician, 59(1):14–18, February 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Boland:2003:PST

[Bol03] Philip J. Boland. Promoting statistical thinking among secondary school students in the national context. The American Statistician, 57(2):85–??, May 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://oberon.ingentaselect.com/cgi-bin/linker?
ini=asa&reqidx=/cw/asa/00031305/v57n2/s4/p85.

Bond:2002:BRB

[Bon02] Marjorie E. Bond. Book review: Teach/Me Data Analysis CD: Single User Edition. The American Statistician, 56(4):335–??, November 2002. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://oberon.ingentaselect.
com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/00031305/v56n4/
s23/p335.

Bonett:2004:CCI

[Bon04] Douglas G. Bonett. Corrections: Confidence intervals for mean absolute deviations. The American Statistician, 58(2):178, May 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/27643544.

Bordley:2002:RTU

[Bor02] Robert F. Bordley. Representing trees using Microsoft doughnut charts. The American Statistician, 56(2):139–??, May 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v56n2/s10/p139.

Box:2005:RRF

[Box05] Joan Fisher Box. A reminiscence of R. A. Fisher. The American Statistician, 59(4):312–314, November 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Boyles:2008:RLI

[Boy08] Russell A. Boyles. The role of likelihood in interval estimation. The American Statistician, 62(1):22–26, February 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bradstreet:2002:INP

[BP02] Thomas E. Bradstreet and Milton N. Parnes. Illustrating the Neyman–Pearson lemma with a stopping rule of order k. The American Statistician, 56(2):113–??, May 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v56n2/s7/p113.

Blume:2003:ILL

[BR03]  Jeffrey D. Blume and Richard M. Royall.  Illustrating

the law of large numbers (and confidence intervals). The American Statistician, 57(1):51–??, February 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://oberon.ingentaselect.com/cgi-bin/linker?
ini=asa&reqidx=/cw/asa/00031305/v57n1/s8/p51.

Brackstone:2007:DOS

[Bra07] Gordon Brackstone. Discussion: The official statistics Olympic challenge. The American Statistician, 61(1):9–12, February 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bradstreet:2008:GLD

[Bra08a] Thomas E. Bradstreet. Graphics of large datasets: Visualizing a million. The American Statistician, 62(2):180–181, May 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bradstreet:2008:PSU

[Bra08b] Thomas E. Bradstreet. Pharmaceutical statistics using SAS: a practical guide. The American Statistician, 62(2):183, May 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Browne:2001:SPU

[Bro01] Richard H. Browne. Statistical practice: Using the sample range as a basis for calculating sample size in power calculations. The American Statistician, 55(4):293–298, November 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685690.

Brown:2007:SLB

[Bro07] Aaron Brown. Strong language on black swans. The American Statistician, 61(3):195–197, August 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Broers:2008:HSB

[Bro08] Nick J. Broers. Helping students to build a conceptual understanding of elementary statistics. The American Statistician, 62(2):161–166, May 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Barker:2001:SPE

[BRRB01] Lawrence Barker, Henry Rolka, Deborah Rolka, and Cedric Brown. Statistical practice: Equivalence testing for binomial random variables: Which test to use? The American Statistician, 55(4):279–287, November 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685688.

Bingham:2000:EHG

[BS00a] Derek R. Bingham and Tim B. Swartz. Equitable handicapping in golf. The American Statistician, 54(3):170–177, August 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.amstat.org/publications/tas/
Bingham.htm;   http://www.jstor.org/stable/2685586.

Blaker:2000:PII

[BS00b] Helge Blaker and Emil Spjotvoll. Paradoxes and improvements in interval estimation. The American Statistician, 54(4):242–247, November 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

amstat.org/publications/tas/Blaker.htm;   http://www.
jstor.org/stable/2685774.

Bonett:2003:CIM

[BS03] Douglas G. Bonett and Edith Seier. Confidence intervals for mean absolute deviations. The American Statistician, 57(4):233– ??, November 2003. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://oberon.ingentaselect.
com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/00031305/v57n4/
s2/p233.

Bathke:2009:GGA

[BSTM09] Arne C. Bathke, Oliver Schabenberger, Randall D. Tobias, and Laurence V. Madden. Greenhouse–Geisser adjustment and the ANOVA-type statistic: Cousins or twins? The American Statistician, 63(3):239–246, August 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Berenson:2008:ASR

[BUK^{+}08] Mark L. Berenson, Jessica Utts, Karen A. Kinard, Deborah J. Rumsey, Albyn Jones, and Leonard M. Gaines. Assessing student retention of essential statistical ideas. The American Statistician, 62(1):54–61, February 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bull:2005:ES

[Bul05] Brian Bull. Exemplar sampling. The American Statistician, 59(2):166–172, May 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bhatti:2008:CPH

[BW08] Chad R. Bhatti and Jennifer L. Wightman. Conditional probability and HIV testing. The American Statistician, 62(3):238–241, August 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Callaert:2003:CSS

[Cal03] Herman Callaert. Comparing statistical software packages: The case of the logrank test in StatXact. The American Statistician, 57(3):214–??, August 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://oberon.

ingentaselect.com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/
00031305/v57n3/s15/p214.

Cannon:2001:BRB

[Can01] Ann Cannon. Book review: Essential Statistics (3rd ed.). The American Statistician, 55(1):83, February 2001. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685537.

Carlton:2000:BRB

[Car00] Matthew A. Carlton. Book review: Data Analysis: Statistical and Computational Methods for Scientists and Engineers. The American Statistician, 54(2):155, May 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2686041.

Carlton:2003:D

[Car03] Matthew A. Carlton. Discussion. The American Statistician, 57(3):179–??, August 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://oberon.

ingentaselect.com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/
00031305/v57n3/s7/p179.

Carlton:2008:PSC

[Car08] Matthew A. Carlton. Probability and statistics for computer scientists. The American Statistician, 62(3):271–272, August 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cavanaugh:2007:STS

[Cav07a] Joseph Cavanaugh. Seeing through statistics and mind on statistics. The American Statistician, 61(1):95, February 2007. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cavanaugh:2007:SEA

[Cav07b] Joseph Cavanaugh. Statistics: The exploration and analysis of data. The American Statistician, 61(1):97–98, February 2007.

CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Chen:2009:RPL

[CBH09] Aiyou Chen, Thomas Bengtsson, and Tin Kam Ho. A regression paradox for linear models: Sufficient conditions and relation to Simpson’s Paradox. The American Statistician, 63(3):218–225, August 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cohen:2006:SDP

[CC06] Dale J. Cohen and Jon Cohen. The sectioned density plot. The American Statistician, 60(2):167–174, May 2006. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cabilio:2001:TCC

[CF01] P. Cabilio and P. J. Farrell. Teacher’s corner: a computer-based lab supplement to courses in introductory statistics. The American Statistician, 55(3):228–232, August 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2685806.

Chatterjee:2007:GDI

[CF07] Sangit Chatterjee and Aykut Firat. Generating data with identical statistics but dissimilar graphics. The American Statistician, 61(3):248–254, August 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Christensen:2008:PPO

[CF08] William F. Christensen and Lindsay W. Florence. Predicting presidential and other multistage election outcomes using statelevel pre-election polls. The American Statistician, 62(1):1–10, February 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Caffo:2006:UFI

[CG06] Brian Caffo and Michael Griswold. A user-friendly introduction to link-probit-normal models. The American Statistician, 60(2): 139–145, May 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cho:2007:BBL

[CG07] Wendy K. Tam Cho and Brian J. Gaines. Breaking the (Benford) law: statistical fraud detection in campaign finance. The American Statistician, 61(3):218–223, August 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

pubs.amstat.org/doi/pdfplus/10.1198/000313007X223496;
 http://www.tandfonline.com/doi/abs/10.1198/000313007X223496.

Chan:2004:BRB

[Cha04] Grace Chan. Book review: SAS Programming: The One-Day Course, by Neil H. Spencer. The American Statistician, 58(3): 260, August 2004. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
27643570.

Chapman:2008:ESP

[Cha08] Jessica Chapman. Elementary statistics: Picturing the world. The American Statistician, 62(4):357–358, November 2008. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Chen:2000:BRB

[Che00] Victoria Chen. Book review: [untitled]. The American Statistician, 54(1):81–82, February 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685618.

Chitturi:2008:IBS

[Chi08] Pallavi Chitturi. Introduction to business statistics. The American Statistician, 62(3):269, August 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Christensen:2008:PNS

[CHJ08] Ronald Christensen, Timothy Hanson, and Alejandro Jara. Parametric nonparametric statistics. The American Statistician, 62 (4):296–306, November 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Christensen:2003:SIT

[Chr03] Ronald Christensen. Significantly insignificant F tests. The American Statistician, 57(1):27–??, February 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://oberon.ingentaselect.com/cgi-bin/linker?
ini=asa&reqidx=/cw/asa/00031305/v57n1/s4/p27.

Christensen:2005:TFN

[Chr05] Ronald Christensen. Testing Fisher, Neyman, Pearson, and Bayes. The American Statistician, 59(2):121–126, May 2005. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cantwell:2004:USM

[CHS04] Patrick J. Cantwell, Howard Hogan, and Kathleen M. Styles. The use of statistical methods in the U.S. Census: “Utah v. Evans”. The American Statistician, 58(3):203–212, August 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/27643558.

Chen:2001:SPN

[CK01] Zhen Chen and Lynn Kuo. Statistical practice: a note on the estimation of the multinomial logit model with random effects. The American Statistician, 55(2):89–95, May 2001. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685993.

Chernick:2002:STB

[CL02] Michael R. Chernick and Christine Y. Liu. The sawtoothed behavior of power versus sample size and software solutions: Single binomial proportion using exact methods. The American Statistician, 56(2):149–??, May 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v56n2/s12/p149.

Chakraborti:2007:CIE

[CL07] S. Chakraborti and J. Li. Confidence interval estimation of a normal percentile. The American Statistician, 61(4):331–336, November 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cleary:2000:BRBa

[Cle00a] Richard Cleary. Book review: How to Think About Statistics. The American Statistician, 54(4):322, November 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685787.

Cleary:2000:BRBb

[Cle00b] Richard Cleary. Book review: Statistics: An Introduction, 5th edition. The American Statistician, 54(4):324, November 2000.

CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2685790.

Cleary:2001:RBT

[Cle01] Richard J. Cleary. Reviews of books and teaching materials: Quantitative Investigations in the Biosciences using MINITAB^{T M}. The American Statistician, 55(2):167–168, May 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2686007.

Cleary:2003:BRBb

[Cle03a] Richard J. Cleary. Book review: Statistical Rules of Thumb. The American Statistician, 57(2):142–??, May 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v57n2/s14/p142.

Cleary:2003:BRBa

[Cle03b] Richard J. Cleary. Book review: Statistics with Confidence (2nd ed.). The American Statistician, 57(1):66–??, February 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v57n1/s11/p66.

Cleary:2003:BRBc

[Cle03c] Richard J. Cleary. Book review: The First Measured Century. The American Statistician, 57(2):143–??, May 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v57n2/s16/p143.

Cleary:2003:BRBd

[Cle03d] Richard J. Cleary. Book reviews: Duelling Idiots and Other Probability Puzzlers and What are the Chances? Voodoo Deaths and Office Gossip and Other Adventures in Probability. The American Statistician, 57(3):218–??, August 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v57n3/s18/p218.

Cleary:2004:BRB

[Cle04] Richard J. Cleary. Book review: Teaching Statistics Using Baseball, by Jim Albert. The American Statistician, 58(4):358, November 2004. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
27643596.

Cleary:2005:GTS

[Cle05a] Richard J. Cleary. Gauss: Titan of science and Karl Pearson: The scientific life in a statistical age. The American Statistician, 59(4):349, November 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://pubs.amstat.org/

doi/pdfplus/10.1198/tas.2005.s31.

Cleary:2005:TSB

[Cle05b] Richard J. Cleary. Teaching statistics: a bag of tricks. The American Statistician, 59(3):275, August 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cleary:2006:ITS

[Cle06a] Richard J. Cleary. Innovations in teaching statistics and achieving quantitative literacy: An urgent challenge for higher education. The American Statistician, 60(1):99–100, February 2006. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cleary:2006:PVG

[Cle06b] Richard J. Cleary. Probability is the very guide of life: The philosophical uses of chance. The American Statistician, 60(3): 290, August 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cleary:2006:WNS

[Cle06c] Richard J. Cleary. Working with numbers and statistics: a handbook for journalists. The American Statistician, 60(2):203, May 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cleary:2008:FPS

[Cle08a] Richard J. Cleary. Fundamentals of probability and statistics for engineers, statistical techniques for data analysis and statistics: An introduction. The American Statistician, 62(1):90–91, February 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cleary:2008:SID

[Cle08b] Richard J. Cleary. Statistics: Informed decisions using data and an introduction to statistical analysis for business and industry: a problem solving approach. The American Statistician, 62(2): 183–184, May 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Chang:2008:NIA

[CLPC08] Ching-Hui Chang, Jyh-Jiuan Lin, Nabendu Pal, and Miao-Chen Chiang. A note on improved approximation of the binomial distribution by the skew-normal distribution. The American Statistician, 62(2):167–170, May 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Casella:2001:EPS

[CLR01] George Casella, Michael Lavine, and Christian P. Robert. Explaining the perfect sampler. The American Statistician, 55(4): 299–305, November 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685691.

Caffo:2013:CSM

[CLR13] Brian Caffo, Carolyn Lauzon, and Joachim Röhmel. Correction to “Easy Multiplicity Control in Equivalence Testing Using Two One-Sided Tests”. The American Statistician, 67(2):115– 116, 2013. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See [LC09].

Chung:2004:DDI

[CLS04] Hwan Chung, Eric Loken, and Joseph L. Schafer. Difficulties in drawing inferences with finite-mixture models: a simple example with a simple solution. The American Statistician, 58(2):152–158, May 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/27643526.

Capotorti:2007:RAR

[CLS07] Andrea Capotorti, Frank Lad, and Giuseppe Sanfilippo. Reassessing accuracy rates of median decisions. The American Statistician, 61(2):132–138, May 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cobb:2000:BRB

[Cob00] George W. Cobb. Book review: [untitled]. The American Statistician, 54(1):78–81, February 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685617.

Cochran:2005:CYR

[Coc05a] James J. Cochran. Can you really learn basic probability by playing a sports board game? The American Statistician, 59 (3):266–272, August 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cochran:2005:NGB

[Coc05b] James J. Cochran. The numbers game: Baseball’s lifelong fascination with statistics. The American Statistician, 59(2):203, May 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cohen:2002:BRB

[Coh02] Arthur Cohen. Book review: An Introduction to Probability Theory and Mathematical Statistics, (2nd ed.). The American Statistician, 56(4):331–??, November 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v56n4/s17/p331.

Collings:2001:RBT

[Col01] Patti B. Collings. Reviews of books and teaching materials: Fathom: Dynamic Systems Software, by William Finzer; Tim Erickson; Jill Bunker. The American Statistician, 55(3):258– 259, August 2001. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685814.

Colleys:2004:DSC

[Col04] Wes Colleys. Discussion: Statistics and the college football championship. The American Statistician, 58(3):191–192, August 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cotsonis:2006:SYS

[Cot06] George Cotsonis. Sharpening your SAS skills. The American Statistician, 60(4):346, November 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cowles:2003:BRB

[Cow03] Mary Kathryn Cowles. Book review: An R and S-PLUS Companion to Applied Regression. The American Statistician, 57(4):316– ??, November 2003. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://oberon.ingentaselect.
com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/00031305/v57n4/

s16/p316. Review of [Fox02].

Cowles:2004:BRBa

[Cow04a] Mary Kathryn Cowles. Book review: SAS for Linear Models (4th ed.). The American Statistician, 58(1):89–??, February 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v58n1/s21/p89;
 http://oberon.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/00031305/v58n1/s22/p89.

Cowles:2004:BRBb

[Cow04b] Mary Kathryn Cowles. Book review: Statistical Computing: An Introduction to Data Analysis using SPlus. The American Statistician, 58(1):89–??, February 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v58n1/s21/p89;
 http://oberon.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/00031305/v58n1/s22/p89.

Cowles:2004:RW

[Cow04c] Mary Kathryn Cowles. Review of WinBUGS 1.4. The American Statistician, 58(4):330–336, November 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/27643593.

Cowles:2006:PSC

[Cow06] Mary Kathryn Cowles. Probability and statistics for computer science. The American Statistician, 60(1):98, February 2006. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Chance:2001:TCS

[CR01] Beth L. Chance and Allan J. Rossman. Teacher’s corner: Sequencing topics in introductory statistics: a debate on what to teach when. The American Statistician, 55(2):140–144, May 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2686001.

Craigmile:2005:ASC

[Cra05] Peter F. Craigmile. All of statistics: a concise course in statistical inference. The American Statistician, 59(2):203–204, May 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Crawford:2008:UCG

[Cra08] Carol A. Gotway Crawford. Unlocking the census with GIS. The American Statistician, 62(4):363, November 2008. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cesana:2001:SPS

[CRM01] B. M. Cesana, G. Reina, and E. Marubini. Statistical practice: Sample size for testing a proportion in clinical trials: a “twostep” procedure combining power and confidence interval expected width. The American Statistician, 55(4):288–292, November 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685689.

Carlton:2005:MBF

[CS05] Matthew A. Carlton and William D. Stansfield. Making babies by the flip of a coin? The American Statistician, 59(2):180–182, May 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cook:2004:POD

[CW04] R. Dennis Cook and Sanford Weisberg. Partial one-dimensional regression models. The American Statistician, 58(2):110–116, May 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/27643518.

Dobler:2005:BRTb

[DAC05] Carolyn Pillers Dobler and Christine M. Anderson-Cook. Brief reviews of teaching materials. The American Statistician, 59(3):

278, August 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Dobler:2006:BRT

[DAC06] Carolyn Pillers Dobler and Christine M. Anderson-Cook. Brief reviews of teaching materials. The American Statistician, 60(2): 207, May 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

David:2005:TRN

[Dav05] Herbert A. David. Tables related to the normal distribution. The American Statistician, 59(4):309–311, November 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

David:2006:IMA

[Dav06] H. A. David. The introduction of matrix algebra into statistics. The American Statistician, 60(2):162, May 2006. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Davis:2007:CCS

[Dav07] J. Wade Davis. A crash course in SPSS for windows: Updated for versions 10, 11, 12 and 13. The American Statistician, 61(1): 99, February 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

David:2008:BRT

[Dav08a] Herbert A. David. The beginnings of randomization tests. The American Statistician, 62(1):70–72, February 2008. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Davis:2008:HUM

[Dav08b] J. Wade Davis. Handbook of univariate and multivariate data analysis and interpretation with SPSS. The American Statistician, 62(3):268, August 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Davis:2008:MST

[Dav08c] J. Wade Davis. Medical statistics: a textbook for the health sciences. The American Statistician, 62(4):362, November 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Davis:2008:MND

[Dav08d] J. Wade Davis. Multidimensional nonlinear descriptive analysis. The American Statistician, 62(3):271, August 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

David:2009:HNZ

[Dav09] Herbert A. David. A historical note on zero correlation and independence. The American Statistician, 63(2):185–186, May 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Dielman:2001:LE

[DBD^{+}01] Terry E. Dielman, Roger L. Berger, Jimmy Doi, Robert G. Lehr, Rand R. Wilcox, Paul Vos, Josef Schmee, and John Wendell. Letters to the Editor. The American Statistician, 55(4):373–376, November 2001. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685709.

DelCastillo:2001:TCC

[DC01] Enrique Del Castillo and Suntara Cahya. A tool for computing confidence regions on the stationary point of response surface. The American Statistician, 55(4):358–365, November 2001. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685701.

Das:2006:BAG

[DD06] Sourish Das and Dipak K. Dey. On Bayesian analysis of generalized linear models using the Jacobian technique. The American Statistician, 60(3):264–268, August 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

DeSantis:2006:PPT

[De 06] Fulvio De Santis. Power priors and their use in clinical trials. The American Statistician, 60(2):122–129, May 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

DePaula:2008:CMI

[De 08] Áureo De Paula. Conditional moments and independence. The American Statistician, 62(3):219–221, August 2008. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

deLaubenfels:2006:VLS

[deL06] Ralph deLaubenfels. The victory of least squares and orthogonality in statistics. The American Statistician, 60(4):315–321, November 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Deming:2006:ESE

[Dem06] W. Edwards Deming. On errors in surveys (an excerpt). The American Statistician, 60(1):34–38, February 2006. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Devore:2001:RBT

[Dev01] Jay Devore. Reviews of books and teaching materials: Stat Labs: Mathematical Statistics Through Applications, by Deborah Nolan; Terry Speed. The American Statistician, 55(3):257–258, August 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685813.

Devore:2002:BRB

[Dev02] Jay Devore. Book review: A First Course in Probability. The American Statistician, 56(3):247–??, August 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v56n3/s14/p247.

Devore:2006:EMS

[Dev06a] Jay Devore. Essential mathematics and statistics for science. The American Statistician, 60(4):344–345, November 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Devore:2006:PSS

[Dev06b] Jay Devore. Probability, statistics, and stochastic processes. The American Statistician, 60(3):290, August 2006. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Devore:2006:SBE

[Dev06c] Jay Devore. Statistics for business and economics. The American Statistician, 60(4):342–343, November 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Devore:2007:ARA

[Dev07a] Jay Devore. Applied regression analysis: a second course in business and economic statistics. The American Statistician, 61(4): 367, November 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Devore:2007:IBS

[Dev07b] Jay Devore. Introduction to business statistics. The American Statistician, 61(1):97, February 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Devore:2007:MSD

[Dev07c] Jay Devore. Making sense of data: a practical guide to exploratory data analysis and data mining. The American Statistician, 61(4):370–371, November 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Devore:2008:S

[Dev08] Jay Devore. Statistics. The American Statistician, 62(4):362– 363, November 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

David:2007:SMK

[DF07] Herbert A. David and Wayne A. Fuller. Sir Maurice Kendall (1907–1983). The American Statistician, 61(1):41–46, February 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Datta:2007:CFC

[DG07] Gauri Sankar Datta and Malay Ghosh. Characteristic functions without contour integration. The American Statistician, 61(1):67–70, February 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Danaher:2005:BYE

[DH05] Peter J. Danaher and Bruce G. S. Hardie. Bacon with your eggs? Applications of a new bivariate beta-binomial distribution. The American Statistician, 59(4):282–286, November 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Davi:2005:RTT

[DHN^{+}05] Angelique Davi, Dominique Haughton, Nada Nasr, Gaurav Shah, Maria Skaletsky, and Ruth Spack. A review of two text-mining packages. The American Statistician, 59(1):89–103, February 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Dias:2005:SAF

[Dia05] Alexandra Dias. Statistical analysis of financial data in S-Plus. The American Statistician, 59(4):351, November 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Dipietro:2004:BRR

[Dip04] Michele Dipietro. Bayesian randomized response as a class project. The American Statistician, 58(4):303–309, November 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/27643588.

Dixon:2007:LSB

[Dix07] Philip Dixon. The little SAS book for enterprise guide 3.0. The American Statistician, 61(3):277, August 2007. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Duchateau:2005:UHG

[DJ05] Luc Duchateau and Paul Janssen. Understanding heterogeneity in generalized mixed and frailty models. The American Statistician, 59(2):143–146, May 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Dodd:2007:BVS

[DK07] Lori E. Dodd and Edward L. Korn. The bootstrap variance of the square of a sample mean. The American Statistician, 61(2): 127–131, May 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Deng:2000:RNG

[DL00] Lih-Yuan Deng and Dennis K. J. Lin. Random number generation for the New Century. The American Statistician, 54(2):145–150, May 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.amstat.org/publications/
tas/Deng.htm;   http://www.jstor.org/stable/2686034.

Dobler:2004:BRBe

[DL04a] Carolyn Pillers Dobler and Russell V. Lenth. Book review: Common Statistical Methods for Clinical Research with SAS Examples, by Glenn A. Walker. The American Statistician, 58(2):175, May 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/27643539.

Dobler:2004:BRBg

[DL04b] Carolyn Pillers Dobler and Russell V. Lenth. Book review: Elementary Probability, by David Strizaker. The American Statistician, 58(2):175, May 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/27643541.

Dobler:2004:BRBc

[DL04c] Carolyn Pillers Dobler and Russell V. Lenth. Book review: Environmental Mathematics in the Classroom, by B. A. Fusaro; P. C. Kenschaft. The American Statistician, 58(2):175, May 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/27643537.

Dobler:2004:BRBd

[DL04d] Carolyn Pillers Dobler and Russell V. Lenth. Book review: International Mathematical Olympiads 1986–1999, by Marcin E. Kuczma; Mathematical Olympiads 2003–2001: Problems and Solutions From Around the World by Titu Andreescu; Zuming Feng; George Lee, Jr. The American Statistician, 58(2):175, May 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/27643538.

Dobler:2004:BRBb

[DL04e] Carolyn Pillers Dobler and Russell V. Lenth. Book review: Mathematical Diamonds, by Ross Honsberger. The American Statistician, 58(2):175, May 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/27643536.

Dobler:2004:BRBf

[DL04f] Carolyn Pillers Dobler and Russell V. Lenth. Book review: Nonparametric Statistical Inference, by Jean Dickinson Gibbons; Subhabrata Chakraborti. The American Statistician, 58(2):175, May 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/27643540.

Dobler:2004:BRT

[DL04g] Carolyn Pillers Dobler and Russell V. Lenth. Brief reviews of teaching materials. The American Statistician, 58(2):175, May 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Dobler:2005:BRTa

[DL05] Carolyn Pillers Dobler and Russell V. Lenth. Brief reviews of teaching materials. The American Statistician, 59(1):114–115, February 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Denby:2001:SPE

[DLM01] L. Denby, J. M. Landwehr, and C. L. Mallows. Statistical practice: An exercise in the real world of design and analysis. The American Statistician, 55(4):263–271, November 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685686.

Dawson:2001:NHL

[DM01] Don Dawson and Lonnie Magee. The National Hockey League Entry Draft, 1969–1995: An application of a weighted pooladjacent-violators algorithm. The American Statistician, 55(3): 194–199, August 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685798.

David:2006:SSW

[DM06] Herbert A. David and Donald F. Morrison. Samuel Stanley Wilks (1906–1964). The American Statistician, 60(1):46–49, February 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

deMast:2007:AKT

[dM07] Jeroen de Mast. Agreement and kappa-type indices. The American Statistician, 61(2):148–153, May 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

deMicheaux:2009:UCC

[dML09] Pierre Lafaye de Micheaux and Benoit Liquet. Understanding convergence concepts: a visual-minded and graphical simulationbased approach. The American Statistician, 63(2):173–178, May 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Dette:2004:EED

[DO04] Holger Dette and Timothy E. O’Brien. Efficient experimental design for the Behrens–Fisher problem with application to bioassay. The American Statistician, 58(2):138–143, May 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/27643522.

Dobler:2002:BRB

[Dob02] Carolyn Pillers Dobler. Book review: Mathematical Statistics: Basic Ideas and Selected Topics (vol. 1, 2nd ed.). The American Statistician, 56(4):332–??, November 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v56n4/s18/p332.

Dobler:2003:BRBb

[Dob03a] Carolyn Pillers Dobler. Book review: Problems in Probability. The American Statistician, 57(3):218–??, August 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v57n3/s17/p218.

Dobler:2003:BRBa

[Dob03b] Carolyn Pillers Dobler. Book review: The Practice of Statistics: Putting the Pieces Together. The American Statistician, 57(2): 142–??, May 2003. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://oberon.ingentaselect.
com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/00031305/v57n2/
s15/p142.

Dobler:2004:SPP

[Dob04a] Carolyn Pillers Dobler. 101 special practice problems in probability and statistics. The American Statistician, 58(2):173, May 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Dobler:2004:BRBa

[Dob04b] Carolyn Pillers Dobler. Book review: 101 Special Practice Problems in Probability and Statistics, by Paul D. Berger; Samuel C. Hanna; Robert E. Maurer. The American Statistician, 58(2):173, May 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/27643533.

Dobler:2006:EMT

[Dob06a] Carolyn Pillers Dobler. The essence of multivariate thinking: Basic themes and methods. The American Statistician, 60(3): 292, August 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Dobler:2006:SGR

[Dob06b] Carolyn Pillers Dobler. SAS guide to report writing: Examples. The American Statistician, 60(3):294, August 2006. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Dorsey-Palmateer:2004:BHH

[DPS04] Reid Dorsey-Palmateer and Gary Smith. Bowlers’ hot hands. The American Statistician, 58(1):38–??, February 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v58n1/s10/p38.

Dodge:2001:APC

[DR01] Yadolah Dodge and Valentin Rousson. On asymmetric properties of the correlation coefficient in the regression setting. The American Statistician, 55(1):51–54, February 2001. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.amstat.org/publications/tas/Dodge.htm;
 http://www.jstor.org/stable/2685529.

Dumbgen:2007:FAB

[DR07] Lutz Dümbgen and Hans Riedwyl. On fences and asymmetry in box-and-whiskers plots. The American Statistician, 61(4):356– 359, November 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Duckworth:2002:BTS

[DS02] William M. Duckworth and W. Robert Stephenson. Beyond traditional statistical methods. The American Statistician, 56(3):230– ??, August 2002. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://oberon.ingentaselect.
com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/00031305/v56n3/
s11/p230.

Datta:2008:GPS

[DS08] Gauri S. Datta and Sanat K. Sarkar. A general proof of some known results of independence between two statistics. The American Statistician, 62(2):141–143, May 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Doane:2000:UBF

[DT00] David P. Doane and Ronald L. Tracy. Using beam and fulcrum displays to explore data. The American Statistician, 54(4):289–290, November 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:/

/www.amstat.org/publications/tas/Doane.htm;   http://
www.jstor.org/stable/2685780.

Dunn:2003:WHW

[Dun03] Peter K. Dunn. What happens when a 1 × 1 × r die is rolled? The American Statistician, 57(4):258–??, November 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v57n4/s8/p258.

Duncan:2004:BRB

[Dun04] Kristin A. Duncan. Book review: Grade Inflation: A Crisis in College Education, by Valen E. Johnson. The American Statistician, 58(3):259, August 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/27643568.

Easterling:2004:TED

[Eas04] Robert G. Easterling. Teaching experimental design. The American Statistician, 58(3):244–252, August 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/27643565.

Edwards:2007:ESB

[Edw07] Phillip Edwards. Essentials of statistics for business and economics. The American Statistician, 61(2):182–183, May 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Erosheva:2002:CDR

[EKG02] Elena A. Erosheva, Patricia D. Kroboth, and Joel B. Greenhouse. Characterizing the diurnal rhythm of DHEA. The American Statistician, 56(4):273–??, November 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v56n4/s3/p273.

Elian:2000:SFB

[Eli00] Silvia N. Elian. Simple forms of the best linear unbiased predictor in the general linear regression model. The American Statistician, 54(1):25–28, February 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685606.

Emerson:2009:AJB

[ESL09] John W. Emerson, Miki Seltzer, and David Lin. Assessing judging bias: An example from the 2000 Olympic Games. The American Statistician, 63(2):124–131, May 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Eubank:2002:EBC

[EW02] R. L. Eubank and Suojin Wang. The equivalence between the Cholesky decomposition and the Kalman filter. The American Statistician, 56(1):39–43, February 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/3087325.

Fay:2002:MBR

[Fay02] Michael P. Fay. Measuring a binary Response’s range of influence in logistic regression. The American Statistician, 56(1):5– 9, February 2002. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
3087321.

Froelich:2009:DYI

[FDC09] Amy G. Froelich, William M. Duckworth, and Jessica Culhane. Does your iPod really play favorites? The American Statistician, 63(3):263–268, August 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Froelich:2005:TST

[FDS05] Amy G. Froelich, William M. Duckworth, and W. Robert Stephenson. Training statistics teachers at Iowa State University. The American Statistician, 59(1):8–10, February 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Fellingham:2008:STS

[Fel08] Gilbert W. Fellingham. Statistical thinking in sports. The American Statistician, 62(4):360–361, November 2008. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Ferketich:2004:BRB

[Fer04] Amy K. Ferketich. Book review: Introductory Biostatistics for the Health Sciences: Modern Applications Including Bootstrap, by Michael R. Chernick; Robert H. Friis. The American Statistician, 58(2):172, May 2004. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
27643531.

Fewster:2009:SEB

[Few09] R. M. Fewster. A simple explanation of Benford’s Law. The American Statistician, 63(1):26–32, February 2009. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://pubs.amstat.org/doi/pdfplus/10.1198/tast.2009.
0005.

Fay:2002:DMC

[FF02] Michael P. Fay and Dean A. Follmann. Designing Monte Carlo implementations of permutation or bootstrap hypothesis tests. The American Statistician, 56(1):63–70, February 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/3087329.

Freidlin:2000:SMT

[FG00] Boris Freidlin and Joseph L. Gastwirth. Should the median test be retired from general use? The American Statistician, 54(3):161–164, August 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.amstat.org/publications/tas/Freidlin.htm;   http:
//www.jstor.org/stable/2685584.

Finney:2006:CGC

[Fin06] David J. Finney. Calibration guidelines challenge outlier practices. The American Statistician, 60(4):309–314, November 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Friendly:2009:WWV

[FK09] Michael Friendly and Ernest Kwan. Where’s Waldo? visualizing collinearity diagnostics. The American Statistician, 63(1):56–65, February 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Fong:2003:UST

[FKLL03] Daniel Y. T. Fong, C. W. Kwan, K. F. Lam, and Karen S. L. Lam. Use of the sign test for the median in the presence of ties. The American Statistician, 57(4):237–??, November 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v57n4/s3/p237.

Fitzmaurice:2007:AMR

[FLP07] Garrett M. Fitzmaurice, Stuart R. Lipsitz, and Michael Parzen. Approximate median regression via the Box–Cox transformation. The American Statistician, 61(3):233–238, August 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Frey:2008:SMI

[FM08] Jesse Frey and Osvaldo Marrero. A surprising MLE for intervalcensored binomial data. The American Statistician, 62(2):135– 137, May 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Faes:2009:ESS

[FMA^{+}09] Christel Faes, Geert Molenberghs, Marc Aerts, Geert Verbeke, and Michael G. Kenward. The effective sample size and an alternative small-sample degrees-of-freedom method. The American Statistician, 63(4):389–399, November 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Ford:2005:PES

[For05] E. David Ford. A primer of ecological statistics. The American Statistician, 59(4):350, November 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Foster:2008:IB

[Fos08] Eric D. Foster. Introduction to bioinformatics. The American Statistician, 62(3):268–269, August 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Fotopoulos:2002:BRB

[Fot02] Stergios B. Fotopoulos. Book review: Applied Probability Models. The American Statistician, 56(3):248–??, August 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v56n3/s15/p248.

Fox:2002:RPC

[Fox02] John Fox. An R and S-Plus companion to applied regression. Sage Publications, Thousand Oaks, CA, USA, 2002. ISBN 0-7619-2279-2, 0-7619-2280-6 (paperback). xvi + 312 pp. LCCN QA278.2 .F628 2002.

Fisher:2004:RWM

[FR04] Lynn A. Fisher and Donald St. P. Richards. Random walks as motivational material in introductory statistics and probability courses. The American Statistician, 58(4):310–316, November 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/27643589.

Franklin:2000:BRB

[Fra00a] Christine Franklin. Book review: Intermediate Statistics: A Modern Approach (2nd ed.). The American Statistician, 54(3): 225, August 2000. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685595.

Franklin:2000:BRBb

[Fra00b] Christine Franklin. Book review: Intermediate Statistics: A Modern Approach (2nd ed.). The American Statistician, 54(3):225–??, August 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Finner:2007:CDC

[FRD07] Helmut Finner, Markus Roters, and Thorsten Dickhaus. Characterizing density crossing points. The American Statistician, 61 (1):28–33, February 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Frey:2005:RMB

[Fre05] Jesse Frey. A ranking method based on minimizing the number of in-sample errors. The American Statistician, 59(3):207–216, August 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Freedman:2006:CHS

[Fre06] David A. Freedman. On the so-called “Huber Sandwich Estimator” and “Robust Standard Errors”. The American Statistician, 60(4):299–302, November 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Freedman:2007:HCS

[Fre07a] David A. Freedman. How can the score test be inconsistent? The American Statistician, 61(4):291–295, November 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Frey:2007:BIP

[Fre07b] Jesse Frey. Bayesian inference on a proportion believed to be a simple fraction. The American Statistician, 61(3):201–206, August 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Frey:2007:HBT

[Fre07c] Jesse Frey. Is an .833 hitter better than a .338 hitter? The American Statistician, 61(2):105–111, May 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Freedman:2008:SA

[Fre08] David A. Freedman. Survival analysis. The American Statistician, 62(2):110–119, May 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Friedman:2002:SGI

[Fri02a] Lynn Friedman. Shieh, G. (2001), “The Inequality between the Coefficient of Determination and the Sum of Squared Simple Correlation Coefficients,” The American Statistician, 55, 121-124: Comment by Friedman. The American Statistician, 56(1): 82, February 2002. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
3087341.

Friendly:2002:CED

[Fri02b] Michael Friendly. Corrgrams: Exploratory displays for correlation matrices. The American Statistician, 56(4):316–??, November 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://oberon.ingentaselect.com/
cgi-bin/linker?ini=asa&reqidx=/cw/asa/00031305/v56n4/
s10/p316.

Friedman:2015:NCC

[Fri15] Emil M. Friedman. Nontransitivity, correlation, and causation: Langford, E., Schwertman, N., and Owens, M. (2001), “Is the Property of Being Positively Correlated Transitive?” The American Statistician, 55, 322–325: Comment by Friedman. The American Statistician, 69(3):257, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/00031305.2015.1056382.

See [LSO01].

Fisher:2001:GAD

[FS01] N. I. Fisher and P. Switzer. Graphical assessment of dependence: Is a picture worth 100 tests? The American Statistician, 55 (3):233–239, August 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685807.

Foster:2006:BWB

[FS06] Dean P. Foster and Robert A. Stine. Being Warren Buffett. The American Statistician, 60(1):53–60, February 2006. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Friedman:2005:GVS

[FW05] Lynn Friedman and Melanie Wall. Graphical views of suppression and multicollinearity in multiple linear regression. The American Statistician, 59(2):127–136, May 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Flury:2000:EE

[FZ00] Bernard Flury and Alice Zoppe. Exercises in EM. The American Statistician, 54(3):207–209, August 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.amstat.org/publications/tas/Flury.htm;
 http://www.jstor.org/stable/2685591.

Gadbury:2001:RIB

[Gad01] Gary L. Gadbury. Randomization inference and bias of standard errors. The American Statistician, 55(4):310–313, November 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685693.

Gal:2003:TSL

[Gal03] Iddo Gal. Teaching for statistical literacy and services of statistics agencies. The American Statistician, 57(2):80–??, May 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v57n2/s3/p80.

Gambino:2009:DEC

[Gam09] Jack G. Gambino. Design effect caveats. The American Statistician, 63(2):141–146, May 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Genest:2003:DDK

[GB03] Christian Genest and Jean-Claude Boies. Detecting dependence with Kendall plots. The American Statistician, 57(4):275–??, November 2003. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://oberon.ingentaselect.
com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/00031305/v57n4/
s11/p275.

Gullion:2006:WSC

[GB06] Christina M. Gullion and Nancy Berman. What statistical consultants do. The American Statistician, 60(2):130–138, May 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Guo:2004:SJM

[GC04] Xu Guo and Bradley P. Carlin. Separate and joint modeling of longitudinal and event time data using standard computer packages. The American Statistician, 58(1):16–??, February 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v58n1/s7/p16.

Grewal:2008:URG

[GDL08] Rajdeep Grewal, James A. Dearden, and Gary L. Llilien. The university rankings game. The American Statistician, 62(3):232– 237, August 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Glen:2001:APP

[GEL01] Andrew G. Glen, Diane L. Evans, and Lawrence M. Leemis. APPL: a probability programming language. The American Statistician, 55(2):156–166, May 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2686005.

Gelfand:2004:DSR

[Gel04] Alan E. Gelfand. Discussion of “Statistical Research: Some Advice for Beginners” by M. Hamada and R. Sitter. The American Statistician, 58(3):197–198, August 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/27643553. See [HS04b].

Gelman:2005:CTS

[Gel05] Andrew Gelman. A course on teaching statistics at the university level. The American Statistician, 59(1):4–7, February 2005. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Gelman:2006:BWC

[Gel06] Andrew Gelman. The boxer, the wrestler, and the coin flip. The American Statistician, 60(2):146–150, May 2006. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Gelman:2008:TBG

[Gel08] Andrew Gelman. Teaching Bayes to graduate students in political science, sociology, public health, education, economics, …. The American Statistician, 62(3):202–205, August 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Gentleman:2003:BRB

[Gen03] Robert Gentleman. Book review: Resampling Methods: A Practical Guide to Data Analysis (2nd ed.). The American Statistician, 57(1):70–??, February 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://oberon.

ingentaselect.com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/
00031305/v57n1/s18/p70.

Gentle:2004:CSC

[Gen04] James E. Gentle. Courses in statistical computing and computational statistics. The American Statistician, 58(1):2–??, February 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://oberon.ingentaselect.com/
cgi-bin/linker?ini=asa&reqidx=/cw/asa/00031305/v58n1/
s2/p2.

Gartside:2001:LE

[GGF^{+}01] Peter S. Gartside, Joseph L. Gastwirth, Boris Freidlin, W. Fred Webber, Walter A. Rosenkrantz, Sander Greenland, Alan Agresti, Herman Callaert, Reinhard Bergmann, and I. J. Good. Letters to the Editor. The American Statistician, 55(2):171–174, May 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2686010.

Giron:2005:BAM

[GGR05] Javier Girón, Josep Ginebra, and Alex Riba. Bayesian analysis of a multinomial sequence and homogeneity of literary style. The American Statistician, 59(1):19–30, February 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Ghosh:2002:PIR

[Gho02] B. K. Ghosh. Probability inequalities related to Markov’s theorem. The American Statistician, 56(3):186–??, August 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v56n3/s4/p186.

Gill:2000:LGR

[Gil00] Paramjit S. Gill. Late-game reversals in professional basketball, football, and hockey. The American Statistician, 54(2):94–99, May 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.amstat.org/publications/
tas/Gill.htm;   http://www.jstor.org/stable/2686024.

Gilbert:2008:PGE

[Gil08] Gregory E. Gilbert. A pocket guide to epidemiology. The American Statistician, 62(1):92–93, February 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Gonen:2005:BTS

[GJLW05] Mithat Gönen, Wesley O. Johnson, Yonggang Lu, and Peter H. Westfall. The Bayesian two-sample t test. The American Statistician, 59(3):252–257, August 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Ganju:2000:AUD

[GL00] Jitendra Ganju and James M. Lucas. Analysis of unbalanced data from an experiment with random block effects and unequally spaced factor levels. The American Statistician, 54(1): 5–11, February 2000. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685603.

Glass:2004:BRB

[Gla04] D. Glass. Book review: Proofs That Really Count: The Art of Combinatorial Proof, by Arthur T. Benjamin; Jennifer J. Quinn. The American Statistician, 58(4):360, November 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/27643599.

Gelman:2002:YCL

[GN02] Andrew Gelman and Deborah Nolan. You can load a die, but you can’t bias a coin. The American Statistician, 56(4):308–??, November 2002. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://oberon.ingentaselect.
com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/00031305/v56n4/
s8/p308.

Gonzalez:2003:BRB

[Gon03] Richard Gonzalez. Book review: Applied Multivariate Statistics for the Social Sciences (4th ed.). The American Statistician, 57(1):68–??, February 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://oberon.

ingentaselect.com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/
00031305/v57n1/s15/p68.

Gelman:2009:SPU

[GP09] Andrew Gelman and David K. Park. Splitting a predictor at the upper quarter or third and the lower quarter or third. The American Statistician, 63(1):1–8, February 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Gelman:2002:LPW

[GPD02] Andrew Gelman, Cristian Pasarica, and Rahul Dodhia. Let’s practice what we preach: Turning tables into graphs. The American Statistician, 56(2):121–??, May 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://oberon.ingentaselect.com/cgi-bin/linker?
ini=asa&reqidx=/cw/asa/00031305/v56n2/s8/p121.

Gervini:2004:CED

[GR04] Daniel Gervini and Valentin Rousson. Criteria for evaluating dimension-reducing components for multivariate data. The American Statistician, 58(1):72–??, February 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v58n1/s14/p72.

Gibbs:2009:CWS

[GR09] Alison Gibbs and Nancy Reid. Comment: What is statistics? The American Statistician, 63(2):112–113, May 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Grambsch:2008:RMM

[Gra08] Patricia Grambsch. Regression to the mean, murder rates, and shall-issue laws. The American Statistician, 62(4):289–295, November 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Gromping:2007:ERI

[Grö07] Ulrike Grömping. Estimators of relative importance in linear regression based on variance decomposition. The American Statistician, 61(2):139–147, May 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Gromping:2009:VIA

[Grö09] Ulrike Grömping. Variable importance assessment in regression: Linear regression versus random forest. The American Statistician, 63(4):308–319, November 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Ghosh:2002:SDW

[GS02] Malay Ghosh and Bimal K. Sinha. A simple derivation of the Wishart distribution. The American Statistician, 56(2):100–??, May 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://oberon.ingentaselect.com/
cgi-bin/linker?ini=asa&reqidx=/cw/asa/00031305/v56n2/
s4/p100.

Gelman:2006:DBS

[GS06] Andrew Gelman and Hal Stern. The difference between “significant” and “not significant” is not itself statistically significant. The American Statistician, 60(4):328–331, November 2006. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Gajewski:2008:OHT

[GS08] Byron J. Gajewski and Stephen D. Simon. A one-hour training seminar on Bayesian statistics for nursing graduate students. The American Statistician, 62(3):190–194, August 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Guerra:2006:DDC

[Gue06] Rudy Guerra. Dicing with death: Chance, risk, and health. The American Statistician, 60(2):204, May 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Gupta:2002:BRB

[Gup02] Ghanshyam Gupta. Book review: A Handbook of Statistical Analyses using SAS, 2nd ed. The American Statistician, 56(4):334– ??, November 2002. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://oberon.ingentaselect.
com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/00031305/v56n4/
s22/p334.

Gurka:2006:SBL

[Gur06] Matthew J. Gurka. Selecting the best linear mixed model under REML. The American Statistician, 60(1):19–26, February 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Gurwit:2008:LSE

[Gur08] Margie R. Gurwit. Learning SAS by example: a programmer’s guide. The American Statistician, 62(2):182–183, May 2008. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Graziani:2009:HCM

[GV09] Rebecca Graziani and Piero Veronese. How to compute a mean? the Chisini approach and its applications. The American Statistician, 63(1):33–36, February 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hagwood:2009:ARC

[Hag09] Charles Hagwood. An application of the residue calculus: The distribution of the sum of nonhomogeneous gamma variates. The American Statistician, 63(1):37–39, February 2009. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hahn:2002:DPS

[Hah02] Gerald J. Hahn. Deming and the proactive statistician. The American Statistician, 56(4):290–??, November 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v56n4/s5/p290.

Haigh:2001:RBT

[Hai01] John Haigh. Reviews of books and teaching materials: Mathematics of Chance. The American Statistician, 55(4):372, November 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685707.

Haigh:2003:BRB

[Hai03] John Haigh. Book review: Risk-Benefit Analysis (2nd ed.). The American Statistician, 57(1):69–??, February 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v57n1/s16/p69.

Haigh:2007:PMA

[Hai07] John Haigh. Probability: Modeling and applications to random processes. The American Statistician, 61(3):276–277, August 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hall:2008:ARO

[Hal08] Daniel B. Hall. Analyzing receiver operating characteristic curves with SAS. The American Statistician, 62(4):362, November 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hamer:2000:BRB

[Ham00] Robert M. Hamer. Book review: Statistics in Psychiatry. The American Statistician, 54(4):322–323, November 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685788.

Hanley:2004:TWM

[Han04] James A. Hanley. “transmuting” women into men: Galton’s family data on human stature. The American Statistician, 58(3):237– 243, August 2004. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
27643564.

Hansen:2005:USW

[Han05] John. Hansen. Using SPSS for Windows and Macintosh: Analyzing and understanding data. The American Statistician, 59(1): 113, February 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hardin:2000:BRB

[Har00] James W. Hardin. Book review: LogXact 4.1 for Windows. The American Statistician, 54(4):320–321, November 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685786.

Harville:2004:DSC

[Har04] David A. Harville. Discussion: Statistics and the college football championship. The American Statistician, 58(3):187–189, August 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hardin:2006:IST

[Har06] Jo Hardin. Introduction to statistics through resampling methods and R/S-PLUS. The American Statistician, 60(4):343–344, November 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Harvill:2008:FPC

[Har08] Jane L. Harvill. Fundamental probability: a computational approach. The American Statistician, 62(2):179–180, May 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hazelton:2003:GTA

[Haz03] Martin L. Hazelton. A graphical tool for assessing normality. The American Statistician, 57(4):285–??, November 2003.

CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v57n4/s12/p285.

Hazelton:2004:CGT

[Haz04] Martin L. Hazelton. Corrections: a graphical tool for assessing normality. The American Statistician, 58(2):178, May 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/27643545.

Hubbard:2003:CME

[HB03] Raymond Hubbard and M. J. Bayarri. Confusion over measures of evidence (p’s) versus errors (α’s) in classical statistical testing. The American Statistician, 57(3):171–??, August 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v57n3/s5/p171.

Harris:2005:MRI

[HB05a] Ian R. Harris and Brent D. Burch. Measuring relative importance of sources of variation without using variance. The American Statistician, 59(3):217–222, August 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hubbard:2005:LE

[HB05b] Raymond. Hubbard and M. J. Bayarri. Letters to the Editor. The American Statistician, 59(4):353–354, November 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Horton:2004:URT

[HBQ04] Nicholas J. Horton, Elizabeth R. Brown, and Linjuan Qian. Use of R as a toolbox for mathematical statistics exploration. The American Statistician, 58(4):343–357, November 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/27643595.

Hin:2007:CWC

[HCW07] Lin-Yee Hin, Vincent J. Carey, and You-Gan Wang. Criteria for working-correlation-structure selection in GEE. The American Statistician, 61(4):360–364, November 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Haughton:2003:RSP

[HDE^{+}03] Dominique Haughton, Joel Deichmann, Abdolreza Eshghi, Selin Sayek, Nicholas Teebagy, and Heikki Topi. A review of software packages for data mining. The American Statistician, 57(4):290–??, November 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://oberon.ingentaselect.com/cgi-bin/linker?
ini=asa&reqidx=/cw/asa/00031305/v57n4/s14/p290.

Hatzinikolaou:2004:RSS

[HF04] Dimitris Hatzinikolaou and Kosmas Ferentinos. On the role of semielasticity in statistics. The American Statistician, 58(2):150– 151, May 2004. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
27643525.

Heller:2008:BD

[HFDJ08] Gillian Z. Heller, Andrew B. Forbes, Keith B. G. Dear, and Erica Jobling. Biostatistics @ distance. The American Statistician, 62 (4):321–328, November 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilton:2001:TCS

[HGA01] Sterling C. Hilton, Scott D. Grimshaw, and Genan T. Anderson. Teacher’s corner: Statistics in preschool. The American Statistician, 55(4):332–336, November 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2685697.

Hoenig:2001:APP

[HH01] John M. Hoenig and Dennis M. Heisey. The abuse of power: The pervasive fallacy of power calculations for data analysis. The American Statistician, 55(1):19–24, February 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.amstat.org/publications/tas/
Hoenig.htm;   http://www.jstor.org/stable/2685525.

Hansen:2004:PNS

[HH04] Morris H. Hansen and William N. Hurwitz. The problem of nonresponse in sample surveys. The American Statistician, 58(4): 292–294, November 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/27643585.

Hardin:2006:RSW

[HH06] James W. Hardin and Joseph M. Hilbe. A review of Scientific WorkPlace 5.5. The American Statistician, 60(1):93–96, February 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hodge:2002:WMM

[HHV^{+}02] Susan E. Hodge, Jian Huang, Paul Vos, Knut M. Wittkowski, and Melanie M. Wall. Wall, M. M., Boen, J., and Tweedie, R. (2001), “An Effective Confidence Interval for the Mean with Samples of Size One and Two,” The American Statistician, 55, 102-105. Comments by Hodge and Huang, Vos, and Wittkowski and reply. The American Statistician, 56(1):79–80, February 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/3087337.

Hothorn:2006:LSC

[HHvdWZ06] Torsten Hothorn, Kurt Hornik, Mark A. van de Wiel, and Achim Zeileis. A Lego system for conditional inference. The American Statistician, 60(3):257–263, August 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2000:SENa

[Hil00a] Joseph Hilbe. Section Editor’s notes. The American Statistician, 54(1):71, February 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685615.

Hilbe:2000:SENb

[Hil00b] Joseph Hilbe. Section Editor’s notes. The American Statistician, 54(4):319, November 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685785.

Hilbe:2001:BRB

[Hil01] Joseph Hilbe. Reviews of books and teaching materials: XploRe: Learning Guide. The American Statistician, 55(2):169–170, May 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2686009.

Hilbe:2002:SENb

[Hil02a] Joseph Hilbe. Section Editor’s notes. The American Statistician, 56(2):148–??, May 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://oberon.

ingentaselect.com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/
00031305/v56n2/s11/p148.

Hilbe:2002:SENc

[Hil02b] Joseph Hilbe. Section Editor’s notes. The American Statistician, 56(3):234–??, August 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://oberon.

ingentaselect.com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/
00031305/v56n3/s12/p234.

Hilbe:2002:SCSa

[Hil02c] Joseph Hilbe. Statistical computing software reviews: Section Editor’s notes. The American Statistician, 56(1):71, February 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/3087330.

Hilbe:2002:SCSb

[Hil02d] Joseph Hilbe. Statistical computing software reviews: Section Editor’s notes. The American Statistician, 56(4):325–??, November 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://oberon.ingentaselect.com/
cgi-bin/linker?ini=asa&reqidx=/cw/asa/00031305/v56n4/
s11/p325.

Hilbe:2002:XEP

[Hil02e] Joseph M. Hilbe. XPro 5.1: Exact procedures for parametric statistical inference. The American Statistician, 56(4):326–??, November 2002. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://oberon.ingentaselect.
com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/00031305/v56n4/
s12/p326.

Hilbe:2003:SENa

[Hil03a] Joseph Hilbe. Section Editor’s notes. The American Statistician, 57(3):200–??, August 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://oberon.

ingentaselect.com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/
00031305/v57n3/s13/p200.

Hilbe:2003:SENb

[Hil03b] Joseph Hilbe. Section Editor’s notes. The American Statistician, 57(4):289–??, November 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://oberon.

ingentaselect.com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/
00031305/v57n4/s13/p289.

Hilbe:2003:RCS

[Hil03c] Joseph M. Hilbe. A review of current SPSS products: SPSS 12, SigmaPlot 8.02, SigmaStat 3.0, Part 1. The American Statistician, 57(4):310–??, November 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://oberon.ingentaselect.com/cgi-bin/linker?
ini=asa&reqidx=/cw/asa/00031305/v57n4/s15/p310.

Hilbe:2004:SENa

[Hil04a] Joseph Hilbe. Section Editor’s notes. The American Statistician, 58(1):85–??, February 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://oberon.

ingentaselect.com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/
00031305/v58n1/s16/p85.

Hilbe:2004:RSP

[Hil04b] Joseph M. Hilbe. A review of SPSS 12.01, part 2. The American Statistician, 58(2):168–171, May 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/27643530.

Hilbe:2004:SENb

[Hil04c] Joseph M. Hilbe. Section Editor’s notes. The American Statistician, 58(2):159, May 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/27643527.

Hilbe:2004:SENc

[Hil04d] Joseph M. Hilbe. Section Editor’s notes. The American Statistician, 58(3):253, August 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/27643566.

Hilbe:2004:SENd

[Hil04e] Joseph M. Hilbe. Section Editor’s notes. The American Statistician, 58(4):328–329, November 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/27643592.

Hilbe:2004:SCS

[Hil04f] Joseph M. Hilbe. Statistical computing software reviews. The American Statistician, 58(2):92, May 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2005:RSb

[Hil05a] Joseph M. Hilbe. Review of SigmaPlot 9.0. The American Statistician, 59(1):111–112, February 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2005:RSP

[Hil05b] Joseph M. Hilbe. A review of SPSS, part 3. The American Statistician, 59(2):185–186, May 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2005:RSc

[Hil05c] Joseph M. Hilbe. A review of Stata 9.0. The American Statistician, 59(4):335–348, November 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2005:RSa

[Hil05d] Joseph M. Hilbe. A review of Systat 11. The American Statistician, 59(1):104–110, February 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2005:SENa

[Hil05e] Joseph M. Hilbe. Section Editor’s notes. The American Statistician, 59(1):88, February 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2005:SENb

[Hil05f] Joseph M. Hilbe. Section Editor’s notes. The American Statistician, 59(2):183–184, May 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2005:SENc

[Hil05g] Joseph M. Hilbe. Section Editor’s notes. The American Statistician, 59(3):273, August 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2005:SENd

[Hil05h] Joseph M. Hilbe. Section Editor’s notes. The American Statistician, 59(4):334, November 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2005:S

[Hil05i] Joseph M. Hilbe. SigmaStat 3.1. The American Statistician, 59(2):187–191, May 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2006:M

[Hil06a] Joseph M. Hilbe. Mathematica 5.2. The American Statistician, 60(2):176–186, May 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2006:RLN

[Hil06b] Joseph M. Hilbe. A review of LIMDEP 9.0 and NLOGIT 4.0. The American Statistician, 60(2):187–202, May 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2006:SENa

[Hil06c] Joseph M. Hilbe. Section Editor’s notes. The American Statistician, 60(1):92, February 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2006:SENb

[Hil06d] Joseph M. Hilbe. Section Editor’s notes. The American Statistician, 60(2):175, May 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2006:SENc

[Hil06e] Joseph M. Hilbe. Section Editor’s notes. The American Statistician, 60(3):271, August 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2006:SENd

[Hil06f] Joseph M. Hilbe. Section Editor’s notes. The American Statistician, 60(4):340, November 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2007:EPC

[Hil07a] Joseph M. Hilbe. ePrint 5 professional conversion utility. The American Statistician, 61(2):179–180, May 2007. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2007:G

[Hil07b] Joseph M. Hilbe. GenStat 9. The American Statistician, 61 (3):269–273, August 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2007:SENa

[Hil07c] Joseph M. Hilbe. Section Editor’s notes. The American Statistician, 61(2):178, May 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2007:SENb

[Hil07d] Joseph M. Hilbe. Section Editor’s notes. The American Statistician, 61(3):260–261, August 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2007:SENc

[Hil07e] Joseph M. Hilbe. Section Editor’s notes. The American Statistician, 61(4):365–366, November 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2007:S

[Hil07f] Joseph M. Hilbe. STATISTICA 7. The American Statistician, 61(1):91–94, February 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2007:SCS

[Hil07g] Joseph M. Hilbe. Statistical computing software reviews. The American Statistician, 61(1):78, February 2007. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2008:HLS

[Hil08a] Joseph M. Hilbe. Hawkes learning systems. The American Statistician, 62(1):85–89, February 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2008:SENd

[Hil08b] Joseph M. Hilbe. Section Editor’s note. The American Statistician, 62(4):355, November 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2008:SENa

[Hil08c] Joseph M. Hilbe. Section Editor’s notes. The American Statistician, 62(1):73, February 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2008:SENb

[Hil08d] Joseph M. Hilbe. Section Editor’s notes. The American Statistician, 62(2):171–172, May 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2008:SENc

[Hil08e] Joseph M. Hilbe. Section Editor’s notes. The American Statistician, 62(3):267, August 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2008:S

[Hil08f] Joseph M. Hilbe. Systat 12.2. The American Statistician, 62(2): 177–178, May 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2009:SEN

[Hil09] Joseph M. Hilbe. Section Editor’s notes. The American Statistician, 63(1):75–77, February 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hitchcock:2003:HMH

[Hit03] David B. Hitchcock. A history of the Metropolis–Hastings algorithm. The American Statistician, 57(4):254–257, November 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://oberon.ingentaselect.com/
cgi-bin/linker?ini=asa&reqidx=/cw/asa/00031305/v57n4/
s7/p254;   http://www.jstor.org/stable/30037292.

Hanley:2008:SWI

[HJM08] James A. Hanley, Marilyse Julien, and Erica E. M. Moodie. Student’s z, t, and s: What if Gosset had R? The American Statistician, 62(1):64–69, February 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hanley:2001:TCV

[HJP^{+}01] James A. Hanley, Lawrence Joseph, Robert W. Platt, Moo K. Chung, and Patrick Belisle. Teacher’s corner: Visualizing the median as the minimum-deviation location. The American Statistician, 55(2):150–152, May 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2686003.

Hyde:2006:ICO

[HJS06] Valerie Hyde, Wolfgang Jank, and Galit Shmueli. Investigating concurrency in online auctions through visualization. The American Statistician, 60(3):241–250, August 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Horton:2007:MAA

[HK07] Nicholas J. Horton and Ken P. Kleinman. Much ado about nothing. The American Statistician, 61(1):79–90, February 2007. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Haughton:2006:RTD

[HKS06] Dominique Haughton, Arnold Kamis, and Patrick Scholten. A review of three directed acyclic graphs software packages. The American Statistician, 60(3):272–286, August 2006. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Horton:2001:SCS

[HL01] Nicholas J. Horton and Stuart R. Lipsitz. Statistical computing software reviews: Multiple imputation in practice: Comparison of software packages for regression models with missing variables. The American Statistician, 55(3):244–254, August 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685809.

Hunter:2004:TMA

[HL04] David R. Hunter and Kenneth Lange. A tutorial on MM algorithms. The American Statistician, 58(1):30–??, February 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v58n1/s9/p30.

Horton:2003:PBW

[HLP03] Nicholas J. Horton, Stuart R. Lipsitz, and Michael Parzen. A potential for bias when rounding in multiple imputation. The American Statistician, 57(4):229–??, November 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v57n4/s1/p229.

Haughton:2009:RTL

[HLW09] Dominique Haughton, Pascal Legrand, and Sam Woolford. Review of three latent class cluster analysis packages: Latent Gold, poLCA, and MCLUST. The American Statistician, 63(1):81–91, February 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Henderson:2001:TCE

[HM01] Michael Henderson and Mary C. Meyer. Teacher’s corner: Exploring the confidence interval for a binomial parameter in a first course in statistical computing. The American Statistician, 55(4): 337–344, November 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685698.

Hamada:2001:SPF

[HMRW01] M. Hamada, H. F. Martz, C. S. Reese, and A. G. Wilson. Statistical practice: Finding near-optimal Bayesian experimental designs via genetic algorithms. The American Statistician, 55(3):175– 181, August 2001. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685795.

Holt:2007:OSO

[Hol07] D. Tim Holt. The official statistics Olympic challenge. The American Statistician, 61(1):1–8, February 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Horton:2006:MLM

[Hor06] Nicholas J. Horton. Multilevel and longitudinal modeling using stata. The American Statistician, 60(3):293–294, August 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Ho:2008:IPI

[HQ08] Daniel E. Ho and Kevin M. Quinn. Improving the presentation and interpretation of online ratings data with model-based figures. The American Statistician, 62(4):279–288, November 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Holcomb:2000:UTL

[HR00] John P. Holcomb, Jr. and Rochelle L. Ruffer. Using a term-long project sequence in introductory statistics. The American Statistician, 54(1):49–53, February 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685611.

Harkness:2005:TGS

[HR05] W. L. Harkness and J. L. Rosenberger. Training graduate students at Penn State University in teaching statistics. The American Statistician, 59(1):11–13, February 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

He:2006:TDM

[HR06] Yulei He and Trivellore E. Raghunathan. Tukey’s gh distribution for multiple imputation. The American Statistician, 60(3):251– 256, August 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hedayat:2003:SDE

[HS03] A. S. Hedayat and Bikas K. Sinha. On a sampling design for estimation of negligible accident rates involving electronic toys. The American Statistician, 57(4):249–??, November 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v57n4/s5/p249.

Hamada:2004:R

[HS04a] Michael Hamada and Randy Sitter. Response. The American Statistician, 58(3):202, August 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hamada:2004:SRS

[HS04b] Michael Hamada and Randy Sitter. Statistical research: Some advice for beginners. The American Statistician, 58(2):93–101, May 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/27643516.

See discussion [Gel04].

Hedayat:2009:CWS

[HS09] Samad Hedayat and John Stufken. Comment: What is statistics? The American Statistician, 63(2):115–116, May 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Heiberger:2002:DDC

[HT02] Richard M. Heiberger and Paulo Teles. Displays for direct comparison of ARIMA models. The American Statistician, 56(2):131– ??, May 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://oberon.ingentaselect.com/
cgi-bin/linker?ini=asa&reqidx=/cw/asa/00031305/v56n2/
s9/p131.

Hanley:2006:PFR

[HT06] James A. Hanley and Dana Teltsch. The PDF of a function of a random variable. The American Statistician, 60(1):61–67, February 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hu:2003:BRB

[Hu03] X. Joan Hu. Book review: Generalized Linear Models: With Applications in Engineering and the Sciences. The American Statistician, 57(1):67–??, February 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://oberon.ingentaselect.com/cgi-bin/linker?
ini=asa&reqidx=/cw/asa/00031305/v57n1/s14/p67.

Hu:2007:MAV

[Hu07] Xiaomi Hu. Multivariate analysis without “vec” and “⊗”. The American Statistician, 61(1):64–66, February 2007. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hu:2008:TCC

[Hu08] Xiaomi Hu. A three-condition characterization of the Moore– Penrose generalized inverse. The American Statistician, 62(3): 216–218, August 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hunter:2005:TCS

[Hun05] David R. Hunter. Teaching computing in statistical theory courses. The American Statistician, 59(4):327–333, November 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hand:2003:LVG

[HV03] David J. Hand and Veronica Vinciotti. Local versus global models for classification problems: Fitting models where it matters. The American Statistician, 57(2):124–??, May 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v57n2/s11/p124.

Hamedani:2009:LE

[HV09] G. G. Hamedani and H. W. Volkmer. Letter to the Editor. The American Statistician, 63(3):295, August 2009. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hwang:2009:SPB

[Hwa09] Leng-Cheng Hwang. A simple proof of the binomial theorem using differential calculus. The American Statistician, 63(1):43– 44, February 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hyden:2001:RBT

[Hyd01] Paul Hyden. Reviews of books and teaching materials: Statistical Quality Assurance Methods for Engineers. The American Statistician, 55(2):167, May 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2686006.

Hyung:2001:RBT

[Hyu01] Namwon Hyung. Reviews of books and teaching materials: Elements of Forecasting, 2nd Edition, by Francis X. Diebold. The American Statistician, 55(4):371–372, November 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685705.

Inoue:2005:RBB

[IBP05] Lurdes Y. T. Inoue, Donald A. Berry, and Giovanni Parmigiani. Relationship between Bayesian and frequentist sample size determination. The American Statistician, 59(1):79–87, February 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Irony:2000:AOS

[IdBPT00] Telba Z. Irony, Carlos A. de B. Pereira, and Ram C. Tiwari. Analysis of opinion swing: Comparison of two correlated proportions. The American Statistician, 54(1):57–62, February 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685613.

Iglewicz:2007:EFP

[Igl07] Boris Iglewicz. Einstein’s first published paper. The American Statistician, 61(4):339–342, November 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Iman:2006:SAF

[IJW06] Ronald L. Iman, Mark E. Johnson, and Charles C. Watson, Jr. Statistical aspects of forecasting and planning for hurricanes. The American Statistician, 60(2):105–121, May 2006. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Iversen:2001:TCD

[IM01] Philip Iversen and Mervyn Marasinghe. Teacher’s corner: Dynamic graphical tools for teaching experimental design and analysis concepts. The American Statistician, 55(4):345–351, November 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685699.

Irizarry:2001:LRM

[Iri01] Rafael A. Irizarry. Local regression with meaningful parameters. The American Statistician, 55(1):72–79, February 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.amstat.org/publications/tas/
Irizarry.htm;   http://www.jstor.org/stable/2685532.

Jackson:2008:DUS

[Jac08] Christopher H. Jackson. Displaying uncertainty with shading. The American Statistician, 62(4):340–347, November 2008. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Jaditz:2000:DPI

[Jad00] Ted Jaditz. Are the digits of π an independent and identically distributed sequence? The American Statistician, 54(1):12–16, February 2000. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685604.

Jammalamadaka:2000:BRB

[Jam00] Sreenivasa Rao Jammalamadaka. Book review: Statistical Distributions in Engineering. The American Statistician, 54(3): 226, August 2000. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685598.

Jammalamadaka:2001:BRBa

[Jam01a] Sreenivasa Rao Jammalamadaka. Book review: [untitled]. The American Statistician, 55(4):372, November 2001. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685706.

Jammalamadaka:2001:BRBb

[Jam01b] Sreenivasa Rao Jammalamadaka. Book review: [untitled]. The American Statistician, 55(4):372, November 2001. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685708.

Jammalamadaka:2001:RBTb

[Jam01c] Sreenivasa Rao Jammalamadaka. Reviews of books and teaching materials: Statistics in Psychology: An Historical Perspective, Second Edition. The American Statistician, 55(4):372–??, 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Jammalamadaka:2001:RBTa

[Jam01d] Sreenivasa Rao Jammalamadaka. Reviews of books and teaching materials: The Nothing That Is: a Natural History of Zero. The American Statistician, 55(4):372, 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Jammalamadaka:2002:BRB

[Jam02] Sreenivasa Rao Jammalamadaka. Book review: Applied Nonparametric Statistical Methods, 3rd Edition. The American Statistician, 56(3):251–??, August 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://oberon.

ingentaselect.com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/
00031305/v56n3/s20/p251.

Jammalamadaka:2003:BRB

[Jam03] Sreenivasa Rao Jammalamadaka. Book review: Introduction to Linear Regression Analysis (3rd ed.). The American Statistician, 57(1):67–??, February 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://oberon.

ingentaselect.com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/
00031305/v57n1/s13/p67.

Janky:2000:TCS

[Jan00]  Denis G. Janky.  Teacher’s corner: Sometimes pooling

for analysis of variance hypothesis tests: a review and study of a split-plot model. The American Statistician, 54 (4):269–279, November 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.amstat.org/publications/tas/Agresti.htm;   http:/
/www.amstat.org/publications/tas/Janky.htm;   http://
www.jstor.org/stable/2685778.

Jeske:2004:TAM

[JB04] Daniel R. Jeske and Todd Blessinger. Tunable approximations for the mean and variance of the maximum of heterogeneous geometrically distributed random variables. The American Statistician, 58(4):322–327, November 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/27643591.

Jersky:2003:BRB

[Jer03] Brian Jersky. Book review: Hyperstat (2nd ed.). The American Statistician, 57(4):316–??, November 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v57n4/s17/p316.

Jones:2004:HMG

[JH04] M. C. Jones and Martin L. Hazelton. Hazelton, M. L. (2003), “A Graphical Tool for Assessing Normality,” The American Statistician,“ 57, 285–288: Comment by Jones and Reply. The American Statistician, 58(2):176–177, May 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/27643543.

Johnstone:2009:CWS

[Joh09] Iain Johnstone. Comment: What is statistics? The American Statistician, 63(2):111, May 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Jones:2002:KTP

[Jon02] M. C. Jones. On Khintchine’s theorem and its place in random variate generation. The American Statistician, 56(4):304– ??, November 2002. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://oberon.ingentaselect.
com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/00031305/v56n4/
s7/p304.

Jeske:2001:MLE

[JP01] Daniel R. Jeske and Hoang Pham. On the maximum likelihood estimates for the Goel–Okumoto software reliability model. The American Statistician, 55(3):219–222, August 2001. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685804.

Jeske:2003:REI

[JS03] Daniel R. Jeske and Ashwin Sampath. A real example that illustrates interesting properties of bootstrap bias correction. The American Statistician, 57(1):62–??, February 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v57n1/s10/p62.

Joffe:2004:MSC

[JTFK04] Marshall M. Joffe, Thomas R. Ten Have, Harold I. Feldman, and Stephen E. Kimmel. Model selection, confounder control, and marginal structural models: Review and new applications. The American Statistician, 58(4):272–279, November 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/27643582.

Johnson:2004:FRD

[JW04] H. Dean Johnson and Dennis A. Warner. Factors relating to the degree to which statistical consulting clients deem their consulting experiences to be a success. The American Statistician, 58(4): 280–289, November 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/27643583.

Kagan:2001:ALC

[Kag01] Abram Kagan. Another look at the Cramer–Rao inequality. The American Statistician, 55(3):211–212, August 2001. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685802.

Kahn:2006:DMC

[Kah06] Michael Kahn. Data matters: Conceptual statistics for a random world. The American Statistician, 60(3):292–293, August 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kang:2000:BRB

[Kan00] Heejoon Kang. Book review: [untitled]. The American Statistician, 54(1):84, February 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685623.

Kuindersma:2007:TBM

[KB07] Scott R. Kuindersma and Brian S. Blais. Teaching Bayesian model comparison with the three-sided coin. The American Statistician, 61(3):239–244, August 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Koehler:2009:AMC

[KBH09] Elizabeth Koehler, Elizabeth Brown, and Sebastien J.-P. A. Haneuse. On the assessment of Monte Carlo error in simulationbased statistical analyses. The American Statistician, 63(2):155– 162, May 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Khuri:2002:EFN

[KC02] André Khuri and George Casella. The existence of the first negative moment revisited. The American Statistician, 56(1):44– 47, February 2002. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
3087326.

Kuiper:2009:GLI

[KC09] Shonda Kuiper and Linda Collins. Guided labs that introduce statistical techniques used in research from multiple disciplines. The American Statistician, 63(4):343–347, November 2009. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kelly:2004:RSP

[Kel04] Patrick J. Kelly. A review of software packages for analyzing correlated survival data. The American Statistician, 58(4):337– 342, November 2004. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
27643594.

Kern:2007:IRA

[Ker07] John C. Kern II. Introduction to regression analysis. The American Statistician, 61(1):101, February 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kesling:2000:MBP

[Kes00] George D. Kesling. The mighty bonus point. The American Statistician, 54(3):215–216, August 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:/

/www.amstat.org/publications/tas/Kesling.htm;   http:
//www.amstat.org/publications/tas/Long.htm;   http://
www.jstor.org/stable/2685593.

Kish:2004:RPR

[KH04] Leslie Kish and Irene Hess. A “replacement” procedure for reducing the bias of nonresponse. The American Statistician, 58(4): 295–297, November 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/27643586.

Khatry:2004:SCB

[Kha04] Deepak B. Khatry. Statistical contributions in biotechnology and nonclinical drug-discovery research. The American Statistician, 58(3):199, August 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/27643555.

Khan:2008:SRL

[Kha08] M. G. M. Khan. Statistics for real-life sample surveys: Nonsimple-random samples and weighted data. The American Statistician, 62(4):361, November 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kiessler:2006:PSE

[Kie06] Peter C. Kiessler. Probability and statistics for engineers and scientists. The American Statistician, 60(4):343, November 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kirmani:2000:BRB

[Kir00] Syed N. U. A. Kirmani. Book review: Engineering Statistics and Applied Statistics and Probability for Engineers (2nd ed.). The American Statistician, 54(3):226, August 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685597.

Karr:2006:FEU

[KKO^{+}06] A. F. Karr, C. N. Kohnen, A. Oganian, J. P. Reiter, and A. P. Sanil. A framework for evaluating the utility of data altered to protect confidentiality. The American Statistician, 60(3):224–232, August 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Koenker:2009:MMW

[Koe09] Roger Koenker. The median is the message: Wilson and Hilfertys experiments on the Law of Errors. The American Statistician, 63 (1):20–25, February 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Koschat:2005:CST

[Kos05] Martin A. Koschat. A case for simple tables. The American Statistician, 59(1):31–40, February 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Krause:2007:MSA

[KP07] Andreas Krause and José Pinheiro. Modeling and simulation to adjust p values in presence of a regression to the mean effect. The American Statistician, 61(4):302–307, November 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

King:2002:PIM

[KR02] Elizabeth N. King and Thomas P. Ryan. A preliminary investigation of maximum likelihood logistic regression versus exact logistic regression. The American Statistician, 56(3):163–??, August 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://oberon.ingentaselect.com/
cgi-bin/linker?ini=asa&reqidx=/cw/asa/00031305/v56n3/
s1/p163.

Kreft:2001:RBT

[Kre01] Ita G. G. Kreft. Reviews of books and teaching materials: Statistical Concepts: Education and Behavioral Sciences (2nd ed.). The American Statistician, 55(3):259, August 2001. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685815.

Kagan:2005:SP

[KS05] Abram Kagan and Lawrence A. Shepp. A sufficiency paradox. The American Statistician, 59(1):54–56, February 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kotz:2008:VPT

[KS08] Samuel Kotz and Edith Seier. Visualizing peak and tails to introduce kurtosis. The American Statistician, 62(4):348–354, November 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Krishnamoorthy:2002:HTA

[KT02] K. Krishnamoorthy and Jessica Thomson. Hypothesis testing about proportions in two finite populations. The American Statistician, 56(3):215–??, August 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://oberon.

ingentaselect.com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/
00031305/v56n3/s9/p215.

Kuk:2009:NWD

[Kuk09] Anthony Y. C. Kuk. A new way to derive locally most powerful rank tests. The American Statistician, 63(3):278–280, August 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kuonen:2000:SAC

[Kuo00] Diego Kuonen. A saddlepoint approximation for the Collector’s Problem. The American Statistician, 54(3):165–169, August 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.amstat.org/publications/
tas/Kuonen.htm;   http://www.jstor.org/stable/2685585.

Kvam:2000:EAL

[Kva00] Paul H. Kvam. The effect of active learning methods on student retention in engineering statistics. The American Statistician, 54(2):136–140, May 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.amstat.org/publications/tas/Kvam.htm;   http://
www.jstor.org/stable/2686032.

Kennedy:2007:OBS

[KYA^{+}07] Richard E. Kennedy, Sharon D. Yeatts, Kellie J. Archer, Chris Gennings, and Viswanathan Ramakrishnan. Opportunities for biostatistics students. The American Statistician, 61(2):120–126, May 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lachenbruch:2000:BRB

[Lac00] Peter A. Lachenbruch. Book review: A Handbook of Statistical Analyses Using Stata. The American Statistician, 54(2):153–154, May 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2686039.

LaMotte:2008:SFP

[LaM08] Lynn Roy LaMotte. Sufficiency in finite parameter and sample spaces. The American Statistician, 62(3):211–215, August 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lange:2004:CSO

[Lan04] Kenneth Lange. Computational statistics and optimization theory at UCLA. The American Statistician, 58(1):9–??, February 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://oberon.ingentaselect.com/
cgi-bin/linker?ini=asa&reqidx=/cw/asa/00031305/v58n1/
s4/p9.

Lane:2006:SM

[Lan06a] David Lane. Statistics without math. The American Statistician, 60(2):205–206, May 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Langsrud:2006:ECP

[Lan06b] Øyvind Langsrud. Explaining correlations by plotting orthogonal contrasts. The American Statistician, 60(4):335–339, November 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Landes:2009:PPP

[Lan09] Reid D. Landes. Passing on the passion for the profession. The American Statistician, 63(2):163–172, May 2009. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lawson:2003:OSS

[Law03] John Lawson. One-step screening and process optimization experiments. The American Statistician, 57(1):15–??, February 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v57n1/s2/p15.

Lindsay:2000:MDT

[LB00] Bruce G. Lindsay and Prasanta Basak. Moments determine the tail of a distribution (but not much else). The American Statistician, 54(4):248–251, November 2000. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.amstat.org/publications/tas/
Lindsay.htm;   http://www.jstor.org/stable/2685775.

Langbehn:2000:LEa

[LBH^{+}00] Douglas R. Langbehn, Vance Berger, James J. Higgins, R. Clifford Blair, Colin L. Mallows, John Ludbrook, Hugh A. F. Dudley, Carles M. Cuadras, David Rindskopf, Alan Agresti, and Brent Coull. Letters to the Editor. The American Statistician, 54 (1):85–88, February 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685624.

Li:2003:DSE

[LBTM03] Xiuhong Li, Jennifer M. Buechner, Patrick M. Tarwater, and Alvaro Muñoz. A diamond-shaped equiponderant graphical display of the effects of two categorical predictors on continuous outcomes. The American Statistician, 57(3):193–??, August 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v57n3/s12/p193.

Lauzon:2009:EMC

[LC09] Carolyn Lauzon and Brian Caffo. Easy multiplicity control in equivalence testing using two one-sided tests. The American Statistician, 63(2):147–154, May 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See correction [CLR13].

Langbehn:2000:LEb

[LDH^{+}00] Douglas R. Langbehn, Rafe M. J. Donahue, Klaus Hinkelmann, Russell D. Wolfinger, Walter W. Stroup, Dan Voss, and Antonio Martin Andres. Letters to the Editor. The American Statistician, 54(3):227–230, August 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685599.

Long:2000:UHC

[LE00] J. Scott Long and Laurie H. Ervin. Using heteroscedasticity consistent standard errors in the linear regression model. The American Statistician, 54(3):217–224, August 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2685594.

Leary:2004:TSU

[Lea04] Richard J. C. Leary. Teaching statistics using baseball. Jim Albert. The American Statistician, 58(4):358, November 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Ledolter:2003:BRB

[Led03] Johannes Ledolter. Book review: The Statistical Sleuth (2nd ed.). The American Statistician, 57(2):145–??, May 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v57n2/s18/p145.

Lee:2007:CCC

[Lee07] Herbert K. H. Lee. Chocolate chip cookies as a teaching aid. The American Statistician, 61(4):351–355, November 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lee:2008:BMS

[Lee08] Herbert K. H. Lee. Bayesian methods: a social and behavioral sciences approach. The American Statistician, 62(4):356, November 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lee:2009:RWL

[Lee09] Jaechoul Lee. A reformulation of weighted least squares estimators. The American Statistician, 63(1):49–55, February 2009. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lenth:2000:BRB

[Len00] Russell V. Lenth. Book review: [untitled]. The American Statistician, 54(1):83–84, February 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685622.

Lenth:2001:BRBa

[Len01a] Russell V. Lenth. Book review: [Introduction to Design and Analysis of Experiments]. The American Statistician, 55(1):80– 81, February 2001. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685533.

Lenth:2001:BRBb

[Len01b] Russell V. Lenth. Book review: [untitled]. The American Statistician, 55(3):256–257, August 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685812.

Lenth:2001:BRBc

[Len01c] Russell V. Lenth. Book review: [untitled]. The American Statistician, 55(4):370, November 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685703.

Lenth:2001:RBTa

[Len01d] Russell V. Lenth. Reviews of books and teaching materials: A First Course in the Design of Experiments: a Linear Models Approach. The American Statistician, 55(4):370–??, 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lenth:2001:RBTb

[Len01e] Russell V. Lenth. Reviews of books and teaching materials: So Much Data So Little Math: How to Predict Data Trends, 5 Easy Profitable Methods. The American Statistician, 55(3):256, 2001.

CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lenth:2001:SPS

[Len01f] Russell V. Lenth. Statistical practice: Some practical guidelines for effective sample size determination. The American Statistician, 55(3):187–193, August 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685797.

Lenth:2002:RBT

[Len02] Russell V. Lenth. Reviews of books and teaching materials: Statistics on the Table: The History of Statistical Concepts and Methods. The American Statistician, 56(1):76, February 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/3087332.

Lenth:2004:BR

[Len04a] Russell V. Lenth. Book reviews. The American Statistician, 58 (2):92, May 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lenth:2004:CCC

[Len04b] Russell V. Lenth. Corrections: CAST (computer-assisted statistics teaching). The American Statistician, 58(2):178, May 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/27643546.

Lesser:2004:WHB

[Les04] Lawrence M. Lesser. Wainer, H., and Brown, L. M. (2004), “Two Statistical Paradoxes in the Interpretation of Group Differences: Illustrated with Medical School Admission and Licensing Data,” The American Statistician, 58, 117–123: Comment by Lesser. The American Statistician, 58(4):362, November 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/27643601.

Levine:2002:HJJ

[Lev02] Douglas W. Levine. Hanley, J. A., Joseph, L., Platt, R. W., Chung, M. K., and Bélisle, P. (2001), “Visualizing the Median as the Minimum-Deviation Location,” The American Statistician, 55, 150-152: Comment by Levine. The American Statistician, 56 (1):81–82, February 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/3087340.

Levine:2000:BRB

[LF00] Richard A. Levine and Juanjuan Fan. Book review: Practical Statistics by Example Using Microsoft Excel. The American Statistician, 54(2):151–153, May 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2686037.

Lann:2006:TMM

[LF06] Avital Lann and Ruma Falk. Tell me the method, I’ll give you the mean. The American Statistician, 60(4):322–327, November 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lovett:2000:ACT

[LG00] Marsha C. Lovett and Joel B. Greenhouse. Applying cognitive theory to statistics instruction. The American Statistician, 54(3):196–206, August 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

amstat.org/publications/tas/Lovett.htm;   http://www.
jstor.org/stable/2685590.

Langohr:2005:LMU

[LG05] Klaus Langohr and Guadalupe Gómez. Likelihood maximization using Web-based optimization tools. The American Statistician, 59(2):192–202, May 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lyles:2009:FLD

[LGH09] Robert H. Lyles, Ying Guo, and Andrew N. Hill. A fresh look at the discriminant function approach for estimating crude or adjusted odds ratios. The American Statistician, 63(4):320–327, November 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Love:2002:SEM

[LH02] Thomas E. Love and David K. Hildebrand. Statistics education and the making statistics more effective in schools of business conferences. The American Statistician, 56(2):107–??, May 2002.

CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v56n2/s6/p107.

Lee:2000:EAE

[LHD00] J. Jack Lee, Kenneth R. Hess, and Joel A. Dubin. Extensions and applications of event charts. The American Statistician, 54 (1):63–70, February 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685614.

Lehr:2000:LE

[LIdBP^{+}00] Robert G. Lehr, Telba Z. Irony, Carlos A. de B. Pereira, Ram C. Tiwari, Michael P. Cohen, Carles M. Cuadras, Lloyd S. Nelson, M. Alvo, P. Cabilio, Simo Puntanen, George P. H. Styan, and Hans Joachim Werner. Letters to the Editor. The American Statistician, 54(4):325–328, November 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2685791.

Liss:2008:MAP

[Lis08] Charles L. Liss. Multivariable analysis: a practical guide for clinicians and study design and statistical analysis: a practical guide for clinicians. The American Statistician, 62(1):92, February 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Little:2006:CB

[Lit06] Roderick J. Little. Calibrated Bayes. The American Statistician, 60(3):213–223, August 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lesch:2009:SST

[LJ09] Scott M. Lesch and Daniel R. Jeske. Some suggestions for teaching about normal approximations to Poisson and binomial distribution functions. The American Statistician, 63(3):274–277, August 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Larreamendy-Joerns:2005:SOS

[LJLC05] Jorge Larreamendy-Joerns, Gaea Leinhardt, and Javier Corredor. Six online statistics courses: Examination and review. The American Statistician, 59(3):240–251, August 2005. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Liao:2003:ACD

[LM03] J. G. Liao and Dan McGee. Adjusted coefficients of determination for logistic regression. The American Statistician, 57(3):161–??, August 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://oberon.ingentaselect.com/
cgi-bin/linker?ini=asa&reqidx=/cw/asa/00031305/v57n3/
s3/p161.

Leemis:2008:UDR

[LM08] Lawrence M. Leemis and Jacquelyn T. McQueston. Univariate distribution relationships. The American Statistician, 62(1):45– 53, February 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lohr:2004:SFS

[Loh04] Sharon L. Lohr. Statistical frontiers in survey sampling. The American Statistician, 58(2):145–149, May 2004. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/27643524.

Losinger:2004:HDS

[Los04a] Willard C. Losinger. How to do statistical research. The American Statistician, 58(3):200, August 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/27643556.

Losinger:2004:RGW

[Los04b] Willard C. Losinger. A review of the GUM workbench. The American Statistician, 58(2):165–167, May 2004. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/27643529.

Leon:2000:UCH

[LP00] Ramón V. León and William C. Parr. Use of course home pages in teaching statistics. The American Statistician, 54(1):44–48, February 2000. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685610.

Liao:2001:FSA

[LR01] J. G. Liao and Ori Rosen. Fast and stable algorithms for computing and sampling from the noncentral hypergeometric distribution. The American Statistician, 55(4):366–369, November 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685702.

Loken:2006:PCM

[LR06] Eric Loken and Michael J. Rovine. Peirce’s 19th Century mixture model approach to rater agreement. The American Statistician, 60(2):158–161, May 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Larocque:2008:CID

[LR08] Denis Larocque and Ronald H. Randles. Confidence intervals for a discrete population median. The American Statistician, 62 (1):32–39, February 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lad:2001:DAE

[LS01] Frank Lad and Romano Scozzafava. Distributions agreeing with exchangeable sequential forecasting. The American Statistician, 55(2):131–139, May 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2686000.

Landon:2008:CCP

[LS08] Joshua Landon and Nozer D. Singpurwalla. Choosing a coverage probability for prediction intervals. The American Statistician, 62(2):120–124, May 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Leemis:2000:SDS

[LSE00] Lawrence M. Leemis, Bruce W. Schmeiser, and Diane L. Evans. Survival distributions satisfying Benford’s Law. The American Statistician, 54(4):236–241, November 2000. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.amstat.org/publications/tas/
Leemis.htm;   http://www.jstor.org/stable/2685773.

Langford:2001:PBP

[LSO01] Eric Langford, Neil Schwertman, and Margaret Owens. Is the property of being positively correlated transitive? The American Statistician, 55(4):322–325, November 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2685695. See comment [Fri15].

Ludbrook:2005:RFL

[Lud05] John Ludbrook. R. A. Fisher’s life and death in Australia, 1959– 1962. The American Statistician, 59(2):164–165, May 2005. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lund:2002:BRBb

[Lun02a] Robert Lund. Book review: Stochastic Processes. An Introduction. The American Statistician, 56(4):332–??, November 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v56n4/s19/p332.

Lund:2002:BRBa

[Lun02b] Ulric J. Lund. Book review: Understanding Statistical Concepts Using S-Plus. The American Statistician, 56(3):249–??, August 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://oberon.ingentaselect.com/cgi-
bin/linker?ini=asa&reqidx=/cw/asa/00031305/v56n3/s18/
p249.

Lund:2004:BRBb

[Lun04a] Robert Lund. Book review: Stochastic Processes and Their Applications, by Frank E. Beichelt; L. Paul Fatti. The American Statistician, 58(3):261, August 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/27643571.

Lund:2004:SPT

[Lun04b] Robert Lund. Stochastic processes and their applications. The American Statistician, 58(3):261, August 2004. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lund:2004:BRBa

[Lun04c] Robert B. Lund. Book review: Elementary Probability Theory with Stochastic Processes and an Introduction to Mathematical Finance, by Kai Lai Chung; Farid Aitsahlia. The American Statistician, 58(2):173–174, May 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/27643534.

Lund:2007:RWS

[Lun07a] Robert Lund. Revenge of the white swan. The American Statistician, 61(3):189–192, August 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lund:2007:BRB

[Lun07b] Ulric J. Lund. Book review: A Handbook of Statistical Analyses Using R. The American Statistician, 61(3):277–278, August 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lynn:2003:SCA

[Lyn03] Henry S. Lynn. Suppression and confounding in action. The American Statistician, 57(1):58–??, February 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v57n1/s9/p58.

Mackisack:2001:RBT

[Mac01] Margaret Mackisack. Reviews of books and teaching materials: Chance Encounters: a First Course in Data Analysis and Inference, by Christopher J. Wild; George A. F. Seber. The American Statistician, 55(3):255–256, August 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2685811.

Machler:2007:SIU

[Mäc07] Martin Mächler. Statistics: An introduction using R. The American Statistician, 61(1):100–101, February 2007. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Mammen:2006:SM

[Mam06] Enno Mammen. Statistical models. The American Statistician, 60(2):204–205, May 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Manola:2003:BRB

[Man03] Judith Manola. Book review: A Handbook of Statistical Analyses using S-Plus (2nd ed.). The American Statistician, 57(2):146–??, May 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://oberon.ingentaselect.com/

cgi-bin/linker?ini=asa&reqidx=/cw/asa/00031305/v57n2/
s19/p146.

Mandel:2007:CTH

[Man07] Micha Mandel. Censoring and truncation — highlighting the differences. The American Statistician, 61(4):321–324, November 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Marasinghe:2007:FCS

[Mar07] Mervyn Marasinghe. A first course in statistical methods. The American Statistician, 61(1):95–97, February 2007. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Massey:2004:DSC

[Mas04] Kenneth Massey. Discussion: Statistics and the college football championship. The American Statistician, 58(3):185–187, August 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Morrell:2000:LRE

[MB00] Christopher H. Morrell and Larry J. Brant. Lines in random effects plots from the linear mixed-effects model. The American Statistician, 54(1):1–4, February 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2685602.

McGee:2006:APS

[MB06] Monnie McGee and Nora V. Bergasa. Analysis of a pilot study for amelioration of itching in liver disease. The American Statistician, 60(4):303–308, November 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

McCue:2001:SFM

[McC01] Kenneth F. McCue. The statistical foundations of the EI method. The American Statistician, 55(2):106–110, May 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685996.

McDonald:2004:IS

[McD04] Elaine T. McDonald. Intro stats. The American Statistician, 58(1):86–??, February 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://oberon.

ingentaselect.com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/
00031305/v58n1/s18/p86.

McGee:2005:ISA

[McG05] Daniel L. McGee. An introduction to survival analysis using Stata (rev.). The American Statistician, 59(1):113, February 2005. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

McIntyre:2005:MUS

[McI05] G. A. McIntyre. A method for unbiased selective sampling, using ranked sets. The American Statistician, 59(3):230–232, August 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Mease:2003:PML

[Mea03] David Mease. A penalized maximum likelihood approach for the ranking of college football teams independent of victory margins. The American Statistician, 57(4):241–??, November 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v57n4/s4/p241.

Mease:2004:DSC

[Mea04] David Mease. Discussion: Statistics and the college football championship. The American Statistician, 58(3):192–194, August 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Mehlum:2009:IPR

[Meh09] Halvor Mehlum. The island problem revisited. The American Statistician, 63(3):269–273, August 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Menard:2000:CDM

[Men00] Scott Menard. Coefficients of determination for multiple logistic regression analysis. The American Statistician, 54(1):17–24, February 2000. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685605.

Menard:2004:CSA

[Men04a] Scott Menard. Correction: Six approaches to calculating standardized logistic regression coefficients. The American Statistician, 58(4):364, November 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/27643603. See [Men04b].

Menard:2004:SAC

[Men04b] Scott Menard. Six approaches to calculating standardized logistic regression coefficients. The American Statistician, 58(3):218– 223, August 2004. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/

27643560. See correction [Men04a].

Meng:2009:DFW

[Men09] Xiao-Li Meng. Desired and feared — what do we do now and over the next 50 years? The American Statistician, 63(3):202– 210, August 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Miller:2007:FDD

[MF07] Gregory K. Miller and Stephanie L. Fridell. A forgotten discrete distribution? reviving the negative hypergeometric model. The American Statistician, 61(4):347–350, November 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Murad:2003:SSO

[MFS^{+}03] Havi Murad, Anat Fleischman, Siegal Sadetzki, Orna Geyer, and Laurence S. Freedman. Small samples and ordered logistic regression: Does it help to collapse categories of outcome? The American Statistician, 57(3):155–??, August 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v57n3/s2/p155.

Miao:2004:EDC

[MG04] Weiwen Miao and Joseph L. Gastwirth. The effect of dependence on confidence intervals for a population proportion. The American Statistician, 58(2):124–130, May 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/27643520.

Madigan:2009:CWS

[MG09] David Madigan and Andrew Gelman. Comment: What is statistics? The American Statistician, 63(2):114, May 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Miller:2001:RBT

[Mil01] Jacqueline B. Miller. Reviews of books and teaching materials: An Electronic Companion to Statistics, by G. W. Cobb; J. A. Witmer; J. D. Cryer; An Electronic Companion to Biostatistics, by A. A. Afifi; M. Lee; G. W. Cobb; An Electronic Companion to Business Statistics by J. D. Cryer; G. W. Cobb; J. A. Witmer. The American Statistician, 55(2):168–169, May 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2686008.

Miller:2003:BRB

[Mil03] Stephen J. Miller. Book reviews: Basic Statistics and Data Analysis (with InfoTrac and CD-ROM) (1st ed.), Introduction to Probability and Statistics (11th ed.) and Just the Essentials of Elementary Statistics (3rd ed.). The American Statistician, 57(3):219–??, August 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://oberon.ingentaselect.com/
cgi-bin/linker?ini=asa&reqidx=/cw/asa/00031305/v57n3/
s19/p219.

Miller:2004:BRBb

[Mil04a] Stephen J. Miller. Book review: CAST (Computer-Assisted Statistics Teaching). The American Statistician, 58(1):87–??, February 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://oberon.ingentaselect.com/
cgi-bin/linker?ini=asa&reqidx=/cw/asa/00031305/v58n1/
s19/p87.

Miller:2004:BRBa

[Mil04b] Stephen J. Miller. Book review: Fifty Fathoms: Statistics Demonstrations for Deeper Understanding. The American Statistician, 58(1):86–??, February 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://oberon.

ingentaselect.com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/
00031305/v58n1/s17/p86.

Miller:2004:BRBc

[Mil04c] Stephen J. Miller. Book review: Statistics in Action: Understanding a World of Data, by Anne E. Watkins; Richard L. Schaeffer; George W. Cobb. The American Statistician, 58(3):259–260, August 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/27643569.

Minnotte:2007:BRB

[Min07] Michael Minnotte. Book review: Introduction to Modern Nonparametric Statistics, by James J. Higgins. The American Statistician, 61(2):184, May 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Mirkin:2001:EWL

[Mir01] Boris Mirkin. Eleven ways to look at the chi-squared coefficient for contingency tables. The American Statistician, 55(2):111–120, May 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685997.

Missiakoulis:2006:PLA

[Mis06] Spyros Missiakoulis. Plutarch and Lucullus’ applied sampling. The American Statistician, 60(3):269–270, August 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Mills:2004:EAS

[MJ04] Jamie D. Mills and Elisa L. Johnson. An evaluation of activstats for SPSS for teaching and learning. The American Statistician, 58(3):254–258, August 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/27643567.

Morris:2009:UNN

[ML09] Carl N. Morris and Kari F. Lock. Unifying the named natural exponential families and their relatives. The American Statistician, 63(3):247–253, August 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Moerbeek:2005:RCVa

[Moe05a] Mirjam Moerbeek. Randomization of clusters versus randomization of persons within clusters. The American Statistician, 59 (1):72–78, February 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Moerbeek:2005:RCVb

[Moe05b] Mirjam Moerbeek. Randomization of clusters versus randomization of persons within clusters. The American Statistician, 59(2): 173–179, May 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Monahan:2004:TSC

[Mon04] John Monahan. Teaching statistical computing at North Carolina State University. The American Statistician, 58(1):6–??, February 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://oberon.ingentaselect.com/
cgi-bin/linker?ini=asa&reqidx=/cw/asa/00031305/v58n1/
s3/p6.

Monahan:2006:PCR

[Mon06] John F. Monahan. Professor C. R. Mudgeon and the “order,” or writing regression questions so that students do not need a calculator. The American Statistician, 60(1):50–52, February 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Moore:2001:UPF

[Moo01] David S. Moore. Undergraduate programs and the future of academic statistics. The American Statistician, 55(1):1–6, February 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685522.

Moore:2005:PGS

[Moo05] David S. Moore. Preparing graduate students to teach statistics. The American Statistician, 59(1):1–3, February 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Moore:2007:SEI

[Moo07] Page C. Moore. Statistics explained: An introductory guide for life scientists. The American Statistician, 61(3):274, August 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Moreno:2007:ISC

[Mor07] Jerry L. Moreno. Investigating statistical concepts, applications, and methods. The American Statistician, 61(1):98–99, February 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Morgan:2007:NST

[MPR07] B. J. T. Morgan, K. J. Palmer, and M. S. Ridout. Negative score test statistic. The American Statistician, 61(4):285–288, November 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Mandel:2009:SBC

[MR09] Micha Mandel and Yosef Rinott. A selection bias conflict and frequentist versus Bayesian viewpoints. The American Statistician, 63(3):211–217, August 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Mrug:2007:EDS

[Mru07] Sylvie Mrug. Experimental design and statistics for psychology: a first course. The American Statistician, 61(3):275–276, August 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

McCready:2001:SPS

[MS01] Thomas A. McCready and Neil C. Schwertman. Statistical practice: The statistical paleontology of Charles Lyell and the Coupon Problem. The American Statistician, 55(4):272–278, November 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685687.

Michael:2002:MAS

[MS02] John R. Michael and William R. Schucany. The mixture approach for simulating bivariate distributions with specified correlations. The American Statistician, 56(1):48–54, February 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/3087327.

Maity:2006:TST

[MS06] Arnab Maity and Michael Sherman. The two-sample T test with one variance unknown. The American Statistician, 60(2):163–166, May 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

McCarthy:2008:PBV

[MSL^{+}08] John McCarthy, Howard Stanislevic, Mark Lindeman, Arlene S. Ash, Vittorio Addona, and Mary Batcher. Percentage-based versus statistical-power-based vote tabulation audits. The American Statistician, 62(1):11–16, February 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Murdoch:2008:VRV

[MTA08] Duncan J. Murdoch, Yu-Ling Tsai, and James Adcock. P -values are random variables. The American Statistician, 62(3):242–245, August 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Mukhopadhyay:2006:MMO

[Muk06] Nitis Mukhopadhyay. MVUE for the mean with one observation. The American Statistician, 60(1):71–74, February 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Molenberghs:2007:LRS

[MV07] Geert Molenberghs and Geert Verbeke. Likelihood ratio, score, and Wald tests in a constrained parameter space. The American Statistician, 61(1):22–27, February 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

McGrath:2005:QCT

[MY05] Richard N. McGrath and Arthur B. Yeh. A quick, compact, twosample dispersion test. The American Statistician, 59(1):47–53, February 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Nadarajah:2009:PDP

[Nad09] Saralees Nadarajah. PDFs and dual PDFs. The American Statistician, 63(1):45–48, February 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Nash:2006:SSP

[Nas06] J. C. Nash. Spreadsheets in statistical practice: Another look. The American Statistician, 60(3):287–289, August 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Nayak:2002:RCT

[Nay02] Tapan K. Nayak. Rao–Cramer type inequalities for mean squared error of prediction. The American Statistician, 56(2):102–??, May 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://oberon.ingentaselect.com/
cgi-bin/linker?ini=asa&reqidx=/cw/asa/00031305/v56n2/
s5/p102.

Nelsen:2001:BRB

[Nel01] Roger Nelsen. Book review: Probability and Random Variables: a Beginner’s Guide. The American Statistician, 55(1):82–83, February 2001. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685536.

Nelson:2007:IMM

[Nel07] Kerrie Nelson. Introduction to mixed modelling: Beyond regression and analysis of variance. The American Statistician, 61(4): 368–369, November 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Newcombe:2001:LCI

[New01] Robert G. Newcombe. Logit confidence intervals and the inverse sinh transformation. The American Statistician, 55(3):200– 202, August 2001. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685799.

Niermann:2005:OOT

[Nie05] Stefan Niermann. Optimizing the ordering of tables with evolutionary computation. The American Statistician, 59(1):41–46, February 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Nolan:2009:CWS

[NL09] Deborah Nolan and Duncan Temple Lang. Comment: What is statistics? The American Statistician, 63(2):117–121, May 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Norwood:2007:DOS

[Nor07] Janet L. Norwood. Discussion: The official statistics Olympic challenge. The American Statistician, 61(1):13–15, February 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Nordman:2009:UHE

[NVB09] Daniel J. Nordman, Stephen B. Vardeman, and Melissa A. Bingham. Uniformly hyper-efficient Bayes inference in a class of nonregular problems. The American Statistician, 63(3):234–238, August 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

OBrien:2003:GIO

[OF03] Timothy E. O’Brien and Gerald M. Funk. A gentle introduction to optimal design for regression models. The American Statistician, 57(4):265–??, November 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v57n4/s9/p265.

Oster:2008:ESS

[OH08a] Robert A. Oster and Joseph M. Hilbe. An examination of statistical software packages for parametric and nonparametric data analyses using exact methods. The American Statistician, 62 (1):74–84, February 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See rejoinder [OH08b].

Oster:2008:RES

[OH08b] Robert A. Oster and Joseph M. Hilbe. Rejoinder to “An Examination of Statistical Software Packages for Parametric and Nonparametric Data Analyses Using Exact Methods”. The American Statistician, 62(2):173–176, May 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See [OH08a].

Ord:2007:ECD

[OI07] J. Keith Ord and Harvey J. Iglarsh. The estimation of conditional distributions from large databases. The American Statistician, 61 (4):308–314, November 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Owen:2000:PDA

[OSC00] W. J. Owen, D. Sinha, and M. H. Capozzoli. A paireddata analysis for a lifetime distribution. The American Statistician, 54(4):252–256, November 2000. CODEN ASTAAJ.

ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.amstat.org/publications/tas/Owen.htm;   http://
www.jstor.org/stable/2685776.

Oster:2002:ESS

[Ost02] Robert A. Oster. An examination of statistical software packages for categorical data analysis using exact methods. The American Statistician, 56(3):235–??, August 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v56n3/s13/p235.

Oster:2003:ESS

[Ost03] Robert A. Oster. An examination of statistical software packages for categorical data analysis using exact methods — Part II. The American Statistician, 57(3):201–??, August 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v57n3/s14/p201.

Oster:2009:SENa

[Ost09a] Robert A. Oster. Section Editor’s notes. The American Statistician, 63(2):187–188, May 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Oster:2009:SENb

[Ost09b] Robert A. Oster. Section Editor’s notes. The American Statistician, 63(4):400, November 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Owen:2001:BRB

[Owe01] William B. Owen. Book review: Introduction to Business Statistics: a Computer Integrated, Data Analysis Approach. The American Statistician, 55(1):83–84, February 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2685538.

Padgett:2008:VSP

[Pad08] Lakshmi Padgett. Variations on split plot and split block experiment designs. The American Statistician, 62(2):184, May 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Pagano:2000:BRB

[Pag00] Marcello Pagano. Book review: StatConcepts: a Visual Tour of Statistical Ideas. The American Statistician, 54(4):323–324, November 2000. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685789.

Palacios:2008:SSR

[Pal08] José Luis Palacios. On the simple symmetric random walk and its maximal function. The American Statistician, 62(2):138–140, May 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Pan:2002:NUM

[Pan02] Wei Pan. A note on the use of marginal likelihood and conditional likelihood in analyzing clustered data. The American Statistician, 56(3):171–??, August 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://oberon.ingentaselect.com/cgi-bin/linker?
ini=asa&reqidx=/cw/asa/00031305/v56n3/s2/p171.

Pardoe:2004:MSS

[Par04] Iain Pardoe. Multidimensional scaling for selecting small groups in college courses. The American Statistician, 58(4):317–321, November 2004. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
27643590.

Pawitan:2000:RFW

[Paw00] Yudi Pawitan. A reminder of the fallibility of the Wald statistic: Likelihood explanation. The American Statistician, 54(1):54– 56, February 2000. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685612.

Parker:2003:SSM

[PB03] Robert A. Parker and Nancy G. Berman. Sample size: More than calculations. The American Statistician, 57(3):166–??, August 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://oberon.ingentaselect.com/
cgi-bin/linker?ini=asa&reqidx=/cw/asa/00031305/v57n3/
s4/p166.

Pardoe:2002:GMA

[PC02] Iain Pardoe and R. Dennis Cook. A graphical method for assessing the fit of a logistic regression model. The American Statistician, 56(4):263–??, November 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v56n4/s2/p263.

Piepho:2005:SCI

[PE05] Hans-Peter Piepho and Katharina Emrich. Simultaneous confidence intervals for two estimable functions and their ratio under a linear model. The American Statistician, 59(4):292–300, November 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Peng:2002:QAL

[PH02] Roger D. Peng and Nicolas W. Hengartner. Quantitative analysis of literary styles. The American Statistician, 56(3):175–??, August 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://oberon.ingentaselect.com/
cgi-bin/linker?ini=asa&reqidx=/cw/asa/00031305/v56n3/
s3/p175.

Potthoff:2008:FFC

[PHSN08] Richard F. Potthoff, Susan Halabi, Joellen M. Schildkraut, and Beth Newman. Flexible frames and control sampling in case-control studies. The American Statistician, 62(4):307–313, November 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Pickle:2006:GDT

[Pic06] Linda Pickle. Graphic discovery: a trout in the milk and other visual adventures. The American Statistician, 60(2):203, May 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Piegorsch:2001:BRB

[Pie01] Walter W. Piegorsch. Book review: Statistics and Experimental Design for Toxicologists (3rd ed.). The American Statistician, 55(1):81, February 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685534.

Pan:2000:NML

[PLC00] Wei Pan, Thomas A. Louis, and John E. Connett. A note on marginal linear regression with correlated response data. The American Statistician, 54(3):191–195, August 2000. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.amstat.org/publications/tas/Pan.htm;
http://www.jstor.org/stable/2685589.

Peikes:2008:PSM

[PMO08] Deborah N. Peikes, Lorenzo Moreno, and Sean Michael Orzol. Propensity score matching. The American Statistician, 62(3): 222–231, August 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Piepho:2007:SSS

[PO07] Hans-Peter Piepho and Joseph Ochieng Ogutu. Simple statespace models in a mixed model framework. The American Statistician, 61(3):224–232, August 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Pourahmadi:2002:RBT

[Pou02] Mohsen Pourahmadi. Reviews of books and teaching materials: A Course in Time Series Analysis. The American Statistician, 56(1):77, February 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/3087334.

Pavlides:2009:HLS

[PP09] Marios G. Pavlides and Michael D. Perlman. How likely is Simpson’s Paradox? The American Statistician, 63(3):226–233, August 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Prado:2008:IDE

[Pra08a] Raquel Prado. Introduction to design of experiments. The American Statistician, 62(4):362, November 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Pratter:2008:SGJ

[Pra08b] Frederick M. Pratter. SAS graphics for Java: Examples using SAS AppDev Studio and the output delivery system. The American Statistician, 62(4):359, November 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Prins:2005:LMR

[Pri05a] Samantha C. Bates Prins. Linear models with R. The American Statistician, 59(3):276, August 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Prins:2005:SAD

[Pri05b] Samantha C. Bates Prins. Statistical analysis and data display: An intermediate course with examples in S-Plus, R, and SAS. The American Statistician, 59(3):275–276, August 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Proschan:2008:NAB

[Pro08] Michael A. Proschan. The normal approximation to the binomial. The American Statistician, 62(1):62–63, February 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Pearce:2006:PSR

[PW06] N. D. Pearce and M. P. Wand. Penalized splines and reproducing kernel methods. The American Statistician, 60(3):233–240, August 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Rovine:2004:PBA

[RA04] Michael J. Rovine and Douglas R. Anderson. Peirce and Bowditch: An American contribution to correlation and regression. The American Statistician, 58(3):232–236, August 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/27643563.

Richards:2007:BRT

[RACK07] Donald Richards, Christine M. Anderson-Cook, and Alina Kline. Brief reviews of teaching materials. The American Statistician, 61(1):102, February 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Rakovski:2006:RSD

[Rak06] Carter Cleveland Rakovski. Regression with social data: Modeling continuous and limited response variables. The American Statistician, 60(2):206, May 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Randles:2001:SPN

[Ran01] Ronald H. Randles. Statistical practice: On neutral responses (zeros) in the sign test and ties in the Wilcoxon–Mann–Whitney test. The American Statistician, 55(2):96–101, May 2001. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685994.

Rosoff:2005:BRT

[RD05] Jeffrey Rosoff and Carolyn Pillers Dobler. Brief reviews of teaching materials. The American Statistician, 59(4):352, November 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Reed:2000:BRB

[Ree00] Gina F. Reed. Book review: [untitled]. The American Statistician, 54(1):83, February 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685621.

Rees:2002:RBT

[Ree02] Kevin D. Rees. Reviews of books and teaching materials: Workshop Statistics: Discovery With Data, A Bayesian Approach. The American Statistician, 56(1):77–78, February 2002. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/3087336.

Reiss:2002:BRB

[RH02] Elayne Reiss and Lorrie L. Hoffman. Book review: The Elements of Statistics: With Applications to Economics and the Social Sciences. The American Statistician, 56(4):333–??, November 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v56n4/s20/p333.

Reich:2006:SAB

[RHCR06] Brian J. Reich, James S. Hodges, Bradley P. Carlin, and Adam M. Reich. A spatial analysis of basketball shot chart data. The American Statistician, 60(1):3–12, February 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Richardson:2008:WSD

[Ric08] Mary Richardson. Workshop statistics: Discovery with data. The American Statistician, 62(4):363, November 2008. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Ridout:2009:SAC

[Rid09] Martin S. Ridout. Statistical applications of the complex-step method of numerical differentiation. The American Statistician, 63(1):66–74, February 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Ries:2006:SMC

[Rie06] Lawrence D. Ries. Statistical methods for communication science. The American Statistician, 60(3):293, August 2006. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Rothman:2004:MDP

[RM04] David Rothman and David Mease. Mease, D. (2003), “A Penalized Maximum Likelihood Approach for the Ranking of College Football Teams Independent of Victory Margins,” “The American Statistician, 57, 241–248: Comment by Rothman and Reply. The American Statistician, 58(2):176, May 2004. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/27643542.

Roback:2003:BRB

[Rob03] Paul Roback. Book review: Weighing the Odds: A Course in Probability and Statistics. The American Statistician, 57(2):144– ??, May 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://oberon.ingentaselect.com/
cgi-bin/linker?ini=asa&reqidx=/cw/asa/00031305/v57n2/
s17/p144.

Robinson:2007:DEA

[Rob07a] Timothy J. Robinson. Design of experiments for agriculture and the natural sciences. The American Statistician, 61(3):275, August 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Robinson:2007:ICI

[Rob07b] Timothy J. Robinson. Introduction to computer-intensive methods of data analysis in biology. The American Statistician, 61(4):

371, November 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Robinson:2008:RB

[Rob08] Timothy J. Robinson. The R book. The American Statistician, 62(3):272–273, August 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Rodriguez:2004:FSD

[Rod04] Ricardo J. Rodriguez. First significant digit patterns from mixtures of uniform distributions. The American Statistician, 58 (1):64–71, February 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://oberon.

ingentaselect.com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/
00031305/v58n1/s13/p64.

Rosenkrantz:2000:CBQ

[Ros00a]  Walter A. Rosenkrantz.  Confidence bands for quantile

functions: a parametric and graphic alternative for testing goodness of fit. The American Statistician, 54(3):185– 190, August 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.amstat.org/

publications/tas/Rosenkrantz.htm;   http://www.jstor.
org/stable/2685588.

Rossman:2000:BRB

[Ros00b] Allan Rossman. Book review: Statistics You Can’t Trust: a Friendly Guide to Clear Thinking About Statistics in Everyday Life. The American Statistician, 54(2):151, May 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2686035.

Rosenbaum:2001:REB

[Ros01] Paul R. Rosenbaum. Replicating effects and biases. The American Statistician, 55(3):223–227, August 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2685805.

Rosenbaum:2003:ECI

[Ros03] Paul R. Rosenbaum. Exact confidence intervals for nonconstant effects by inverting the signed rank test. The American Statistician, 57(2):132–??, May 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://oberon.ingentaselect.com/cgi-bin/linker?
ini=asa&reqidx=/cw/asa/00031305/v57n2/s12/p132.

Rosenbaum:2005:HC

[Ros05] Paul R. Rosenbaum. Heterogeneity and causality. The American Statistician, 59(2):147–152, May 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Rosalsky:2007:SPP

[Ros07] Andrew Rosalsky. A simple and probabilistic proof of the binomial theorem. The American Statistician, 61(2):161–162, May 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Rowell:2002:BRB

[Row02] Ginger Holmes Rowell. Book review: Statistics for Psychologists: An Intermediate Course. The American Statistician, 56(3):248– ??, August 2002. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://oberon.ingentaselect.
com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/00031305/v56n3/
s16/p248.

Ritter:2001:APE

[RSH01] Mary Ann Ritter, Robert R. Starbuck, and Robert V. Hogg. Advice from prospective employers on training BS statisticians. The American Statistician, 55(1):14–18, February 2001. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685524.

Root:2001:TCC

[RT01] Rob Root and Trisha Thorme. Teacher’s corner: Communitybased projects in applied statistics: Using service-learning to enhance student understanding. The American Statistician, 55(4): 326–331, November 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685696.

Rinott:2003:MRR

[RT03] Yosef Rinott and Michael Tam. Monotone regrouping, regression, and Simpson’s Paradox. The American Statistician, 57(2):139– ??, May 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://oberon.ingentaselect.com/

cgi-bin/linker?ini=asa&reqidx=/cw/asa/00031305/v57n2/
s13/p139.

Rubin:2004:DGF

[Rub04a] Donald B. Rubin. The design of a general and flexible system for handling nonresponse in sample surveys. The American Statistician, 58(4):298–302, November 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/27643587.

Rubin:2004:ABS

[Rub04b] Donald B. Rubin. On advice for beginners in statistical research. The American Statistician, 58(3):196–197, August 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/27643552.

Rue:2004:BRB

[Rue04] Montserrat Rue. Book review: Statistical Methods for the Analysis of Biomedical Data, by Robert F. Woolson; William R. Clarke. The American Statistician, 58(2):174, May 2004. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/27643535.

Rumsey:2001:RBT

[Rum01] Deborah J. Rumsey. Reviews of books and teaching materials: Statistics and Data Analysis: From Elementary to Intermediate. The American Statistician, 55(3):255, August 2001. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685810.

Rumsey:2006:IBS

[Rum06] Deborah J. Rumsey. Introduction to Bayesian statistics. The American Statistician, 60(1):98–99, February 2006. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Rump:2007:CGS

[Rum07] Christopher M. Rump. Capital growth and the St. Petersburg Game. The American Statistician, 61(3):213–217, August 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Russo:2008:IPR

[Rus08] Ralph P. Russo. Intuitive probability and random processes using MATLAB. The American Statistician, 62(2):181–182, May 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Ryan:2009:ACI

[Rya09] Kenneth J. Ryan. Approximate confidence intervals for p when double sampling. The American Statistician, 63(2):132–140, May 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Reiczigel:2005:BTS

[RZR05] Jenő Reiczigel, Ildikó Zakariás, and Lajos Rózsa. A bootstrap test of stochastic equality of two populations. The American Statistician, 59(2):156–161, May 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Sansgiry:2000:TLD

[SA00] Prashant Sansgiry and Olcay Akman. Transformations of the lognormal distribution as a selection model. The American Statistician, 54(4):307–309, November 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.amstat.org/publications/tas/Sansgiry.htm;   http:
//www.jstor.org/stable/2685783.

Stokes:2002:ERb

[SA02] Lynne Stokes and James H. Albert. Editor’s reports. The American Statistician, 56(4):261–??, November 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v56n4/s1/p261.

Spiegelman:2003:NNM

[SA03] Clifford Spiegelman and Eun Sug Ark. Nearly nonparametric multivariate density estimates that incorporate marginal parametric density information. The American Statistician, 57(3): 183–??, August 2003. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://oberon.ingentaselect.
com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/00031305/v57n3/
s9/p183.

Sabatti:2002:MDV

[Sab02] Chiara Sabatti. Measuring dependency with volume tests. The American Statistician, 56(3):191–??, August 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v56n3/s5/p191.

Stefanski:2002:CE

[SB02] Leonard A. Stefanski and Dennis D. Boos. The calculus of Mestimation. The American Statistician, 56(1):29–38, February 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/3087324.

Smith:2009:MCD

[SB09] Marlene A. Smith and Peter G. Bryant. Managing case discussions in introductory business statistics classes: Practical approaches for instructors. The American Statistician, 63(4):348– 355, November 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Sellke:2001:CVT

[SBB01] Thomas Sellke, M. J. Bayarri, and James O. Berger. Calibration of p values for testing precise null hypotheses. The American Statistician, 55(1):62–71, February 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.amstat.org/publications/tas/Sellke.htm;   http://
www.jstor.org/stable/2685531;   http://www.tandfonline.
com/doi/abs/10.1198/000313001300339950.

Shi:2009:IMG

[SC09] Lei Shi and Gemai Chen. Influence measures for general linear models with correlated errors. The American Statistician, 63 (1):40–42, February 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Schaffner:2000:BRB

[Sch00] Andrew Schaffner. Book review: Revealing Statistical Principles. The American Statistician, 54(3):225–226, August 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685596.

Scheuren:2003:Ia

[Sch03a] Fritz Scheuren. Introduction. The American Statistician, 57(3): 189–??, August 2003. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://oberon.ingentaselect.
com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/00031305/v57n3/
s10/p189.

Scheuren:2003:Ib

[Sch03b] Fritz Scheuren. Introduction. The American Statistician, 57 (4):253–??, November 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://oberon.

ingentaselect.com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/
00031305/v57n4/s6/p253.

Scheuren:2003:LBW

[Sch03c] Fritz Scheuren. Looking back so we can look forward. The American Statistician, 57(2):94–??, May 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v57n2/s7/p94.

Scheuren:2004:Ia

[Sch04a] Fritz Scheuren. Introduction. The American Statistician, 58 (1):12–??, February 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://oberon.

ingentaselect.com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/
00031305/v58n1/s5/p12.

Scheuren:2004:Ib

[Sch04b] Fritz Scheuren. Introduction. The American Statistician, 58(2): 144, May 2004. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
27643523.

Scheuren:2004:Ic

[Sch04c] Fritz Scheuren. Introduction. The American Statistician, 58(3): 224, August 2004. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
27643561.

Scheuren:2004:Id

[Sch04d] Fritz Scheuren. Introduction. The American Statistician, 58(4): 290–291, November 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/27643584.

Schoolfield:2004:BRB

[Sch04e] Clyde H. Schoolfield, Jr. Book review: Elements of Stochastic Modeling, by K. Borovkov. The American Statistician, 58(4): 360–361, November 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/27643600.

Schwartzbaum:2004:BRB

[Sch04f] Judith Schwartzbaum. Book review: ActivEpi 1.1 (2nd ed.), by David G. Kleinbaum; ActivEpi Companion Textbook: A Supplement for Use with the ActivEpi CD-ROM by David G. Kleinbaum; Kevin M. Sullivan; Nancy D. Barker. The American Statistician, 58(4):359–360, November 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/27643598.

Scheuren:2005:Ia

[Sch05a] Fritz Scheuren. Introduction. The American Statistician, 59(1): 57, February 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Scheuren:2005:Ib

[Sch05b] Fritz Scheuren. Introduction. The American Statistician, 59(2): 162–163, May 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Scheuren:2005:Ic

[Sch05c] Fritz Scheuren. Introduction. The American Statistician, 59(3): 223, August 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Scheuren:2005:Id

[Sch05d] Fritz Scheuren. Introduction. The American Statistician, 59(4): 308, November 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Scheuren:2005:MI

[Sch05e] Fritz Scheuren. Multiple imputation. The American Statistician, 59(4):315–319, November 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Scheuren:2006:I

[Sch06a] Fritz Scheuren. Introduction. The American Statistician, 60 (1):32–33, February 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Schilling:2006:UPC

[Sch06b] Mark Schilling. Understanding probability: Chance rules in everyday life. The American Statistician, 60(1):97–98, February 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Satten:2001:KME

[SD01] Glen A. Satten and Somnath Datta. The Kaplan–Meier estimator as an inverse-probability-of-censoring weighted average. The American Statistician, 55(3):207–210, August 2001. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685801.

Sekar:2004:MEB

[SD04] C. Chandra Sekar and W. Edwards Deming. On a method of estimating birth and death rates and the extent of registration (excerpt). The American Statistician, 58(1):13–??, February 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v58n1/s6/p13.

Senn:2008:NCS

[Sen08] Stephen Senn. A note concerning a selection “paradox” of Dawid’s. The American Statistician, 62(3):206–210, August 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Sparks:2000:UDF

[SF00] Ross S. Sparks and John B. F. Field. Using Deming’s funnel experiment to demonstrate effects of violating assumptions underlying Shewhart’s control charts. The American Statistician, 54(4):291–302, November 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

amstat.org/publications/tas/Sparks.htm;   http://www.
jstor.org/stable/2685781.

Schall:2000:DBP

[SG00] Teddy Schall and Gary Smith General. Do baseball players regress toward the mean? The American Statistician, 54(4):231–235, November 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

amstat.org/publications/tas/Schall.htm;   http://www.
jstor.org/stable/2685772.

Schenker:2001:SPJ

[SG01] Nathaniel Schenker and Jane F. Gentleman. Statistical practice: On judging the significance of differences by examining the overlap between confidence intervals. The American Statistician, 55 (3):182–186, August 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685796.

Stroup:2004:TSP

[SGCS04] Donna F. Stroup, Richard A. Goodman, Ralph Cordell, and Richard Scheaffer. Teaching statistical principles using epidemiology: Measuring the health of populations. The American Statistician, 58(1):77–??, February 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://oberon.ingentaselect.com/cgi-bin/linker?
ini=asa&reqidx=/cw/asa/00031305/v58n1/s15/p77.

Schabenberger:2000:CSE

[SGK00] Oliver Schabenberger, Timothy G. Gregoire, and Fanzhi Kong. Collections of simple effects and their relationship to main effects and interactions in factorials. The American Statistician, 54(3):210–214, August 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.amstat.

org/publications/tas/Schabenberger.htm;   http://www.
jstor.org/stable/2685592.

Shepherd:2007:ECS

[SGM07] Bryan E. Shepherd, Peter B. Gilbert, and Devan V. Mehrotra. Eliciting a counterfactual sensitivity parameter. The American Statistician, 61(1):56–63, February 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Sharpe:2001:RBT

[Sha01] Norean Radke Sharpe. Reviews of books and teaching materials: Statistics for Management and Economics, Fifth Edition, by Gerald Keller; Brian Warrack. The American Statistician, 55(4): 370–371, November 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685704.

Stellman:2000:LE

[SHD^{+}00] Steven D. Stellman, David A. Harville, Michael F. Driscoll, H. J. Keselman, James Algina, Rhonda K. Kowalchuk, Ravindra Khattree, Ananda Sen, Robert M. Hamer, Lyle V. Jones, Bryan E. Adams, Amber G. Adams, Hui-Min Chang, Carol Etzel, Jo-Kang Kuo, Jorge Montemayor, and William R. Schucany. Letters to the Editor. The American Statistician, 54(2):156–160, May 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2686042.

Sherman:2007:PLN

[She07] Michael Sherman. Probabilities: The little numbers that rule our lives. The American Statistician, 61(4):372, November 2007. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Shieh:2001:IBC

[Shi01] Gwowen Shieh. The inequality between the coefficient of determination and the sum of squared simple correlation coefficients. The American Statistician, 55(2):121–124, May 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685998.

Shoemaker:2003:FTE

[Sho03] Lewis H. Shoemaker. Fixing the F test for equal variances. The American Statistician, 57(2):105–??, May 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v57n2/s9/p105.

Stasny:2004:SRT

[SHS04] Elizabeth A. Stasny, Michael Hamada, and Randy Sitter. Some random thoughts on doing research [with response]. The American Statistician, 58(3):200–202, August 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/27643557.

Shuster:2001:DSC

[Shu01] Jonathan Shuster. Dependence in simulation: a case study from the game of contract bridge. The American Statistician, 55(4): 306–309, November 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685692.

Shuster:2006:UTP

[Shu06] Jonathan J. Shuster. Using a two-player coin game paradox in the classroom. The American Statistician, 60(1):68–70, February 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Simon:2007:ARM

[Sim07] Gary Simon. Applied regression modeling: a business approach. The American Statistician, 61(4):367–368, November 2007. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Singpurwalla:2000:PPP

[Sin00] Nozer D. Singpurwalla. The point process paradox: Where should we extend the conversation? The American Statistician, 54(2):119–120, May 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

amstat.org/publications/tas/Singpurwalla.htm;   http:
//www.jstor.org/stable/2686029.

Sinha:2008:IBS

[Sin08] Samiran Sinha. Introduction to Bayesian statistics. The American Statistician, 62(3):268, August 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Sisson:2007:G

[Sis07] S. A. Sisson. Genetics. The American Statistician, 61(2):112–119, May 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Snapinn:2005:IIT

[SJI05] Steven M. Snapinn, Qi Jiang, and Boris Iglewicz. Illustrating the impact of a time-varying covariate with an extended Kaplan–Meier estimator. The American Statistician, 59(4):301– 307, November 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Sklar:2007:UBS

[Skl07] Jeffrey C. Sklar. Understanding business statistics. The American Statistician, 61(4):369–370, November 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Skoulakis:2008:RFC

[Sko08] Georgios Skoulakis. A recursive formula for computing central moments of a multivariate lognormal distribution. The American Statistician, 62(2):147–150, May 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Steiner:2009:TVR

[SM09] Stefan H. Steiner and R. Jock MacKay. Teaching variation reduction using a virtual manufacturing environment. The American Statistician, 63(4):361–365, November 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Schenker:2007:CFF

[SMC^{+}07] Nathaniel Schenker, Katherine L. Monti, George W. Cobb, Ronald S. Fecso, and Joan S. Chmiel. Combining features of a frequency table and a stem-and-leaf plot to summarize the American Statistical Association’s strategic activities. The American Statistician, 61(3):245–247, August 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Stern:2004:SCF

[SMH^{+}04] Hal S. Stern, Kenneth Massey, David A. Harville, Richard Billingsley, Wes Colley, and David Mease. Statistics and the college football championship [with discussions, with reply]. The American Statistician, 58(3):179–185, August 2004. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/27643551.

Serroyen:2009:NML

[SMVD09] Jan Serroyen, Geert Molenberghs, Geert Verbeke, and Marie Davidian. Nonlinear models for longitudinal data. The American Statistician, 63(4):378–388, November 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Snee:2008:WED

[Sne08] Ronald D. Snee. W. Edwards Deming’s “Making AnotherWorld”. The American Statistician, 62(3):251–255, August 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Snyder:2005:PP

[Sny05] Carl Snyder. The problem of prosperity. The American Statistician, 59(1):62–67, February 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Solo:2000:SDS

[Sol00] V. Solo. A simple derivation of the smoothing spline. The American Statistician, 54(1):40–43, February 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2685609.

Sen:2001:LE

[SOS^{+}01] Ananda Sen, W. Jason Owen, Sunil K. Sapra, Robert M. Price, Jr., and Alan Agresti. Letters to the Editor. The American Statistician, 55(3):260–261, August 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2685816.

Sowey:2003:GWE

[Sow03] Eric R. Sowey. The getting of wisdom: Educating statisticians to enhance their clients’ numeracy. The American Statistician, 57 (2):89–??, May 2003. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://oberon.ingentaselect.
com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/00031305/v57n2/
s5/p89.

Sparks:2008:MD

[Spa08] JonDavid Sparks. Missing data. The American Statistician, 62 (3):269–270, August 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Speed:2008:SPE

[Spe08] F. Michael Speed. SAS programming for enterprise guide users. The American Statistician, 62(4):362, November 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Spirling:2007:TPI

[Spi07] Arthur Spirling. “turning points” in the Iraq conflict. The American Statistician, 61(4):315–320, November 2007. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Szekely:2004:SPP

[SR04] Gábor J. Székely and Donald St. P. Richards. The St. Petersburg Paradox and the crash of high-tech stocks in 2000. The American Statistician, 58(3):225–231, August 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/27643562.

Szekely:2005:RRS

[SR05] Gabor J. Székely and Donald St. P. Richards. Response: Remain steadfast with the St. Petersburg Paradox to quantify irrational exuberance. The American Statistician, 59(3):235–239, August 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Strug:2007:IES

[SRC07] Lisa J. Strug, Charles A. Rohde, and Paul N. Corey. An introduction to evidential sample size calculations. The American Statistician, 61(3):207–212, August 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Speckman:2008:DPC

[SRMP08] Paul L. Speckman, Jeffrey N. Rouder, Richard D. Morey, and Michael S. Pratte. Delta plots and coherent distribution ordering. The American Statistician, 62(3):262–266, August 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Singpurwalla:2001:NRB

[SS01] Nozer D. Singpurwalla and Andrew Swift. Network reliability and Borel’s paradox. The American Statistician, 55(3):213–218, August 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685803.

Sinharay:2002:SBF

[SS02] Sandip Sinharay and Hal S. Stern. On the sensitivity of Bayes factors to the prior distributions. The American Statistician, 56(3): 196–??, August 2002. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://oberon.ingentaselect.
com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/00031305/v56n3/
s6/p196.

Scheaffer:2004:SUE

[SS04] Richard L. Scheaffer and Elizabeth A. Stasny. The state of undergraduate education in statistics: a report from the CBMS 2000. The American Statistician, 58(4):265–271, November 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/27643581.

Solow:2005:HSN

[SS05] Andrew R. Solow and Woollcott Smith. How surprising is a new record? The American Statistician, 59(2):153–155, May 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Simon:2006:LL

[SS06] Gary A. Simon and Jeffrey S. Simonoff. “last licks”. The American Statistician, 60(1):13–18, February 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Scott:2008:SHF

[SS08] David W. Scott and Warren R. Scott. Smoothed histograms for frequency data on irregular intervals. The American Statistician, 62(3):256–261, August 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Solow:2009:SPC

[SS09] Andrew R. Solow and Woollcott K. Smith. A statistical problem concerning the Mar Saba letter. The American Statistician, 63 (3):254–257, August 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Stasny:2001:HGJ

[Sta01] Elizabeth A. Stasny. How to get a job in academics. The American Statistician, 55(1):35–40, February 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.amstat.org/publications/tas/Stasny.htm;   http://
www.jstor.org/stable/2685527.

Stamey:2008:MMS

[Sta08] James Stamey. Modern mathematical statistics with applications. The American Statistician, 62(4):358, November 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Stephenson:2001:BRB

[Ste01] W. Robert Stephenson. Book review: Introduction to Probability and Statistics for Engineers and Scientists (2nd ed.). The American Statistician, 55(1):81–82, February 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2685535.

Stefanski:2007:RR

[Ste07] Leonard A. Stefanski. Residual (sur)realism. The American Statistician, 61(2):163–177, May 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Stefanski:2008:NCL

[Ste08] Leonard A. Stefanski. The North Carolina Lottery coincidence. The American Statistician, 62(2):130–134, May 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Stine:2000:BRB

[Sti00] Robert Stine. Book review: Statistics with Mathematica. The American Statistician, 54(2):153, May 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2686038.

Stigler:2008:SHT

[Sti08] Stephen M. Stigler. Statistics at Harvard? A toast to the Harvard Statistics Department. The American Statistician, 62(3):246–250, August 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Stokes:2001:ER

[Sto01] Lynne Stokes. Editor’s report. The American Statistician, 55(2): 87–88, May 2001. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685992.

Stokes:2002:ERa

[Sto02] Lynne Stokes. Editor’s reports. The American Statistician, 56(2): 83–??, May 2002. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://oberon.ingentaselect.
com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/00031305/v56n2/
s1/p83.

Strickland:2002:RBT

[Str02] Pamela A. Ohman Strickland. Reviews of books and teaching materials: SPSS for Windows: An Introduction to Use and Interpretation in Research. The American Statistician, 56(1):77, February 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/3087335.

Stroup:2004:CTS

[Str04] Donna L. Stroup. Corrections: Teaching statistical principles using epidemiology: Measuring the health of populations. The American Statistician, 58(2):178, May 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/27643547.

Stufken:2009:IEI

[Stu09] John Stufken. An invitation: Editor’s introduction to “Desired and Feared — What Do We Do Now and Over the Next 50 Years?”. The American Statistician, 63(3):201, August 2009. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Su:2008:DMM

[Su08] Xiaogang Su. Data mining methods and models. The American Statistician, 62(1):91, February 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Symanzik:2003:TEC

[SV03] Jürgen Symanzik and Natascha Vukasinovic. Teaching experiences with a course on “Web-based statistics”. The American Statistician, 57(1):46–??, February 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://oberon.ingentaselect.com/cgi-bin/linker?
ini=asa&reqidx=/cw/asa/00031305/v57n1/s7/p46.

Spence:2005:WPH

[SW05] Ian Spence and Howard Wainer. William Playfair and his graphical inventions: An excerpt from the introduction to the republication of his atlas and statistical breviary. The American Statistician, 59(3):224–229, August 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Schilling:2002:HHB

[SWW02] Mark F. Schilling, Ann E. Watkins, and William Watkins. Is human height bimodal? The American Statistician, 56(3):223–??, August 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://oberon.ingentaselect.com/
cgi-bin/linker?ini=asa&reqidx=/cw/asa/00031305/v56n3/
s10/p223.

Strasak:2007:USM

[SZM^{+}07] Alexander M. Strasak, Qamruz Zaman, Gerhard Marinell, Karl P. Pfeiffer, and Hanno Ulmer. The use of statistics in medical research. The American Statistician, 61(1):47–55, February 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Taleb:2007:BSD

[Tal07] Nassim Nicholas Taleb. Black swans and the domains of statistics. The American Statistician, 61(3):198–200, August 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Tang:2002:BRB

[Tan02] Boxin Tang. Book review: Experimental Design With Applications in Management, Engineering, and the Sciences. The American Statistician, 56(4):333–??, November 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v56n4/s21/p333.

Tarpey:2000:NPS

[Tar00] Thaddeus Tarpey. A note on the prediction sum of squares statistic for restricted least squares. The American Statistician, 54(2):116–118, May 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

amstat.org/publications/tas/Tarpey.htm;   http://www.
jstor.org/stable/2686028.

Tarpey:2007:LTM

[Tar07] Thaddeus Tarpey. Linear transformations and the k-means clustering algorithm. The American Statistician, 61(1):34–40, February 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Terrell:2003:BRB

[Ter03] George Terrell. Book review: All the Mathematics You Missed [But Need to Know for Graduate School]. The American Statistician, 57(1):69–??, February 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://oberon.ingentaselect.com/cgi-bin/linker?
ini=asa&reqidx=/cw/asa/00031305/v57n1/s17/p69.

Terpstra:2005:SIC

[Ter05] Jeff T. Terpstra. Some illustrative classroom examples regarding sums of discrete random variables with finite support. The American Statistician, 59(3):258–265, August 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Tuyl:2008:CBL

[TGM08] Frank Tuyl, Richard Gerlach, and Kerrie Mengersen. A comparison of Bayes–Laplace, Jeffreys, and other priors. The American Statistician, 62(1):40–44, February 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Tarpey:2000:SB

[TH00] Thaddeus Tarpey and John Holcomb. Spline bottles. The American Statistician, 54(2):129–135, May 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.amstat.org/publications/tas/Tarpey2.htm;   http:/
/www.jstor.org/stable/2686031.

Thompson:2004:IW

[Tho04] Elizabeth Thompson. The importance of “why?”. The American Statistician, 58(3):198, August 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/27643554.

Timm:2004:EES

[Tim04] Neil H. Timm. Estimating effect sizes in exploratory experimental studies when using a linear model. The American Statistician, 58 (3):213–217, August 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/27643559.

Tjur:2009:CDL

[Tju09] Tue Tjur. Coefficients of determination in logistic regression models — a new proposal: The coefficient of discrimination. The American Statistician, 63(4):366–372, November 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Troendle:2004:ESC

[TKM04] James F. Troendle, Edward L. Korn, and Lisa M. McShane. An example of slow convergence of the bootstrap in high dimensions. The American Statistician, 58(1):25–??, February 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v58n1/s8/p25.

Tobias:2004:AAS

[Tob04] Justin L. Tobias. Assessing assessments of school performance: The case of California. The American Statistician, 58(1):55–??, February 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://oberon.ingentaselect.com/
cgi-bin/linker?ini=asa&reqidx=/cw/asa/00031305/v58n1/
s12/p55.

Torrieri:2007:ACC

[Tor07] Nancy K. Torrieri. America is changing, and so is the Census. The American Statistician, 61(1):16–21, February 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Tra:2007:EYC

[Tra07] Yolande Tra. Even you can learn statistics: a guide for everyone who has ever been afraid of statistics. The American Statistician, 61(2):182, May 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Theoharakis:2003:HDS

[TS03] Vasilis Theoharakis and Mary Skordia. How do statisticians perceive statistics journals? The American Statistician, 57(2):115– ??, May 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://oberon.ingentaselect.com/
cgi-bin/linker?ini=asa&reqidx=/cw/asa/00031305/v57n2/
s10/p115.

Tan:2009:SSS

[TTT09] Ming T. Tan, Guo-Liang Tian, and Man-Lai Tang. Sample surveys with sensitive questions: a nonrandomized response approach. The American Statistician, 63(1):9–16, February 2009.

CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Turiel:2007:QRB

[Tur07] Thomas P. Turiel. Quantum random bit generators. The American Statistician, 61(3):255–259, August 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Tuyl:2009:LE

[Tuy09] Frank Tuyl. Letter to the Editor. The American Statistician, 63 (3):294, August 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wu:2001:CGM

[tWGB01] Chi tsung Wu, Marcia L. Gumpertz, and Dennis D. Boos. Comparison of GEE, MINQUE, ML, and REML estimating equations for normally distributed data. The American Statistician, 55(2): 125–130, May 2001. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685999.

Utts:2008:ETB

[UJ08] Jessica Utts and Wesley Johnson. The evolution of teaching Bayesian statistics to nonstatisticians. The American Statistician, 62(3):199–201, August 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Upton:2000:CIM

[Upt00] Graham J. G. Upton. Conditional independence, Mantel– Haenszel test, and the Yates correction. The American Statistician, 54(2):112–115, May 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:/

/www.amstat.org/publications/tas/Upton.htm;   http://
www.jstor.org/stable/2686027.

Upton:2001:TSD

[Upt01] Graham J. G. Upton. A toroidal scatter diagram for ternary variables. The American Statistician, 55(3):240–243, August 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2685808.

Utts:2002:BRB

[Utt02] Jessica Utts. Book review: News and Numbers: A Guide to Reporting Statistical Claims and Controversies in Health and Related Fields, (2nd Ed.). The American Statistician, 56(4):330–??, November 2002. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://oberon.ingentaselect.
com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/00031305/v56n4/
s15/p330.

Utts:2003:SSS

[Utt03a] Jessica Utts. Special section on statistical literacy: How you can help. The American Statistician, 57(2):73–??, May 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v57n2/s1/p73.

Utts:2003:WEC

[Utt03b] Jessica Utts. What educated citizens should know about statistics and probability. The American Statistician, 57(2):74–??, May 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://oberon.ingentaselect.com/
cgi-bin/linker?ini=asa&reqidx=/cw/asa/00031305/v57n2/
s2/p74.

Verrill:2005:DFT

[VD05] Steve Verrill and Mark Durst. The decline and fall of Type II error rates. The American Statistician, 59(4):287–291, November 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

vanDorp:2002:STS

[vDK02] J. René van Dorp and Samuel Kotz. The standard two-sided power distribution and its properties: With applications in financial engineering. The American Statistician, 56(2):90–??, May 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://oberon.ingentaselect.com/
cgi-bin/linker?ini=asa&reqidx=/cw/asa/00031305/v56n2/
s3/p90.

Villa:2006:UMG

[VE06] Enrique R. Villa and Luis A. Escobar. Using moment generating functions to derive mixture distributions. The American Statistician, 60(1):75–80, February 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Verrall:2001:BRB

[Ver01] Richard Verrall. Book review: Introduction to Statistics With Applications in General Insurance. The American Statistician, 55(1):84, February 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685539.

vonHippel:2004:BSM

[vH04] Paul T. von Hippel. Biases in SPSS 12.0 missing value analysis. The American Statistician, 58(2):160–164, May 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/27643528.

Vos:2005:ECD

[VH05] Paul W. Vos and Suzanne Hudson. Evaluation criteria for discrete confidence intervals. The American Statistician, 59(2):137–142, May 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Vickers:2008:DAE

[Vic08] Andrew J. Vickers. Decision analysis for the evaluation of diagnostic tests, prediction models, and molecular markers. The American Statistician, 62(4):314–320, November 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Vardeman:2003:SES

[VM03] Stephen B. Vardeman and Max D. Morris. Statistics and ethics: Some advice for young statisticians. The American Statistician, 57(1):21–??, February 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://oberon.

ingentaselect.com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/
00031305/v57n1/s3/p21.

Verbeke:2007:WCG

[VM07] Geert Verbeke and Geert Molenberghs. What can go wrong with the score test? The American Statistician, 61(4):289–290, November 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Verbeke:2001:CLM

[VSL01] Geert Verbeke, Bart Spiessens, and Emmanuel Lesaffre. Conditional linear mixed models. The American Statistician, 55 (1):25–34, February 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.amstat.org/

publications/tas/Verbeke.htm;   http://www.jstor.org/
stable/2685526.

Waldman:2000:EDC

[Wal00a] Donald M. Waldman. Estimation in discrete choice models with choice-based samples. The American Statistician, 54(4): 303–306, November 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.amstat.org/

publications/tas/Waldman.htm;   http://www.jstor.org/
stable/2685782.

Wallace:2000:BRB

[Wal00b] Edward C. Wallace. Book review: A Basic Course in Statistics, 4th edition. The American Statistician, 54(2):151, May 2000. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2686036.

Walker:2002:BRB

[Wal02] John H. Walker. Book review: Contemporary Business Statistics with Microsoft Excel. The American Statistician, 56(3):250–??, August 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://oberon.ingentaselect.com/
cgi-bin/linker?ini=asa&reqidx=/cw/asa/00031305/v56n3/
s19/p250.

Walker:2005:RAS

[Wal05] Helen M. Walker. The role of the American Statistical Association. The American Statistician, 59(1):68–71, February 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wang:2000:FIW

[Wan00] Y. H. Wang. Fiducial intervals: What are they? The American Statistician, 54(2):105–111, May 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.amstat.org/publications/tas/Wang.htm;   http://
www.jstor.org/stable/2686026.

Wand:2002:VDC

[Wan02] M. P. Wand. Vector differential calculus in statistics. The American Statistician, 56(1):55–62, February 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/3087328.

Warner:2002:RBT

[War02] Bradley A. Warner. Reviews of books and teaching materials: Basic Engineering Data Collection and Analysis, by Stephen B. Vardeman; J. Marcus Jobe. The American Statistician, 56(1):76– 77, February 2002. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
3087333.

Wasserman:2006:ACD

[Was06] Stanley Wasserman. Analyzing categorical data. The American Statistician, 60(3):291–292, August 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

White:2002:TPR

[WB02] Chris White and Scott Berry. Tiered polychotomous regression: Ranking NFL quarterbacks. The American Statistician, 56(1):10– 21, February 2002. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
3087322.

Wainer:2004:TSP

[WB04] Howard Wainer and Lisa M. Brown. Two statistical paradoxes in the interpretation of group differences: Illustrated with medical school admission and licensing data. The American Statistician, 58(2):117–123, May 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/27643519.

Wall:2001:ECI

[WBT01] Melanie M. Wall, James Boen, and Richard Tweedie. An effective confidence interval for the mean with samples of size one and two. The American Statistician, 55(2):102–105, May 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685995.

Wickham:2007:RG

[WC07] Hadley Wickham and Dianne Cook. R graphics. The American Statistician, 61(1):99–100, February 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Webber:2004:CHL

[WCLS04] W. Fred Webber, Hwan Chung, Eric Loken, and Joseph L. Schafer. Chung, H., Loken, E., and Schafer, J. L. (2004), “Difficulties in Drawing Inferences with Finite-Mixture Models: A Simple Example with a Simple Solution,” “The American Statistician”, 58, 152–158: Comment by Webber and Reply. The American Statistician, 58(4):362–363, November 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/27643602.

Whiteside:2008:BEN

[WE08] Mary M. Whiteside and Mark E. Eakin. A better estimate of the number of valid signatures on a petition. The American Statistician, 62(1):17–21, February 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Weir:2007:ISF

[Wei07] Bruce Weir. Introduction to statistics for forensic scientists. The American Statistician, 61(3):274, August 2007. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Westfall:2006:E

[Wes06] Peter Westfall. Editorial. The American Statistician, 60(1):1– 2, February 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Westfall:2007:ER

[Wes07] Peter H. Westfall. Editor’s report. The American Statistician, 61(2):103–104, May 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Westfall:2008:EN

[Wes08] Peter H. Westfall. Editor’s note. The American Statistician, 62 (3):189, August 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wilkinson:2009:HCH

[WF09] Leland Wilkinson and Michael Friendly. The history of the cluster heat map. The American Statistician, 63(2):179–184, May 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Woodward:2009:EFB

[WGHE09] Wayne A. Woodward, Henry L. Gray, James R. Haney, and Alan C. Elliott. Examining factors to better understand autoregressive models. The American Statistician, 63(4):335–342, November 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Westfall:2007:BS

[WH07] Peter H. Westfall and Joseph M. Hilbe. The black swan. The American Statistician, 61(3):193–194, August 2007. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wiegand:2006:SAM

[Wie06] Ryan E. Wiegand. Statistical analysis of medical data using SAS. The American Statistician, 60(4):345–346, November 2006. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wiens:2007:HSD

[Wie07] Brian Wiens. Handbook of statistical distributions with applications. The American Statistician, 61(4):370, November 2007. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wiegand:2008:DPA

[Wie08] Ryan E. Wiegand. Data preparation for analytics using SAS. The American Statistician, 62(4):356–357, November 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Williams:2002:BRB

[Wil02] Calvin Williams. Book review: Applied Multivariate Data Analysis (2nd Edition). The American Statistician, 56(3):248–??, August 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://oberon.ingentaselect.com/
cgi-bin/linker?ini=asa&reqidx=/cw/asa/00031305/v56n3/
s17/p248.

Wilkinson:2006:RPC

[Wil06a] Leland Wilkinson. Revising the Pareto chart. The American Statistician, 60(4):332–334, November 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wilks:2006:USE

[Wil06b] S. S. Wilks. Undergraduate statistical education. The American Statistician, 60(1):39–45, February 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Willink:2008:UPN

[Wil08] Robin Willink. A unique property of the normal distribution associated with perturbing a general random variable. The American Statistician, 62(2):144–146, May 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wolter:2003:RUB

[WJM^{+}03] Kirk Wolter, Diana Jergovic, Whitney Moore, Joe Murphy, and Colm O’Muircheartaigh. Reliability of the uncertified ballots in the 2000 Presidential Election in Florida. The American Statistician, 57(1):1–??, February 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v57n1/s1/p1.

Williamson:2009:RCO

[WMAY09] Patricia Pepple Williamson, Darcy P. Mays, Ghidewon Abay Asmerom, and Yingying Yang. Revisiting the classical occupancy problem. The American Statistician, 63(4):356–360, November 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Withers:2007:RRM

[WN07] Christopher S. Withers and Saralees Nadarajah. A recurrence relation for moments of the noncentral chi square. The American Statistician, 61(4):337–338, November 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wood:2008:CFA

[Woo08] Phil Wood. Confirmatory factor analysis for applied research. The American Statistician, 62(1):91–92, February 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wendell:2001:LCI

[WS01] John P. Wendell and Josef Schmee. Likelihood confidence intervals for proportions in finite populations. The American Statistician, 55(1):55–61, February 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.amstat.

org/publications/tas/Wendell.htm;   http://www.jstor.
org/stable/2685530.

Winkler:2002:SPR

[WSF02] Robert L. Winkler, James E. Smith, and Dennis G. Fryback. Statistical practice: The role of informative priors in zero-numerator problems: Being conservative versus being candid. The American Statistician, 56(1):1–4, February 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/3087320.

Wulff:2003:BRB

[Wul03] Shaun S. Wulff. Book review: A First Course in Design and Analysis of Experiments. The American Statistician, 57(1):66–??, February 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://oberon.ingentaselect.com/
cgi-bin/linker?ini=asa&reqidx=/cw/asa/00031305/v57n1/
s12/p66.

Wulff:2007:SMM

[Wul07] Shaun S. Wulff. SAS for mixed models. The American Statistician, 61(2):184–185, May 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wuthrich:2008:SPM

[Wüt08] Mario V. Wüthrich. Statistical and probabilistic methods in actuarial science. The American Statistician, 62(4):360, November 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wachsmuth:2003:GBP

[WWD03] Amanda Wachsmuth, Leland Wilkinson, and Gerard E. Dallal. Galton’s bend: a previously undiscovered nonlinearity in Galton’s family stature regression data. The American Statistician, 57(3): 190–??, August 2003. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://oberon.ingentaselect.
com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/00031305/v57n3/
s11/p190.

Wu:2004:ECD

[WY04] Samuel S. Wu and Mark C. K. Yang. Evaluation of the current decision rule in figure skating and possible improvements. The American Statistician, 58(1):46–??, February 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v58n1/s11/p46.

Xie:2003:SAD

[XN03] Rongrong Xie and Paul I. Nelson. Separation among distributions related by linear regression. The American Statistician, 57(1):33– ??, February 2003. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://oberon.ingentaselect.
com/cgi-bin/linker?ini=asa&reqidx=/cw/asa/00031305/v57n1/
s5/p33.

Yu:2009:IPI

[YA09] Keming Yu and Abdallah Ally. Improving prediction intervals: Some elementary methods. The American Statistician, 63(1):17– 19, February 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Yalta:2007:NRG

[Yal07] A. Talha Yalta. The numerical reliability of GAUSS 8.0. The American Statistician, 61(3):262–268, August 2007. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Yu:2006:SES

[YCF06] Philip L. H. Yu, Jennifer S. K. Chan, and Wing K. Fung. Statistical exploration from SARS. The American Statistician, 60 (1):81–91, February 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Yucel:2008:UCI

[YHZ08] Recai M. Yucel, Yulei He, and Alan M. Zaslavsky. Using calibration to improve rounding in imputation. The American Statistician, 62(2):125–129, May 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Young:2005:NSW

[You05] Allyn A. Young. National statistics in war and peace. The American Statistician, 59(1):58–61, February 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Yang:2001:ESE

[YT01] Li Yang and Anastasios A. Tsiatis. Efficiency study of estimators for a treatment effect in a pretest-posttest trial. The American Statistician, 55(4):314–321, November 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2685694.

Yozgatligil:2009:RMS

[YW09] Ceylan Yozgatligil and William W. S. Wei. Representation of multiplicative seasonal vector autoregressive moving average models. The American Statistician, 63(4):328–334, November 2009. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Zamba:2008:SDQ

[Zam08] Gideon (K. D.) Zamba. Statistical development of quality in medicine. The American Statistician, 62(4):359–360, November 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Zhou:2000:OSC

[ZG00] Xiao-Hua Zhou and Sujuan Gao. One-sided confidence intervals for means of positively skewed distributions. The American Statistician, 54(2):100–104, May 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.amstat.org/publications/tas/Zhou.htm;   http://
www.jstor.org/stable/2686025.

Zhang:2000:PSC

[Zha00] Li-Chun Zhang. Post-stratification and calibration — a synthesis. The American Statistician, 54(3):178–184, August 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.amstat.org/publications/tas/Li-
Chun.htm;   http://www.jstor.org/stable/2685587.

Zhang:2007:SMS

[Zha07a] Lingyun Zhang. Sample mean and sample variance. The American Statistician, 61(2):159–160, May 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Zhang:2007:FB

[Zha07b] Ying Zhang. Fundamentals of biostatistics. The American Statistician, 61(2):183, May 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Zhou:2001:TCU

[Zho01] Mai Zhou. Teacher’s corner: Understanding the Cox regression models with time-change covariates. The American Statistician, 55(2):153–155, May 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2686004.

Zhu:2008:KE

[Zhu08] Mu Zhu. Kernels and ensembles. The American Statistician, 62(2):97–109, May 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Zimmerman:2000:SCS

[Zim00] Dale L. Zimmerman. Statistical computing software reviews: Viewing the correlation structure of longitudinal data through a PRISM. The American Statistician, 54(4):310–318, November 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.amstat.org/publications/
tas/Zimmerman.htm;   http://www.jstor.org/stable/2685784.

Zimmerman:2007:SPD

[Zim07] M. Bridget Zimmerman. SAS programming and data analysis: a theory and program-driven approach. The American Statistician, 61(3):277–278, August 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Zhou:2009:RMR

[ZL09] Hua Zhou and Kenneth Lange. Rating movies and rating the raters who rate them. The American Statistician, 63(4):297–307, November 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Zaslavsky:2002:CZK

[ZvD02] Alan M. Zaslavsky and David A. van Dyk. Chen, Z., and Kuo, L. (2001), “A Note on the Estimation of the Multinomial Logit Model with Random Effects,” The American Statistician, 55, 89– 95: Comment by Zaslavsky and van Dyk. The American Statistician, 56(1):80–81, February 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/3087338.